### v2

# Avoiding strict saddle points of nonconvex regularized problems

Abstract
In this paper, we consider a class of non-convex and non-smooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms. These problems, moreover, are probably non-lipschitz around sparse solutions. We propose an damped iterative reweighted $\ell_{1}$ algorithm to solve these problems. The algorithm is guaranteed to converge only to local minimizers when randomly initialized. By deriving a second-order sufficient optimal condition under the non-degenerate assumption, we show that the strict saddle property is generic on these sparse optimization problems.
keywords:
strict saddle, iterative reweighted $\ell_{1}$ algorithm, nonconvex regularization, center stable manifold theorem
1 Introduction
In this paper, we consider the following nonsmooth regularization problem,
$\min\limits_{x\in\mathbb{R}^{n}}F(x):=f(x)+\lambda\psi(x),$ (P)
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and the penalty/regularization term $\psi(x)$ takes the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$. We assume the objective $F$ has bounded level sets, the data fitting function $f$ is twice continuously differentiable and possibly nonconvex, and $r$ satisfies the following properties.
Assumption 1.
$r$ is continuous on $[0,+\infty)$ with $r(0)=0$, $C^{2}$-smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
This assumption indicates $r(t)>0$ for all $t>0$, but $r$ could be nonconvex or not Lipschitz around 0.
The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance. This makes it well-suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [1, 2, 3, 4]. Many widely used nonconvex regularization such as the smoothly clipped absolute deviation penalty (SCAD) [5] or the minimax concave penalty (MCP) [6], the $\ell_{p}$ quasi-norm penalty [7] satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability. We provide the explicit form of several commonly used nonconvex regularization in Table 1. It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizer with bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) non-Lipschitz continuous regularizer with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as $\ell_{p}$ quasi-norm (LPN).
Numerous studies focus on gradient-based optimization methods that are guaranteed to find first-order stationary points. However, the nonconvexity of $F$ can lead to multiple local minima and saddle points, which can cause algorithms to converge to stationary points rather than the (local) optimum. Additionally, the presence of saddle points can significantly slow down the convergence of optimization algorithms, especially gradient-based methods. At a saddle point, the gradient provides misleading information about the descent directions. As a result, avoiding saddle points can help optimization algorithms more effectively converge to better solutions. Therefore, it becomes a prevalent topic, particularly in high-dimensional optimization problems, and is also the central focus of this paper.
In smooth optimization, many methods have been developed and analyzed to escape from saddle points, including first-order methods [8, 9, 10], momentum-based methods [11], second-order optimization algorithms [12], the algorithmic frameworks for constrained nonlinear problems [13], stochastic algorithms [14, 15], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [16, 17, 18, 19]. In smooth settings, it is widely assumed that the strict saddle property holds, meaning all saddle points possess at least one strictly negative eigenvalue. This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases and has been shown to be generally satisfied for smooth problems [20, 21].
In nonsmooth optimization, avoiding saddle points poses a greater challenge. At stationary point, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate the avoidance of saddle points. Developing nonsmooth optimization algorithms capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than in smooth optimization. An example of this complexity is demonstrated in the analysis of the Douglas splitting algorithm for nonsmooth problems, as discussed in [22]. Recent research [23] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point. Additionally, several proximal methods, including proximal point method, proximal gradient method and proximal linear method [23], as well as a more efficient algorithm incorporating perturbation [24], have been proposed to converge to local minima in nonsmooth optimization settings. In these studies, the nonsmooth objective function $r(x)$ is assumed to be $\rho$-weakly convex meaning $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. They assume an active-manifold around the limit point and then define the strict saddle points as those of the $f$ restricted to the manifold. The weakly $\rho$-convexity for $r$ plays a crucial role in constructing C2 differentiable Moreau envelope with the same critical points and strict saddle points in proximal methods.
As for the regularized problem (P), several first-order methods have been proposed. [25] introduces a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1. This method is subsequently extended to address group $\ell_{p,q}$ norm regularized problems in [26]. Furthermore, the Difference of Convex (DC) algorithm is enlisted [27] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions. It is important to note, however, that the DC algorithm generally assumes $r^{\prime}(|0^{+}|)<\infty$ to accommodate the DC decomposition, which does not hold true for the $\ell_{p}$ regularized problem. Another avenue of research involves iteratively reweighted algorithms, which approximate the nonconvex regularizers by iteratively reweighted $\ell_{1}$ or $\ell_{2}$ regularizers. Initially applied to solve the unconstrained $\ell_{p}$ regularized problem in [28] without convergence analysis, this approach was later employed to address general constrained nonconvex regularization problems [29]. Subsequent work in [30] established its global convergence and demonstrated its locally stable support property [31]. Furthermore, convergence rate analysis under the Kurdyka-Łojasiewicz property [32] and its extrapolated version was proposed in [33]. While these first-order methods primarily focus on identifying first-order critical points, none of them guarantee avoidance of saddle points.
1.1 The main results
In this paper, our main purpose is to develop an iterative reweighted $\ell_{1}$ algorithm that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weakly convexity. The additional assumption for $r$ needed in our work to guarantee that the existence of an active manifold around an active manifold as assumed by [23].
Assumption 2.
For any stationary point $x^{*}$, $0\in{\rm rint}\partial F(x^{*})$ (where $\partial F$ denotes the regular subdifferential and ${\rm rint}$ denotes the relative interior).
The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) is widely applicable to these non-convex regularization problems. In contrast to the proximal gradient algorithm, ${\rm IRL}_{1}$ solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator. This method offers a favorable local property [31], known as the model identification. At the tail end of the iteration, the sign of the iterates remain unchanged while the nonzero components are kept uniformly bounded away from 0. In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{J}$ as follows after finite number of iterations.
Definition 1.
(Model Subspace) The model subspace defined by $\mathcal{J}\subset\{1,\ldots,n\}$ is
$\mathcal{M}_{\mathcal{J}}:=\{u\in\mathbb{R}^{n}:\forall j\in\mathcal{J},u_{j}=0\}.$
This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{J}}$. If we think the iteration of ${\rm IRL}_{1}$ as a mapping that assigns to every point another point. This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) on $\mathcal{M}_{\mathcal{J}}$, even though the underlying optimization problem is nonsmooth in the entire space.
We can summarize our contribution as follows:
  1. (i)
    We first show the local minimal and second-order optimality of $f$ are equivalent to those of $f$ restricted on $\mathcal{M}_{\mathcal{J}}$, and define the strict saddle points as those on $\mathcal{M}_{\mathcal{J}}$.
  2. (ii)
    We develop damped IRL1 algorithms individually for the two types of regularizers, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning the next iterate is the combination of a traditional IRL1 iterate and its previous iterate. The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ becomes $\infty$ at each iteration.
  3. (iii)
    At the tail end of the iterations, we show this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{J}$ on the subspace $\mathcal{M}_{\mathcal{J}}$ and decrease the rest components to 0 at a linear speed. We then apply the center stable manifold theorem and prove that the algorithm can avoid converge to strict saddle points.
1.2 Outline
[…]
1.3 Notation
[…]
${\rm sign}(t)=\left\{\begin{aligned} &1,\ &t>0,\\ &-1,\ &t<0,\\ &0,\ &t=0.\end{aligned}\right.$ (1)
$DT(x)$ denote the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$. Given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}\subset\{1,\ldots,n\}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants. If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$. Given an index set $\mathcal{I}$, denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively. Given a symmetric matrix $H\in\mathbb{R}^{n\times n}$, let $\lambda_{\min}(H)$ be the smallest eigenvalue of $H$. Let $x\in\mathbb{R}^{n}$ and $w\in\mathbb{R}_{+}^{n}$, the soft-thresholding operator $\mathbb{S}_{w}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with the vector of thresholds $w$ is defined $[\mathbb{S}_{w}(x)]_{i}=\text{sign}(x_{i})\max(|x_{i}|-w_{i},0)$. In particular, if $w_{i}=\infty$, it follows that $[\mathbb{S}_{w}(x)]_{i}=0$.
Regularizer $\psi(x)$ $r^{\prime}(|x_{i}|)$
EXP [34] $\sum\limits_{i=1}^{n}(1-e^{-p|x_{i}|})$ $pe^{-p|x_{i}|}$
LOG [35] $\sum\limits_{i=1}^{n}\log(1+p|x_{i}|)$ $\frac{p}{1+p|x_{i}|}$
FRA [36] $\sum\limits_{i=1}^{n}\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{p}{(|x_{i}|+p)^{2}}$
LPN [7] $\sum\limits_{i=1}^{n}|x_{i}|^{p}$ $p|x_{i}|^{p-1}$
SCAD [5] $r(|x_{i}|)=\lambda\left\{\begin{array}[]{ll}|x_{i}|&{\rm if}\ |x_{i}|\leq\lambda\\ \frac{a\lambda-|x_{i}|}{a-1}&{\rm if}\ \lambda<|x_{i}|\leq a\lambda\\ \frac{(a+1)\lambda^{2}}{2}&{\rm if}\ |x_{i}|>a\lambda\end{array}\right.$ $\lambda\{\mathbb{I}(|x_{i}|\leq\lambda)+\frac{(a\lambda-|x_{i}|)_{+}}{(a-1)\lambda}\mathbb{I}(|x_{i}|>\lambda)\}$
MCP [6] $r(|x_{i}|)=\left\{\begin{array}[]{ll}\lambda|x_{i}|-\frac{\beta^{2}}{2a}&{\rm if}\ |x_{i}|\leq a\lambda\\ \frac{a\lambda^{2}}{2}&{\rm otherwise}.\end{array}\right.$ $(a\lambda-|x_{i}|)_{+}/a$
Table 1: Concrete examples of nonconvex regularization
2 Preliminaries
2.1 Optimality condition
In this subsection, we introduce the first and the second-order optimality condition of problem (P). First, let us recall the definition of subgradient and the first-order optimality condition.
Definition 2.
[…]
  1. (1)
    $v$ is a regular subgradient of $f$ at $\bar{x}$, written $v\in\hat{\partial}f(\bar{x})$ if
    $\partial f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$ (2)
  2. (2)
    $v$ is a (general) subgradient of $f$ at $\bar{x}$, written $v\in\partial f(\bar{x})$, if there are sequences $x^{v}\rightarrow\bar{x}$ and $v^{v}\in\hat{\partial}f(x^{v})$ with $x^{v}\rightarrow v$.
Let $s(\cdot)=r\circ|\cdot|$. It is trivial to see $s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$.
The subgradients of $s$ are given in the following proposition.
Proposition 1.
It holds that $\partial F(x)=\hat{\partial}F(x)=\nabla f(x)+\lambda\partial\psi(x)$ where
$\partial\psi(x)=\hat{\partial}\psi(x)=\partial\sum\limits_{i=1}^{n}s(x_{i})=\partial s(x_{1})\times...\times\partial s(x_{n})$ (3)
and
$\partial s(x_{i})=\hat{\partial}s(x_{i})=\begin{cases}\{\text{sign}(x_{i})r^{\prime}(|x_{i}|)\},&i\in\mathcal{I}(x),\\ [-r^{\prime}(|0^{+}|),+r^{\prime}(|0^{+}|)],&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)<\infty,\\ \mathbb{R},&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)=\infty.\end{cases}$ (4)
Proof.
We only need calculate the subgradients of $s$ at 00, i.e., $i\in\mathcal{J}(x)$. If $r^{\prime}(|0^{+}|)<\infty$, define for any $x\in\mathbb{R}$,
$\hat{s}(x):=\begin{cases}r(|x|)&\text{ if }x\geq 0,\\ r^{\prime}(|0^{+}|)x&\text{ otherwise}.\end{cases}$
Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$. By [37, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$. On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, meaning $r^{\prime}(|0^{+}|),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$. It follows from [37, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$. Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
If $r^{\prime}(|0^{+}|)=\infty$, we have $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{r(|x^{*}|)-r(0)-v(x-0)}{|x-0|}\geq 0$ for all $v\in\mathbb{R}$. Therefore, we have ^$\mathbb{R}\subseteq\hat{\partial}s(0)\subseteq\partial s(0)\subseteq\mathbb{R}$. Hence, ^$\hat{\partial}s(0)=\partial s(0)=\mathbb{R}$.
Overall, we have shown that (4) is true. It then follows from [37, Proposition 10.5] that (3) is true. ∎
We can appeal to the classical rule of Fermat [37, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a critical point of (P), then
$0\in\partial F(x^{*})=\nabla f(x^{*})+\lambda\partial\psi(x^{*})$ (5)
This condition, combined with Assumption 2 and the explicit form of the subgradients, yields the following results.
Proposition 2.
If $x^{*}$ is a local minimum for (P), then the following holds
$\nabla_{i}f(x^{*})=-\text{sign}(x_{i}^{*})\lambda r^{\prime}(|x^{*}_{i}|),$ $i\in\mathcal{I}^{*},$ (6)
$\nabla_{i}f(x^{*})\in(-\lambda r^{\prime}(|0^{+}|),+\lambda r^{\prime}(|0^{+}|)),$ $i\in\mathcal{J}^{*}.$ (7)
Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point satisfy the necessary optimality conditions (6) and (7) are referred as the critical point.
In the following, we demonstrate that the second-order optimality condition can be characterized by the nonzero components.
Theorem 3.
  1. (i)
    (necessary) If $x^{*}$ is a local minimum of $F$, then
    $d^{T}\nabla^{2}f(x^{*})d-\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (8)
    or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))\geq 0$.
  2. (ii)
    (sufficient) A critical point $x^{*}$ of $F$ is a strict local minimum if
    $d^{T}\nabla^{2}f(x^{*})d-\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (9)
    or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>0$.
Proof.
(i) is proved in [38, Theorem 3.3]. We only prove (ii).
Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of non-zero $\{u^{k}\}$ satisfying
$F(x^{*}+u^{k})\leq F(x^{*})\quad{\rm and}\quad\lim\limits_{k\rightarrow\infty}u^{k}=0.$ (10)
Consider $F_{|\mathcal{I}^{*}}$, which is locally smooth around $x^{*}_{\mathcal{I}^{*}}$ and $x^{*}_{\mathcal{I}^{*}}$ is stationary for $F_{|\mathcal{I}^{*}}$. It follows from $u^{k}=u_{\mathcal{I}^{*}}^{k}+u_{\mathcal{J}^{*}}^{k}$ that
$F(x^{*}+u^{k})-F(x^{*})$
$=$ $\ f(x^{*}+u^{k})+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
$=$ $\ f(x^{*}+u^{k})+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)+f(x^{*}+u_{\mathcal{I}^{*}}^{k})$
$-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
$=$ $\ f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}^{k}r(|u_{i}^{k}|).$
It is easy to know that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$ since $x^{*}$ is the critical point of (P) and (8) holds. Therefore for sufficiently large $k$,
$F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})>0.$ (11)
Moreover, by the Lagrangian mean value theorem, there exists sufficiently small $\epsilon^{k}\in(0,\min_{i}\{|u_{i}^{k}|\})$ such that
$r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$
The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}\|_{\infty})$; consequently,
$\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{1}\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{2}.$
This, combined with (11), yields
$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|,$ (12)
where $\eta^{k}:=\frac{f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})}{\|u_{\mathcal{J}^{*}}^{k}\|}$. By the definition of gradient,
$\lim_{u^{k}\to 0}\frac{f(x^{*}+u_{\mathcal{J}^{*}}^{k})-f(x^{*})-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\lim_{u^{k}\to 0}\eta^{k}-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=0,$
implying $\limsup\limits_{k\to\infty}\eta^{k}=\limsup\limits_{k\to\infty}\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\|\nabla f(x^{*})\|_{\infty}$. By (7), $\limsup\limits_{k\to\infty}\eta^{k}=\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$. Therefore, $\eta^{k}<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$. It follows from (12) that for sufficiently large $k$,
$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|>0,$
contradicting (10). ∎
2.2 Strict Saddle Point
In this section, we define the strict saddle point of the problem (P). In smooth settings, the strict saddle can be defined as the points with negative minimum eigenvalue of the Hessian matrix, which indicates a descent direction. For our case, Theorem 3 implies a saddle point can be defined using the nonzeros.
Definition 3.
A critical point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathcal{M}_{\mathcal{I}^{*}}\setminus\{0\}$. If all critical points of $F(x)$ are either strict saddle points or local minimums, $F$ satisfies the strict saddle property.
This definition is related to the “active strict saddle point” proposed in [23], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points. To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the critical point on which the function is partly smooth. Our definition also belongs to this kind, as $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold containing $x^{*}$.
2.3 Center Stable Manifold Theorem
As many other works in smooth setting like [40], we interpret our algorithm as the fixed point iteration of a nonlinear map $T:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$. Likewise, we define the fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as the unstable fixed points, as described below.
Definition 4.
(Unstable fixed point) Let $x^{*},\epsilon^{*}$ be a fixed point of a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. $x^{*},\epsilon^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*},\epsilon^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
The Center Stable Manifold (CSM) Theorem plays a crucial role in deriving the avoidance of strict saddle points. The CSM theorem establishes a favorable property for the fixed point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, which converge to the fixed point with at least one eigenvalue of $DT$ larger than $1$ in magnitude, has Lebesgue measure zero.
Theorem 4.
(The Center Stable Manifold Theorem) Let 00 be a fixed point for the $C^{1}$ local diffeomorphism $T:U\rightarrow\mathbb{R}^{n}$ where $U$ is a neighborhood of zero in $\mathbb{R}^{n}$. Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant splitting of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one, and greater than one. There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$, and $\cap_{k=0}^{\infty}T^{-k}(B)\subset W_{loc}^{cs}$.
The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero. Such result was established in [40, Theorem 2] for a global diffeomorphism map, under which the preimage of a measure zero set also has measure zero. The following corollary in [23] shows the same conclusion for lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous).
Corollary 1.
[23, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $\nabla T(x)$ is invertible. Then, the set of initial points attracted by such fixed points
$W:=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$
has zero Lebesgue measure.
3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
The ${\rm IRL}_{1}$ algorithm is an effective method to solve problem (P). At each iteration, the ${\rm IRL}_{1}$ algorithm solves the following weighted $\ell_{1}$ regularized subproblem,
$x^{k+1}=S(x^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i})|y_{i}|,$
where $w(x_{i})=r^{\prime}(|x_{i}|)$. To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a critical point and a Lipeomorphism globally. However, this is not true in our case since $S$ is generally non-invertible. To see this, one can rewrite the map $S$ using the soft-threshold operator with a changing threshold as
$S(x^{k})=\mathbb{S}_{\lambda w^{k}}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})).$
This algorithm may quickly converges to a local minimizers due to the rapid growth of the weight values, especially for the case when $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered. To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
$\min\limits_{\theta\in\mathbb{R}^{2n}}F(\theta):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}^{2}).$ (13)
with $\theta:=[x;\epsilon]$ for shorthand. Here we also treat $\epsilon$ as another variable to optimize, so that the algorithm updates both $x$ and $\epsilon$ by solving the following subproblem
$x^{k+1}=S(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w_{i}^{k}|y_{i}|,$
with $w_{i}^{k}=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and setting $\epsilon^{k+1}=\mu\epsilon^{k}$ with $\mu\in(0,1)$.
The subproblem solution can be given explicitly using the solft-thresholding operator. This update can be written as
$S(x^{k})=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$ (14)
or explicitly,
$S_{i}(x^{k})=\left\{\begin{aligned} &x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{ if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ &x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{ if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ &0,&\text{otherwise }.\end{aligned}\right.$ (15)
The map $S$ is not a bijection, as it maps multiple values to 00. To address this issue, we need to revise this algorithm to ensure it satisfy these conditions. Motivated by the technique used in [23], we modified the algorithm with following damping iteration,
$x^{k+1}=(1-\alpha)x^{k}+\alpha S(x^{k},\epsilon^{k}),$
where $\alpha\in(0,1)$ is the damping parameter. Intuitively, if $S_{i}(x^{k},\epsilon^{k}_{i})=0$, the damping operator maps $x^{k}_{i}$ to $(1-\alpha)x^{k}_{i}$, resulting in a bijective map. We can use the iterate $(x;\epsilon)\in\mathbb{R}^{2n}$ and equivalently write the update of $(x^{k};\epsilon^{k})$ as
$(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S(x^{k},\epsilon^{k})\\ \mu\epsilon^{k}\end{bmatrix}$ (16)
with $\epsilon_{0}>0$, $\mu\in(0,1)$ and $\alpha\in(0,1)$.
This algorithm, hereinafter nicknamed DIRL1, is presented in Algorithm 10. If the algorithm does not involve the vector of $\epsilon$, we can simply treat this case as setting $\epsilon^{0}=0$ in Algorithm 10 and the iteration simply reverts to
$x^{k+1}=T(x^{k})=(1-\alpha)x^{k}+\alpha S(x^{k}).$
We have the following result to connect the optimality of (P) and the fixed point of our iteration.
Proposition 5.
$x^{*}$ is a critical point of (P) if and only if $x^{*}=S(x^{*},0)$. In addition, $x^{*}$ is a critical point of (P) if and only if $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$.
Proof.
Suppose $x^{*}$ is a critical point. It follows from (6) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}(x^{*},0)=x_{i}^{*}$. This result also holds true by the same argument if $x_{i}^{*}<0$.
As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (7). Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$. Therefore, $S_{i}(x^{*},0)=0=x_{i}^{*}$. Overall, we have shown that $x^{*}=S(x^{*},0)$ in all cases.
Conversely, suppose $x^{*}=S(x^{*},0)$. If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$. Same argument yields $0=x_{i}^{*}-S_{i}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$. Therefore, $\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)=0$ for $i\in\mathcal{I}^{*}$.
If $x_{i}^{*}=0$, it indicates that $|\frac{1}{\beta}\nabla_{i}f(x^{*})|=|x_{i}^{*}-\frac{1}{\beta}\nabla_{i}f(x^{*})|\leq\tfrac{\lambda}{\beta}w(0,0)=\tfrac{\lambda}{\beta}r^{\prime}(|0^{+}|)$. Hence, $\nabla_{i}f(x^{*})\in[-\lambda r^{\prime}(|0^{+}|),-\lambda r^{\prime}(|0^{+}|)]$. Overall, $x^{*}$ satisfies (5) and is a critical point of (P). By Assumption 2, it satisfies (6) and (7).
The rest part of the statement of this proposition is obviously true by noticing that $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$ implies that $\epsilon^{*}=0$. ∎
1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
2:Initialize: set $k=0$.
3:repeat
4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})$.
5:     Compute new iterate:
6:     $y^{k}=\arg\min\limits_{y\in\mathbb{R}^{n}}\{\nabla f(x^{k})^{T}(y-x^{k})+\frac{\beta}{2}\|y-x^{k}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})|y_{i}|\}$.
7:     $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$.
8:     $\epsilon^{k+1}=\mu\epsilon^{k}$.
9:     Set $k=k+1$.
10:until Convergence
Algorithm 1 Damped Iteratively Reweighted $\ell_{1}$ Algorithm (${\rm DIRL}_{1}$)
4 Convergence analysis
The analysis focuses solely on the algorithm involving the update of $\epsilon$. It is noteworthy the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.
Assumption 3.
Given the initial point $(x^{0}$, $\epsilon^{0})$ and $F^{0}=F(x^{0},\epsilon^{0})$, the level set $\mathcal{L}(F^{0}):=\{(x,\epsilon)\mid F(x,\epsilon)\leq F^{0}\}$ is bounded.
This assumption means $f$ is Lipschitz continuous and Lipschitz differentiable on $\mathcal{L}(F^{0})$ with constant $L_{f}$, i.e., $\forall x,y\in\mathbb{R}^{n}$ (we use the same constant for $f$ and $\nabla f$ for simplicity).
$\|f(x)-f(y)\|\leq L_{f}\|x-y\|\text{ \ and \ }\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|.$ (17)
We first show the monotonicity of $F(x,\epsilon)$ along a direction. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
$\Delta F(x^{k+1},\epsilon^{k+1})=$ $\ F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k+1}),$
$\Delta G(x^{k+1};x^{k},\epsilon^{k})=$ $\ G(x^{k};x^{k},\epsilon^{k})-G(x^{k+1};x^{k},\epsilon^{k}).$
Proposition 6.
Suppose Assumption 1-3 hold. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$. It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and the reduction satisfies
$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
Moreover, $\lim\limits_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$ and there exists $C>0$ such that
$\|x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})\|_{\infty}\leq C,$ (18)
for any $k\in\mathbb{N}$.
Proof.
Since the concavity of $r(\cdot)$ on $\mathbb{R}_{++}$ gives $r(x_{1})\leq r(x_{2})+r^{\prime}(x_{2})(x_{1}-x_{2})$ for any $x_{1},x_{2}\in\mathbb{R}_{++}$, we can imply that for $i=1,...,n$,
$r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})+r^{\prime}(|x_{i}^{k}|+(\epsilon^{k}_{i})^{2})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
Summing up the above inequality over $i$ yields
$\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$ (19)
On the other hand, it follows that
$f(x^{k})-f(x^{k+1})\geq\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}.$ (20)
Combining (19) and (20), we have
$\Delta F(x^{k+1},\epsilon^{k+1})\geq F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k})$ (21)
$\geq$ $\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)$
$=$ $\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2},$
where the first inequality is by the fact that $\epsilon^{k+1}\leq\epsilon^{k}$.
For each subproblem, $y^{k}$ satisfies the optimal condition
$\nabla f(x^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})=0.$ (22)
with $\xi^{k}\in\partial\|y^{k}\|_{1}$ and $W(x^{k},\epsilon^{k}):={\rm diag}(w(x_{1}^{k},\epsilon_{1}^{k}),...,w(x_{n}^{k},\epsilon_{n}^{k}))$. Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimal condition (22), the following holds
$\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$ (23)
$=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
$\geq$ $\ \nabla f(x^{k})^{T}(\alpha x^{k}-\alpha y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(\alpha|x_{i}^{k}|-\alpha|y_{i}^{k}|)-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$\geq$ $\ \alpha[\nabla f(x^{k})^{T}(x^{k}-y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})\xi_{i}^{k}(x_{i}^{k}-y_{i}^{k})]-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ \alpha[\nabla f(y^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})]^{T}(x^{k}-y^{k})+\frac{\alpha\beta(2-\alpha)}{2}\|x^{k}-y^{k}\|^{2}$
$=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2},$
where the first inequality is by $\alpha\in(0,1)$ and the second is by the convexity of ||$\cdot$|| ⋅ |. Combine (21), (23) and get
$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
Replacing $k$ with $t$ and summing up from $t=0$ to $k-1$, we have
$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
It follows that $\{(x^{k},\epsilon^{k})\}\in\mathcal{L}(F^{0})$ since $F(x^{k},\epsilon^{k})\leq F(x^{0},\epsilon^{0})$. By Assumption 3, it is bounded and $\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})>-\infty$. Taking $k\rightarrow\infty$, we have
$(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, it follows from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (18) holds for any $k\in\mathbb{N}$. ∎
With the monotonicity of $F$, we derive the global convergence of the ${\rm DIRL}_{1}$ method.
Theorem 7.
(Global convergence) Suppose Assumption 1 and 3 hold. Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$. Then $\mu^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
Proof.
Firstly, boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$. Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$. The optimality condition of the $k$th subproblem implies
$-\beta(y_{i}^{k}-x_{i}^{k})\in\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k})\partial|y_{i}^{k}|.$
Then we consider the two kinds of cases with $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$.
If $r^{\prime}(|0^{+}|)=\infty$, as shown in Proposition 2, we only need to consider $i\in\mathcal{I}^{*}$. Taking the limit on $\mathcal{S}$,
00 0 $=\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(y_{i}^{k})+\beta(y_{i}^{k}-x_{i}^{k})|$
$=\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(x_{i}^{k})+\frac{\beta}{\alpha}(x_{i}^{k+1}-x_{i}^{k})|$
$\geq\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(x_{i}^{k})|-\frac{\beta}{\alpha}\|x^{k}-x^{k+1}\|$
$=|\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0){\rm sign}(x_{i}^{*})|$
$=|\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|){\rm sign}(x_{i}^{*})|.$
which satisfy (5). Therefore, $x^{*}$ is stationary.
If $r^{\prime}(|0^{+}|)<\infty$, we need also to consider $i\in\mathcal{J}^{*}$. By [41, Proposition 2.1.5] and [37, Theorem 5.19], $\partial r$ is outer semi-continuous and there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that
$0=\nabla_{i}f(x^{*})+\lambda w(0,0)\xi^{*}_{i}=\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)\text{sign}(x_{i}^{*}),$
meaning (5) is satisfied. Therefore, $x^{*}$ is stationary. ∎
5 Avoidance of strict saddle point
DIRL1 has the property of so-called “active manifold identification”, which belongs to the research line of active-manifold identification in nonsmooth optimization [42, 43]. It means the iterates fall onto a smooth manifold after finite iterations. The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consists of the nonzeros at a limit point.
Theorem 8.
(Model Identification and Local Smoothness) Under Assumption 1-3, for any critical point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
  1. (i)
    For any point $(x,\epsilon)\in B$, $S(x,\epsilon)\in(x^{*}+\mathcal{M}_{\mathcal{I}^{*}})$ and $\text{sign}(S(x,\epsilon))=\text{sign}(x)=\text{sign}(x^{*})$.
  2. (ii)
    Map $S$ is $C^{1}$-smooth in $B$ and $S(x)=x-\frac{1}{\beta}g(\theta)$ with
    $g_{i}(\theta)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
  3. (iii)
    Map $T$ is $C^{1}$-smooth in $B$.
Proof.
(i) Since $S$ is a continuous map and $x^{*}=S(x^{*},\epsilon^{*})$, it holds that
$\text{sign}(x_{i})=\text{sign}(x^{*}_{i})\text{\ and\ }\text{sign}(S_{i}(x,\epsilon))=\text{sign}(S_{i}(x^{*},0))=\text{sign}(x^{*}_{i}),\quad i\in\mathcal{I}^{*}$
for any $(x,\epsilon)$ sufficiently close to $(x^{*},\epsilon^{*})$. It remains to show that for any $i\in\mathcal{J}^{*}$, $S_{i}(x,\epsilon)=0$ for any $(x,\epsilon)\in B$. By Assumption 2, $|x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})|=|\tfrac{1}{\beta}\nabla_{i}f(x^{*})|<\tfrac{1}{\beta}\lambda w(0,0)$. Therefore, there exists a neighborhood $B$ such that $|x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)|<\tfrac{1}{\beta}\lambda w(x_{i},\epsilon_{i})$, $\forall(x,\epsilon)\in B$, which is equivalent to $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))<0$ and $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))>0$, meaning $S_{i}(x,\epsilon)=0$.
(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agree with following constrain problem
$S(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (24)
$=$ $\ \arg\min\limits_{x\in\mathbb{R}^{n}}\{g(x^{k},\epsilon^{k})^{T}x+\frac{\beta}{2}\|x-x^{k}\|^{2}\},$
where each component of $g:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ is defined by
$g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
Therefore, $S$ is $C^{1}$-smooth in $B$
(iii) It follows from (24) that the iteration $T$ is $C^{1}$-smooth in $B$ for any for any $(x,\epsilon)\in B$. ∎
The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent for the variables in $\mathcal{I}^{*}$ locally around a critical point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$. Specifically, it can be expressed as
$T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &\mu\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$ (25)
around $(x^{*},\epsilon^{*})$. We then proceed to the saddle escape mechanism for DIRL1, which ensures that the algorithm does not converge to the strict saddle points defined in Definition 3 and, instead, converges to local minima. To apply the CSM theorem, we have to show the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
Theorem 9.
(Unstable Fixed Point) Suppose $x^{*}$ is a critical point of (P). $x^{*}$ is a strict saddle point of (P) if and only if the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$.
Proof.
First of all, we calculate the Jabobian of $T$ in the neighborhood $B$ described in Theorem 8. It holds that
$\frac{\partial g}{\partial\epsilon_{i}}(x,\epsilon)=\begin{cases}0&\text{ \ if}i\in\mathcal{J}^{*},\\ \lambda r^{\prime\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*})2\epsilon&\text{ \ if }i\in\mathcal{I}^{*}.\end{cases}$
Let $D_{\epsilon}g(x,\epsilon)$ be the Jacobian matrix of $g(\cdot)$ with respect to $\epsilon$. It follows that $D_{\epsilon}g(x^{*},\epsilon^{*})=\bm{0}$ by the fact that $\epsilon_{i}^{*}=0$ and $r^{\prime\prime}(|x_{i}^{*}|)<+\infty$, $i\in\mathcal{I}^{*}$. On the other hand, $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \bm{0}&\beta I\end{bmatrix}$. Therefore, the Jacobian of $T$ at $(x^{*},\epsilon^{*})$ is given by
$DT^{*}(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}\ &\mu I\end{bmatrix}=\begin{bmatrix}I-\frac{\alpha}{\beta}D_{x}g(x^{*},\epsilon^{*})\ &\bm{0}\\ \bm{0}\ &\mu I\end{bmatrix}$
$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&\bm{0}\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&\mu I\end{bmatrix}$
Now we are ready to prove the equivalence between the unstable fixed point of $T$ and the strict saddle point of (P). Recall Definition 4 and Definition 3 and notice $\mu\in(0,1)$ and $\alpha\in(0,1)$. It suffices to show that $I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has an eigenvalue greater than 1 if and only if $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has a negative eigenvalue. This equivalence is obviously true since
$\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0\iff\lambda_{\max}(I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>1.$
To apply Corollary 1, we next to show $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$. This is equivalent to show $T$ and its inverse $T^{-1}$ is Lipschitz continuous.
Proposition 10.
Suppose Assumption 1-3 hold. Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ defined in Proposition 6, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$. $T$ is a lipeomorphism when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
Proof.
If $r^{\prime}(|0+|)<\infty$, $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|,\epsilon_{i}^{2})$ is Lipschitz continuous in $\mathcal{L}(F^{0})$.
If $r^{\prime}(|0+|)=\infty$, by Proposition 6, $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$. When $r^{\prime}(|x_{i}|)\geq w(x_{i},\epsilon_{i})>C$, then it must be true that $S_{i}(x,\epsilon)=0$. Therefore, for any $S_{i}(x)=0$ with $x\in\mathcal{L}(F^{0})$,
$r^{\prime}(|x_{i}|)\leq C\text{ \ and \ }|x_{i}|\geq\underline{x}:=(r^{\prime})^{-1}(C)>0$ (26)
by Assumption 3.
This implies that we can equivalently write $S$ as
$S_{i}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
with
$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i}^{2})&\text{ if }r^{\prime}(|0+|)<\infty\\ \min(w(\underline{x},0),w(x_{i},\epsilon_{i}))=\min(r^{\prime}(\underline{x}),r^{\prime}(|x_{i}|+\epsilon_{i}^{2}))&\text{ if }r^{\prime}(|0+|)=\infty.\end{cases}$
Therefore, $u(\cdot)=\min(r^{\prime}(\underline{x}),r^{\prime}(\cdot))$ is also Lipschitz continuous in $\mathcal{L}(F^{0})$. Let $L_{r}$ be the Lipschitz constant, meaning for any $\hat{x},\tilde{x}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$
$\|u(\hat{x},\hat{\epsilon})-u(\tilde{x},\tilde{\epsilon})\|\leq L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|.$ (27)
A natural choice is $L_{r}=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$ and $L_{r}=r^{\prime\prime}(|\underline{x}|)$ if $r^{\prime}(|0^{+}|)=\infty$.
Next we prove $S$ is Lipschitz continuous. It follows that
$S(\hat{x},\hat{\epsilon})-S(\tilde{x},\tilde{\epsilon})=$ $\ \mathbb{S}_{\lambda\hat{u}/\beta}(\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x}))-\mathbb{S}_{\lambda\tilde{u}/\beta}(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))$
$\leq$ $\ \|\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x})-(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))\|+\|\hat{u}-\tilde{u}\|$
$\leq$ $\ \|\hat{x}-\tilde{x}\|+\tfrac{1}{\beta}\|\nabla f(\hat{x})-\nabla f(\tilde{x}))\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
$\leq$ $\ (1+\tfrac{1}{\beta}L_{f})\|\hat{x}-\tilde{x}\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
$\leq$ $\ (1+\tfrac{1}{\beta}L_{f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|,$
where the first inequality is by Lemma A.2 and the second is by (17) and (27). Therefore, $\|\alpha(\hat{x}-S(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
Notice that $T=\begin{bmatrix}x\\ \epsilon\end{bmatrix}+\alpha(\begin{bmatrix}S(x,\epsilon)\\ \mu\epsilon\end{bmatrix}-\begin{bmatrix}x\\ \epsilon\end{bmatrix})$, Since $\begin{bmatrix}S(x,\epsilon)\\ \mu\epsilon\end{bmatrix}$ is Lipschitz with constant $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$. Applying [23, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is a lipeomorphism. ∎
Now we are ready to apply Corollary 1 to derive the following global escape theorem.
Theorem 11.
(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3. Let the condition of Proposition 10 holds. Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10. Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure. of (P).
Proof.
It follows directly from Theorem 9 that map $T$ is local $C^{1}$-smooth around any critical points, and all critical points $x^{*}$ are unstable fixed points. By applying Proposition 10, $T$ is a lipeomorphism. By Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure. ∎
Appendix
Lemma A.1.
Given $a,b,c\in\mathbb{R}^{n}_{+}$ with $a_{i}\leq b_{i}+c_{i}\quad\text{for all }i$. It holds that $\|a\|\leq\|b\|+\|c\|$.
Proof.
It follows that $a_{i}^{2}\leq b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2}$, implying $\sum_{i=1}^{n}a_{i}^{2}\leq\sum_{i=1}^{n}(b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2})$, or equivalently, $\|a\|^{2}\leq\|b\|^{2}+2b^{T}c+\|c\|^{2}$. By the Cauchy-Schwarz inequality, $2\sum_{i=1}^{n}b_{i}c_{i}\leq 2\|b\|\|c\|$. Therefore, $\|a\|^{2}\leq(\|b\|+\|c\|)^{2}$, completing the proof. ∎
Lemma A.2.
Given $\hat{w},\tilde{w}\in\mathbb{R}^{n}_{+}$ and $\hat{z},\tilde{z}\in\mathbb{R}^{n}$, it holds that
$\|\mathbb{S}_{\hat{w}}(\hat{z})-\mathbb{S}_{\tilde{w}}(\tilde{z})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$
Proof.
(a) If $|\hat{z}_{i}|\leq\hat{w}_{i}$ and $|\tilde{y}_{i}|\leq\tilde{w}_{i}$, then $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})=0$ and $\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=0$. In this case, $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=|0-0|=0\leq|\hat{z}_{i}-\tilde{z}_{i}|$.
(b) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|\leq\tilde{w}_{i}$ (or vice versa), then $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})=\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})$ and $\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=0$. In this case, $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=||\hat{z}_{i}|-\hat{w}_{i}|$. It follows that
$|\hat{z}_{i}-\tilde{z}_{i}|\geq$ $\ |\hat{z}_{i}|-|\tilde{z}_{i}|\geq|\hat{z}_{i}|-\tilde{w}_{i}$
$=$ $\ ||\hat{z}_{i}|-\hat{w}_{i}+\lambda(\hat{w}_{i}-\tilde{w}_{i})/\beta|$
$\geq$ $\ |\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|-|\lambda(\hat{w}_{i}-\tilde{w}_{i})/\beta|,$
implying
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
(c) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|>\tilde{w}_{i}$, then
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})-\text{sign}(\tilde{z}_{i})(|\tilde{z}_{i}|-\tilde{w}_{i})|.$
If $\text{sign}(\hat{z}_{i})=\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|(|\hat{z}_{i}|-\hat{w}_{i})-(|\tilde{z}_{i}|-\tilde{w}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|$. If $\text{sign}(\hat{z}_{i})=-\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=(|\hat{z}_{i}|-\hat{w}_{i})+(|\tilde{z}_{i}|-\tilde{w}_{i})\leq|\hat{z}_{i}|+|\tilde{z}_{i}|=|\hat{z}_{i}-\tilde{z}_{i}|.$
In all three cases, we show
$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
Therefore, by (A.1),
$\|\mathbb{S}_{\hat{w}}(\hat{w})-\mathbb{S}_{\tilde{w}}(\tilde{w})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$

### v3

@@
-Avoiding strict saddle points of nonconvex regularized problems
+
 Abstract
 In this paper, we consider a class of non-convex and non-smooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms.
-These problems, moreover, are probably non-lipschitz around sparse solutions.
-We propose an damped iterative reweighted $\ell_{1}$ algorithm to solve these problems.
-The algorithm is guaranteed to converge only to local minimizers when randomly initialized.
-By deriving a second-order sufficient optimal condition under the non-degenerate assumption, we show that the strict saddle property is generic on these sparse optimization problems.
+We show the second-order optimality conditions only depend on the nonzeros of the stationary points.
+We propose two damped iterative reweighted algorithms including the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems.
+For DIRL1, we show the reweighted $\ell_{1}$ subproblem has support identification property so that DIRL1 locally reverts to a gradient descent algorithm around a stationary point.
+For DIRL2, we show the solution map of the reweighted $\ell_{2}$ subproblem is differentiable and Lipschitz continuous everywhere.
+Therefore, the map of DIRL1 and DIRL2 and their inverse are Lipschitz continuous, and the strict saddle points are their unstable fixed points.
+By applying the stable manifold theorem, these algorithms are shown to converge only to local minimizers with randomly initialization when the strictly saddle point property is assumed.
 keywords:
@@
 where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and the penalty/regularization term $\psi(x)$ takes the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$.
-We assume the objective $F$ has bounded level sets, the data fitting function $f$ is twice continuously differentiable and possibly nonconvex, and $r$ satisfies the following properties.
+The data fitting function $f$ is twice continuously differentiable and possibly nonconvex, and $r$ satisfies the following properties.
 Assumption 1.
-$r$ is continuous on $[0,+\infty)$ with $r(0)=0$, $C^{2}$-smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
+$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$, C2-smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
 This assumption indicates $r(t)>0$ for all $t>0$, but $r$ could be nonconvex or not Lipschitz around 0.
@@
 In smooth settings, it is widely assumed that the strict saddle property holds, meaning all saddle points possess at least one strictly negative eigenvalue.
-This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases and has been shown to be generally satisfied for smooth problems [20, 21].
+This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases and it has been shown to be generally satisfied for smooth problems [20, 21].
 In nonsmooth optimization, avoiding saddle points poses a greater challenge.
-At stationary point, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate the avoidance of saddle points.
+At stationary points, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate the avoidance of saddle points.
 Developing nonsmooth optimization algorithms capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than in smooth optimization.
@@
 They assume an active-manifold around the limit point and then define the strict saddle points as those of the $f$ restricted to the manifold.
-The weakly $\rho$-convexity for $r$ plays a crucial role in constructing C2 differentiable Moreau envelope with the same critical points and strict saddle points in proximal methods.
+The weakly $\rho$-convexity for $r$ plays a crucial role in constructing C2 differentiable Moreau envelope with the same stationary points and strict saddle points in proximal methods.
 As for the regularized problem (P), several first-order methods have been proposed.
@@
 Furthermore, convergence rate analysis under the Kurdyka-Łojasiewicz property [32] and its extrapolated version was proposed in [33].
-While these first-order methods primarily focus on identifying first-order critical points, none of them guarantee avoidance of saddle points.
+While these first-order methods primarily focus on identifying first-order stationary points, none of them guarantee the avoidance of saddle points.
 1.1 The main results
-In this paper, our main purpose is to develop an iterative reweighted $\ell_{1}$ algorithm that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weakly convexity.
-The additional assumption for $r$ needed in our work to guarantee that the existence of an active manifold around an active manifold as assumed by [23].
+In this paper, our main purpose is to develop and analyze first-order methods that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weakly convexity.
+In particular, we focus on iteratively reweighted methods which approximate the regularization term by simpler weighted $\ell_{1}$ or $\ell_{2}$ norms [34, 35, 36, 30, 37, 29, 31, 33].
+In contrast to the proximal gradient algorithm, iteratively reweighted methods solve a sequence of weighted $\ell_{1}$ or $\ell_{2}$ regularized subproblems and are widely applicable to different non-convex regularization problems.
+The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator.
+This method offers a favorable local property [31], known as the model identification, under the following additional assumption.
 Assumption 2.
 For any stationary point $x^{*}$, $0\in{\rm rint}\partial F(x^{*})$ (where $\partial F$ denotes the regular subdifferential and ${\rm rint}$ denotes the relative interior).
-The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) is widely applicable to these non-convex regularization problems.
-In contrast to the proximal gradient algorithm, ${\rm IRL}_{1}$ solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator.
-This method offers a favorable local property [31], known as the model identification.
+This kind of assumption is also widely made to guarantee that the existence of an active manifold around a stationary point as assumed by [23].
 At the tail end of the iteration, the sign of the iterates remain unchanged while the nonzero components are kept uniformly bounded away from 0.
@@
 This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{J}}$.
-If we think the iteration of ${\rm IRL}_{1}$ as a mapping that assigns to every point another point.
-This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) on $\mathcal{M}_{\mathcal{J}}$, even though the underlying optimization problem is nonsmooth in the entire space.
-We can summarize our contribution as follows:
+This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) on $\mathcal{M}_{\mathcal{J}}$, even though the underlying optimization problem is nonsmooth in $\mathbb{R}^{n}$.
+The iterative reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these non-convex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the solution is trivially given.
+Compared with ${\rm IRL}_{1}$, ${\rm IRL}_{2}$ does not enjoy the model subspace identification/active-manifold identity property.
+In fact, it can be often witnessed that this algorithm can generate iterates with elements asymptotically approaching zero.
+To show the different behaviors of the two algorithms, we provide the following example.
+Example 1.
+Consider the $\ell_{1/2}^{1/2}$ regularization problem
+$\min_{x\in\mathbb{R}^{2}}F(x)=f(x)+\|x\|_{1/2}^{1/2},\ \text{ with }f(x)=x_{1}^{2}+(x_{2}-5/4)^{2}.$
+The global minimizer is $(0,1)$, the local minimizer111(0,0)00(0,0)( 0 , 0 ) is always a dummy local minimizer of $\ell_{p}$ regularization problems is $(0,0)$ and the saddle point minimizer is $(0,\frac{3-2\sqrt{2}}{4})\approx(0,0.043)$ which is also a local maximizer on the $x_{2}$-axis.
+(a) Contour of $F$.
+(b) Contour of $F$ around $(0,0)$.
+Let $g(x;x^{k})=\nabla f(x^{k})^{T}(x-x^{k})+\tfrac{1}{2}\|x-x^{k}\|^{2}$. The iterations of iteratively reweighted algorithms are as follows [29]. We use the ${\rm IRL}_{1}$ algorithm with iteration
+$x^{k+1}=\arg\min_{x_{1},x_{2}}g(x;x^{k})+\tfrac{1}{2}\frac{1}{(|x_{1}^{k}|+0.1^{k})^{1/2}}|x_{1}|+\tfrac{1}{2}\frac{1}{(|x_{2}^{k}|+0.1^{k})^{1/2}}|x_{2}|$
+and the ${\rm IRL}_{2}$ algorithm with iteration
+$x^{k+1}=\arg\min_{x_{1},x_{2}}g(x;x^{k})+\tfrac{1}{4}\frac{1}{((x_{1}^{k})^{2}+0.1^{k})^{3/4}}x_{1}^{2}+\tfrac{1}{4}\frac{1}{((x_{2}^{k})^{2}+0.1^{k})^{3/4}}x_{2}^{2}.$
+The trajectories of the two algorithms are depicted in Figure 2.
+(a) Trajectory of ${\rm IRL}_{1}$
+(b) Trajectory of ${\rm IRL}_{2}$
+Figure 2: Trajectories generated by iteratively reweighted methods. ${\rm IRL}_{1}$ eventually falls into the subspace (the $x_{2}$-axis), while ${\rm IRL}_{2}$ has nonzero elements for all iterates.
+The behavior of ${\rm IRL}_{2}$ brings another challenge of analyzing the avoidance of strict saddle point.
+Current analysis for nonsmooth optimization algorithms ubiquitously depends on the active-manifold identification property, which ensures the algorithm reverts to a smooth algorithm in the active manifold in a local neighbourhood of the stationary point.
+However, there is no answer how a nonsmooth first-order algorithm like ${\rm IRL}_{2}$ would behave in the local region of a strict saddle point.
+We now summarize our contribution as follows:
   1. (i)
@@
     We develop damped IRL1 algorithms individually for the two types of regularizers, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning the next iterate is the combination of a traditional IRL1 iterate and its previous iterate.
-The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ becomes $\infty$ at each iteration.
+The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ from becoming $\infty$ at each iteration.
+At the tail end of the iterations, we show this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{J}$ on the subspace $\mathcal{M}_{\mathcal{J}}$ and decrease the rest components to 0 at a linear speed.
+We then apply the center stable manifold theorem and prove that the algorithm can avoid the convergence to strict saddle points.
   3. (iii)
-    At the tail end of the iterations, we show this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{J}$ on the subspace $\mathcal{M}_{\mathcal{J}}$ and decrease the rest components to 0 at a linear speed.
-We then apply the center stable manifold theorem and prove that the algorithm can avoid converge to strict saddle points.
+    We develop damped IRL2 algorithms for the nonconvex regularized problems and provide the global convergence.
+Under an additional assumption which is satisfied by most common nonconvex regularizers, we show that the algorithm is $C^{1}$ smooth and nondegenerate everywhere.
+We also proved the equivalence between the unstable fixed point of the iteration map and the strict saddle points of the regularization problem.
+Therefore, the avoidance of strict saddle points are derived.
+This is the first algorithm shown to have this property without using the active-manifold identification property nor the stochastic technique in nonsmooth optimization algorithms.
 1.2 Outline
@@
 $DT(x)$ denote the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$.
-Given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}\subset\{1,\ldots,n\}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants.
+Given $x\in\mathbb{R}^{n}$ and $\mathcal{I}\subset\{1,\ldots,n\}$, denote $x_{\mathcal{I}^{*}}\subset\mathbb{R}^{\mathcal{I}^{*}}$ or $[x_{i}]_{i\in\mathcal{I}}$ as the subvector of the elements in $\mathcal{I}$; also, given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants.
 If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$.
-Given an index set $\mathcal{I}$, denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively.
+Denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively.
 Given a symmetric matrix $H\in\mathbb{R}^{n\times n}$, let $\lambda_{\min}(H)$ be the smallest eigenvalue of $H$.
@@
 In particular, if $w_{i}=\infty$, it follows that $[\mathbb{S}_{w}(x)]_{i}=0$.
-Regularizer $\psi(x)$ $r^{\prime}(|x_{i}|)$
-EXP [34] $\sum\limits_{i=1}^{n}(1-e^{-p|x_{i}|})$ $pe^{-p|x_{i}|}$
-LOG [35] $\sum\limits_{i=1}^{n}\log(1+p|x_{i}|)$ $\frac{p}{1+p|x_{i}|}$
-FRA [36] $\sum\limits_{i=1}^{n}\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{p}{(|x_{i}|+p)^{2}}$
-LPN [7] $\sum\limits_{i=1}^{n}|x_{i}|^{p}$ $p|x_{i}|^{p-1}$
-SCAD [5] $r(|x_{i}|)=\lambda\left\{\begin{array}[]{ll}|x_{i}|&{\rm if}\ |x_{i}|\leq\lambda\\ \frac{a\lambda-|x_{i}|}{a-1}&{\rm if}\ \lambda<|x_{i}|\leq a\lambda\\ \frac{(a+1)\lambda^{2}}{2}&{\rm if}\ |x_{i}|>a\lambda\end{array}\right.$ $\lambda\{\mathbb{I}(|x_{i}|\leq\lambda)+\frac{(a\lambda-|x_{i}|)_{+}}{(a-1)\lambda}\mathbb{I}(|x_{i}|>\lambda)\}$
-MCP [6] $r(|x_{i}|)=\left\{\begin{array}[]{ll}\lambda|x_{i}|-\frac{\beta^{2}}{2a}&{\rm if}\ |x_{i}|\leq a\lambda\\ \frac{a\lambda^{2}}{2}&{\rm otherwise}.\end{array}\right.$ $(a\lambda-|x_{i}|)_{+}/a$
+Regularizer $\psi(x)$ $r^{\prime}(|x_{i}|)$ $r^{\prime\prime}(|x_{i}|)$
+EXP [38] $\sum\limits_{i=1}^{n}(1-e^{-p|x_{i}|})$ $pe^{-p|x_{i}|}$ $-\frac{1}{p^{2}}e^{-\frac{|x_{i}|}{p}}{p}$
+LOG [39] $\sum\limits_{i=1}^{n}\log(1+p|x_{i}|)$ $\frac{p}{1+p|x_{i}|}$ $-\frac{1}{(|x_{i}|+p)^{2}}$
+FRA [40] $\sum\limits_{i=1}^{n}\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{p}{(|x_{i}|+p)^{2}}$ $-\frac{2p}{(|x_{i}|+p)^{2}}$
+LPN [7] $\sum\limits_{i=1}^{n}|x_{i}|^{p}$ $p|x_{i}|^{p-1}$ $p(p-1)|x_{i}|^{p-2}$
+TAN [41] $\sum\limits_{i=1}^{n}\arctan\frac{|x_{i}|}{p}$ $\frac{p}{(|x_{i}|^{2}+p^{2})^{2}}$ $-\frac{2p|x_{i}|}{(|x_{i}|^{2}+p^{2})^{2}}$
 Table 1: Concrete examples of nonconvex regularization
-2 Preliminaries
-2.1 Optimality condition
+2 Optimality and strict saddle points
+2.1 Optimality conditions
 In this subsection, we introduce the first and the second-order optimality condition of problem (P). First, let us recall the definition of subgradient and the first-order optimality condition.
@@
     $v$ is a regular subgradient of $f$ at $\bar{x}$, written $v\in\hat{\partial}f(\bar{x})$ if
-    $\partial f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$ (2)
+    ^$\hat{\partial}f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$ (2)
   2. (2)
     $v$ is a (general) subgradient of $f$ at $\bar{x}$, written $v\in\partial f(\bar{x})$, if there are sequences $x^{v}\rightarrow\bar{x}$ and $v^{v}\in\hat{\partial}f(x^{v})$ with $x^{v}\rightarrow v$.
-Let $s(\cdot)=r\circ|\cdot|$. It is trivial to see $s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$.
-The subgradients of $s$ are given in the following proposition.
+It is trivial to see $s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$. The subgradients of $s$ are given in the following proposition.
 Proposition 2.
@@
 Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$.
-By [37, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
+By [42, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
 On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, meaning $r^{\prime}(|0^{+}|),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$.
-It follows from [37, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$.
+It follows from [42, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$.
 Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
@@
 Hence, ^$\hat{\partial}s(0)=\partial s(0)=\mathbb{R}$.
-Overall, we have shown that (4) is true. It then follows from [37, Proposition 10.5] that (3) is true. ∎
-We can appeal to the classical rule of Fermat [37, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a critical point of (P), then
+Overall, we have shown that (4) is true. It then follows from [42, Proposition 10.5] that (3) is true. ∎
+We can appeal to the classical rule of Fermat [42, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a stationary point of (P), then
 $0\in\partial F(x^{*})=\nabla f(x^{*})+\lambda\partial\psi(x^{*})$ (5)
@@
 $\nabla_{i}f(x^{*})\in(-\lambda r^{\prime}(|0^{+}|),+\lambda r^{\prime}(|0^{+}|)),$ $i\in\mathcal{J}^{*}.$ (7)
-Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point satisfy the necessary optimality conditions (6) and (7) are referred as the critical point.
+Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point satisfy the necessary optimality conditions (6) and (7) are referred as the stationary point.
 In the following, we demonstrate that the second-order optimality condition can be characterized by the nonzero components.
@@
     (necessary) If $x^{*}$ is a local minimum of $F$, then
-    $d^{T}\nabla^{2}f(x^{*})d-\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (8)
+    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (8)
     or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))\geq 0$.
   2. (ii)
-    (sufficient) A critical point $x^{*}$ of $F$ is a strict local minimum if
-    $d^{T}\nabla^{2}f(x^{*})d-\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (9)
+    (sufficient) A stationary point $x^{*}$ of $F$ is a strict local minimum if
+    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (9)
     or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>0$.
 Proof.
-(i) is proved in [38, Theorem 3.3]. We only prove (ii).
+(i) is proved in [43, Theorem 3.3]. We only prove (ii).
 Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of non-zero $\{u^{k}\}$ satisfying
@@
 $=$ $\ f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}^{k}r(|u_{i}^{k}|).$
-It is easy to know that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$ since $x^{*}$ is the critical point of (P) and (8) holds. Therefore for sufficiently large $k$,
+It is easy to know that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$ since $x^{*}$ is the stationary point of (P) and (8) holds. Therefore for sufficiently large $k$,
 $F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})>0.$ (11)
@@
 contradicting (10). ∎
-2.2 Strict Saddle Point
+2.2 Strict Saddle Points
 In this section, we define the strict saddle point of the problem (P).
@@
 Definition 3.
-A critical point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathcal{M}_{\mathcal{I}^{*}}\setminus\{0\}$.
-If all critical points of $F(x)$ are either strict saddle points or local minimums, $F$ satisfies the strict saddle property.
+A stationary point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$.
+If all stationary points of $F(x)$ are either strict saddle points or local minimums, $F$ satisfies the strict saddle property.
 This definition is related to the “active strict saddle point” proposed in [23], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points.
-To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the critical point on which the function is partly smooth.
+To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the stationary point on which the function is partly smooth.
 Our definition also belongs to this kind, as $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold containing $x^{*}$.
-2.3 Center Stable Manifold Theorem
-As many other works in smooth setting like [40], we interpret our algorithm as the fixed point iteration of a nonlinear map $T:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$.
+To show that the strict saddle property is mathematical generic, many works show that this property holds true for a full measure of linear perturbation [23, Theorem 2.9], [44, Theorem 2].
+Namely, the perturbed function has the strict saddle property for almost all $v\in\mathbb{R}^{n}$.
+In our case, we also have the same results by considering the following perturbed nonsmooth regularization problem for $v\in\mathbb{R}^{n}$,
+$\min\limits_{x\in\mathbb{R}^{n}}F_{v}(x):=F(x)-\langle v,x\rangle=f(x)+\lambda\psi(x)-\langle v,x\rangle.$ (13)
+To show the desired result, we need the assistance of the famous Sard’s theorem. Our proof of the strict saddle property mainly follows the proof of [45, Lemma 2.7] which is proved firstly in [46].
+Definition 4.
+Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, if the derivative at $x$ is singular (Jacobian is not full rank), then $x$ is called a critical point of $f$.
+Theorem 5.
+(Sard’s Theorem) Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, and let $C$ be the set of stationary points of $f$ in $\mathcal{M}$.
+Then $f(C)$ has measure zero in $\mathcal{N}$.
+The genericity of the strict saddle property is stated below.
+Theorem 6.
+For a full Lebesgue measure set of perturbations $v\in\mathbb{R}^{n}$, the perturbation function $F_{v}(x):=F(x)-\langle v,x\rangle$ satisfies
+$\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d\neq 0,$ (14)
+for all stationary points $x$ of $F_{v}$ with $\mathcal{I}=\mathcal{I}(x)$.
+Proof.
+Let $s\in\{+1,-1,0\}^{n}$ be a vector of sign pattern and
+$\mathcal{M}(s)=\{x\in\mathbb{R}^{n}\mid\text{sign}(x)=s\}$
+be a manifold.
+Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$ with $\mathcal{I}=\mathcal{I}(s)$.
+The Jacobian of $v$ at $x$ is then given by $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$.
+By Sard’s Theorem,
+$\{[\nabla_{\mathcal{I}}F(x);v_{\mathcal{I}^{c}}]\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$
+$\implies$ $\{v\mid v=v(x),\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$
+$\implies$ $\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}\}\text{ has measure 0 in }\mathbb{R}^{n}.$
+Now we consider all the possible sign pattern vectors $s\in\{+1,-1,0\}^{n}$. There are at most $3^{n}$ different patterns. Therefore,
+$\bigcup_{s\in\{+1,-1,0\}^{n}}\mathcal{V}(s)=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathbb{R}^{n}\}$
+has measure 0 in $\mathbb{R}^{n}$. ∎
+Center Stable Manifold Theorem
+As many other works in smooth setting like [47], we interpret our algorithm as the fixed point iteration of a nonlinear map $T$.
 Likewise, we define the fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as the unstable fixed points, as described below.
 Definition 5.
-(Unstable fixed point) Let $x^{*},\epsilon^{*}$ be a fixed point of a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
-$x^{*},\epsilon^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*},\epsilon^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
+(Unstable fixed point) Let $x^{*}$ be a fixed point of a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
+$x^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
 The Center Stable Manifold (CSM) Theorem plays a crucial role in deriving the avoidance of strict saddle points.
@@
 Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant splitting of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one, and greater than one.
-There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$, and $\cap_{k=0}^{\infty}T^{-k}(B)\subset W_{loc}^{cs}$.
+There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$.
+In addition, if $T^{k}(x)\in B$ for all $k\geq 0$, then $x\in W_{loc}^{cs}$.
+If $\nabla T(0)$ is invertible, then $T$ is locally a diffeomorphism around 0.
 The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero.
-Such result was established in [40, Theorem 2] for a global diffeomorphism map, under which the preimage of a measure zero set also has measure zero.
-The following corollary in [23] shows the same conclusion for lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous).
+This requires the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero.
+Such result was established in [47, Theorem 2] for a global diffeomorphism map and then for a globally lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [23].
+Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown by using the fact that $\mathbb{R}^{n}$ is second countable no matter the number of stationary points are countable (see [23, Corollary 2.12] for a standard proof).
+We list their results below.
+Theorem 8.
+[47, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$.
+Let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ Then the set of initial points that converge to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
 Corollary 1.
 [23, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $\nabla T(x)$ is invertible.
-Then, the set of initial points attracted by such fixed points
-$W:=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$
-has zero Lebesgue measure.
+Then, the set of initial points attracted by such fixed points $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim\limits_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
+To summarize, the analysis of escaping from strict saddle points under the strict saddle property generally consists of three steps:
+  1. (1).
+    Interpret the iterative algorithm as a fixed point iteration, and build the equivalence of the unstable fixed points and the strict saddle points.
+  2. (2).
+    Show the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at fixed point), so that the CSM theorem guarantees the avoidance of converging to the strict saddle point.
+  3. (3a).
+    Show the fixed point iteration is globally a diffeomorphism.
+  4. (3b).
+    (Alternative to (3a)) Show the fixed point iteration is globally a lipeomorphism.
 3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
@@
 where $w(x_{i})=r^{\prime}(|x_{i}|)$.
-To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a critical point and a Lipeomorphism globally.
+To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a Lipeomorphism globally.
 However, this is not true in our case since $S$ is generally non-invertible.
-To see this, one can rewrite the map $S$ using the soft-threshold operator with a changing threshold as
-$S(x^{k})=\mathbb{S}_{\lambda w^{k}}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})).$
 This algorithm may quickly converges to a local minimizers due to the rapid growth of the weight values, especially for the case when $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered.
 To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
-$\min\limits_{\theta\in\mathbb{R}^{2n}}F(\theta):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}^{2}).$ (13)
-with $\theta:=[x;\epsilon]$ for shorthand. Here we also treat $\epsilon$ as another variable to optimize, so that the algorithm updates both $x$ and $\epsilon$ by solving the following subproblem
-$x^{k+1}=S(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w_{i}^{k}|y_{i}|,$
-with $w_{i}^{k}=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and setting $\epsilon^{k+1}=\mu\epsilon^{k}$ with $\mu\in(0,1)$.
-The subproblem solution can be given explicitly using the solft-thresholding operator. This update can be written as
-$S(x^{k})=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$ (14)
+$\min\limits_{x,\epsilon\in\mathbb{R}^{n}}F(x,\epsilon):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}).$ (15)
+Here we also treat $\epsilon$ as variables, so that the algorithm updates $x$ and $\epsilon$ by $(x^{k+1},\epsilon^{k+1})=S(x^{k},\epsilon^{k})$
+$\begin{cases}x^{k+1}=S^{x}(x^{k},\epsilon^{k})=\arg\min\limits_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum\limits_{i=1}^{n}w_{i}^{k}|y_{i}|,\\ \epsilon^{k+1}=S^{\epsilon}(x^{k},\epsilon^{k})=\mu\epsilon^{k}\end{cases}$
+with $w_{i}^{k}=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and $\mu\in(0,1)$.
+The subproblem solution can be given explicitly using the solft-thresholding operator,
+$S^{x}(x^{k},\epsilon)=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$ (16)
 or explicitly,
-$S_{i}(x^{k})=\left\{\begin{aligned} &x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{ if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ &x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{ if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ &0,&\text{otherwise }.\end{aligned}\right.$ (15)
-The map $S$ is not a bijection, as it maps multiple values to 00.
-To address this issue, we need to revise this algorithm to ensure it satisfy these conditions.
-Motivated by the technique used in [23], we modified the algorithm with following damping iteration,
-$x^{k+1}=(1-\alpha)x^{k}+\alpha S(x^{k},\epsilon^{k}),$
-where $\alpha\in(0,1)$ is the damping parameter.
-Intuitively, if $S_{i}(x^{k},\epsilon^{k}_{i})=0$, the damping operator maps $x^{k}_{i}$ to $(1-\alpha)x^{k}_{i}$, resulting in a bijective map.
-We can use the iterate $(x;\epsilon)\in\mathbb{R}^{2n}$ and equivalently write the update of $(x^{k};\epsilon^{k})$ as
-$(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S(x^{k},\epsilon^{k})\\ \mu\epsilon^{k}\end{bmatrix}$ (16)
-with $\epsilon_{0}>0$, $\mu\in(0,1)$ and $\alpha\in(0,1)$.
+$S^{x}_{i}(x^{k},\epsilon^{k})=\begin{cases}x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ 0,&\text{otherwise }.\end{cases}$ (17)
+$S$ is not a bijection, as it maps multiple values to 00. Motivated by the technique used in [23], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
+$(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S^{x}(x^{k},\epsilon^{k})\\ (1-\alpha(1-\mu))\epsilon^{k}\end{bmatrix}$ (18)
+with $\epsilon_{0}\in\mathbb{R}^{n}_{++}$ and $\alpha\in(0,1)$.
 This algorithm, hereinafter nicknamed DIRL1, is presented in Algorithm 10.
@@
 Proposition 9.
-$x^{*}$ is a critical point of (P) if and only if $x^{*}=S(x^{*},0)$. In addition, $x^{*}$ is a critical point of (P) if and only if $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$.
+$x^{*}$ is a stationary point of (P) if and only if $x^{*}=S^{x}(x^{*},0)$. In addition, $x^{*}$ is a stationary point of (P) if and only if $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$.
 Proof.
-Suppose $x^{*}$ is a critical point.
-It follows from (6) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}(x^{*},0)=x_{i}^{*}$.
+Suppose $x^{*}$ is a stationary point.
+It follows from (6) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}^{x}(x^{*},0)=x_{i}^{*}$.
 This result also holds true by the same argument if $x_{i}^{*}<0$.
@@
 Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$.
-Therefore, $S_{i}(x^{*},0)=0=x_{i}^{*}$.
-Overall, we have shown that $x^{*}=S(x^{*},0)$ in all cases.
-Conversely, suppose $x^{*}=S(x^{*},0)$.
-If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$.
-Same argument yields $0=x_{i}^{*}-S_{i}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$.
+Therefore, $S_{i}^{x}(x^{*},0)=0=x_{i}^{*}$.
+Overall, we have shown that $x^{*}=S^{x}(x^{*},0)$ in all cases.
+Conversely, suppose $x^{*}=S^{x}(x^{*},0)$.
+If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$.
+Same argument yields $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$.
 Therefore, $\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)=0$ for $i\in\mathcal{I}^{*}$.
@@
 Hence, $\nabla_{i}f(x^{*})\in[-\lambda r^{\prime}(|0^{+}|),-\lambda r^{\prime}(|0^{+}|)]$.
-Overall, $x^{*}$ satisfies (5) and is a critical point of (P).
+Overall, $x^{*}$ satisfies (5) and is a stationary point of (P).
 By Assumption 2, it satisfies (6) and (7).
@@
 3:repeat
-4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})$.
+4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$.
 5:     Compute new iterate:
@@
 7:     $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$.
-8:     $\epsilon^{k+1}=\mu\epsilon^{k}$.
+8:     $\epsilon^{k+1}=(1-\alpha(1-\mu))\epsilon^{k}$.
 9:     Set $k=k+1$.
@@
 Algorithm 1 Damped Iteratively Reweighted $\ell_{1}$ Algorithm (${\rm DIRL}_{1}$)
-4 Convergence analysis
-The analysis focuses solely on the algorithm involving the update of $\epsilon$.
-It is noteworthy the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero.
-We first make the following assumptions for the algorithm.
+3.1 Convergence analysis
+It is noteworthy that the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.
 Assumption 3.
@@
 This assumption means $f$ is Lipschitz continuous and Lipschitz differentiable on $\mathcal{L}(F^{0})$ with constant $L_{f}$, i.e., $\forall x,y\in\mathbb{R}^{n}$ (we use the same constant for $f$ and $\nabla f$ for simplicity).
-$\|f(x)-f(y)\|\leq L_{f}\|x-y\|\text{ \ and \ }\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|.$ (17)
-We first show the monotonicity of $F(x,\epsilon)$ along a direction. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
+$|f(x)-f(y)|\leq L_{f}\|x-y\|\text{ \ and \ }\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|.$ (19)
+This assumption also implies that if $r^{\prime}(|0^{+}|)=\infty$, then for any stationary point $x^{*}$, $|x_{i}^{*}|\geq(r^{\prime})^{-1}(L_{f}/\lambda),\forall i\in\mathcal{I}^{*}$ by (6).
+Therefore, we know $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above and this is also the case for $r^{\prime}(|0^{+}|)<\infty$.
+Overall, we can conclude
+$\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$ (20)
+on all stationary points in $\mathcal{L}(F^{0})$.
+We first show the monotonicity of $F(x,\epsilon)$. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
 $\Delta F(x^{k+1},\epsilon^{k+1})=$ $\ F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k+1}),$
@@
 Since the concavity of $r(\cdot)$ on $\mathbb{R}_{++}$ gives $r(x_{1})\leq r(x_{2})+r^{\prime}(x_{2})(x_{1}-x_{2})$ for any $x_{1},x_{2}\in\mathbb{R}_{++}$, we can imply that for $i=1,...,n$,
-$r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})+r^{\prime}(|x_{i}^{k}|+(\epsilon^{k}_{i})^{2})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
+$r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+\epsilon_{i}^{k})+r^{\prime}(|x_{i}^{k}|+\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
 Summing up the above inequality over $i$ yields
-$\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+(\epsilon_{i}^{k})^{2})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$ (19)
+$\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+\epsilon_{i}^{k})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+\epsilon_{i}^{k})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$ (22)
 On the other hand, it follows that
@@
 Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$.
-Then $\mu^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
+Then $\epsilon^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
 Proof.
@@
 Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$.
+This also implies $\text{sign}(y_{i}^{k})=\text{sign}(x_{i}^{*}),i\in\mathcal{I}^{*}$ for sufficiently large $k\in\mathcal{S}$.
 The optimality condition of the $k$th subproblem implies
 $-\beta(y_{i}^{k}-x_{i}^{k})\in\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k})\partial|y_{i}^{k}|.$
-Then we consider the two kinds of cases with $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$.
-If $r^{\prime}(|0^{+}|)=\infty$, as shown in Proposition 2, we only need to consider $i\in\mathcal{I}^{*}$. Taking the limit on $\mathcal{S}$,
-00 0 $=\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(y_{i}^{k})+\beta(y_{i}^{k}-x_{i}^{k})|$
-$=\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(x_{i}^{k})+\frac{\beta}{\alpha}(x_{i}^{k+1}-x_{i}^{k})|$
-$\geq\lim\limits_{k\rightarrow\infty\atop k\in\mathcal{S}}|\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k}){\rm sign}(x_{i}^{k})|-\frac{\beta}{\alpha}\|x^{k}-x^{k+1}\|$
-$=|\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0){\rm sign}(x_{i}^{*})|$
-$=|\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|){\rm sign}(x_{i}^{*})|.$
-which satisfy (5). Therefore, $x^{*}$ is stationary.
-If $r^{\prime}(|0^{+}|)<\infty$, we need also to consider $i\in\mathcal{J}^{*}$.
-By [41, Proposition 2.1.5] and [37, Theorem 5.19], $\partial r$ is outer semi-continuous and there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that
-$0=\nabla_{i}f(x^{*})+\lambda w(0,0)\xi^{*}_{i}=\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)\text{sign}(x_{i}^{*}),$
-meaning (5) is satisfied. Therefore, $x^{*}$ is stationary. ∎
-5 Avoidance of strict saddle point
-DIRL1 has the property of so-called “active manifold identification”, which belongs to the research line of active-manifold identification in nonsmooth optimization [42, 43].
-It means the iterates fall onto a smooth manifold after finite iterations.
+If $w(x_{i}^{*},0)<+\infty$, it follows from [48, Proposition 2.1.5], [42, Theorem 5.19] and the outer semi-continuity of $\partial r$ that there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that
+$0=\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)\xi^{*}_{i}$
+as $y_{i}^{k}-x_{i}^{k}\overset{\mathcal{S}}{\to}0$, $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$.
+Therefore, for $w(x_{i}^{*},0)=r^{\prime}(|x_{i}^{*}|)<\infty,i\in\mathcal{I}^{*}$, (6) is satisfied; for $w(x_{i}^{*},0)=r^{\prime}(|0^{+}|)<\infty,i\in\mathcal{J}^{*}$, (7) is satisfied by Assumption 2.
+Therefore, $x^{*}$ is stationary.
+∎
+3.2 Avoidance of strict saddle point
+The map $S$ has the property of so-called “active manifold identification”, which means the iterates fall onto a smooth manifold after finite iterations.
 The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consists of the nonzeros at a limit point.
 Theorem 12.
-(Model Identification and Local Smoothness) Under Assumption 1-3, for any critical point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
+(Model Identification and Local Smoothness) Under Assumption 1-3, for any stationary point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
   1. (i)
-    For any point $(x,\epsilon)\in B$, $S(x,\epsilon)\in(x^{*}+\mathcal{M}_{\mathcal{I}^{*}})$ and $\text{sign}(S(x,\epsilon))=\text{sign}(x)=\text{sign}(x^{*})$.
+    For any point $(x,\epsilon)\in B$, $S^{x}(x,\epsilon)\in(x^{*}+\mathcal{M}_{\mathcal{I}^{*}})$ and $\text{sign}(S^{x}(x,\epsilon))=\text{sign}(x)=\text{sign}(x^{*})$.
   2. (ii)
-    Map $S$ is $C^{1}$-smooth in $B$ and $S(x)=x-\frac{1}{\beta}g(\theta)$ with
-    $g_{i}(\theta)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
+    $S$ is $C^{1}$-smooth in $B$ and $S^{x}(x,\epsilon)=x-\frac{1}{\beta}g(x,\epsilon)$ with
+    $g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
   3. (iii)
@@
 Proof.
-(i) Since $S$ is a continuous map and $x^{*}=S(x^{*},\epsilon^{*})$, it holds that
+(i) Since $S$ is a continuous map and $x^{*}=S^{x}(x^{*},\epsilon^{*})$, it holds that
 $\text{sign}(x_{i})=\text{sign}(x^{*}_{i})\text{\ and\ }\text{sign}(S_{i}(x,\epsilon))=\text{sign}(S_{i}(x^{*},0))=\text{sign}(x^{*}_{i}),\quad i\in\mathcal{I}^{*}$
@@
 (ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agree with following constrain problem
-$S(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (24)
+$S^{x}(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (27)
 $=$ $\ \arg\min\limits_{x\in\mathbb{R}^{n}}\{g(x^{k},\epsilon^{k})^{T}x+\frac{\beta}{2}\|x-x^{k}\|^{2}\},$
@@
 (iii) It follows from (27) that the iteration $T$ is $C^{1}$-smooth in $B$ for any for any $(x,\epsilon)\in B$. ∎
-The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent for the variables in $\mathcal{I}^{*}$ locally around a critical point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$.
+The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent for the variables in $\mathcal{I}^{*}$ locally around a stationary point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$.
 Specifically, it can be expressed as
-$T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &\mu\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$ (25)
+$T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &(1-\alpha(1-\mu))\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$ (28)
 around $(x^{*},\epsilon^{*})$.
@@
 Theorem 13.
-(Unstable Fixed Point) Suppose $x^{*}$ is a critical point of (P).
+(Unstable Fixed Point) Suppose $x^{*}$ is a stationary point of (P).
 $x^{*}$ is a strict saddle point of (P) if and only if the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$.
+Moreover, if $\alpha<\frac{\beta}{\rho}$, $DT(x^{*},\epsilon^{*})$ is invertible.
 Proof.
-First of all, we calculate the Jabobian of $T$ in the neighborhood $B$ described in Theorem 8. It holds that
-$\frac{\partial g}{\partial\epsilon_{i}}(x,\epsilon)=\begin{cases}0&\text{ \ if}i\in\mathcal{J}^{*},\\ \lambda r^{\prime\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*})2\epsilon&\text{ \ if }i\in\mathcal{I}^{*}.\end{cases}$
-Let $D_{\epsilon}g(x,\epsilon)$ be the Jacobian matrix of $g(\cdot)$ with respect to $\epsilon$.
-It follows that $D_{\epsilon}g(x^{*},\epsilon^{*})=\bm{0}$ by the fact that $\epsilon_{i}^{*}=0$ and $r^{\prime\prime}(|x_{i}^{*}|)<+\infty$, $i\in\mathcal{I}^{*}$.
-On the other hand, $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \bm{0}&\beta I\end{bmatrix}$.
+First of all, we calculate the Jabobian of $T$ in the neighborhood $B$ described in Theorem 12.
+It holds that $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \bm{0}&\beta I\end{bmatrix}$.
 Therefore, the Jacobian of $T$ at $(x^{*},\epsilon^{*})$ is given by
-$DT^{*}(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}\ &\mu I\end{bmatrix}=\begin{bmatrix}I-\frac{\alpha}{\beta}D_{x}g(x^{*},\epsilon^{*})\ &\bm{0}\\ \bm{0}\ &\mu I\end{bmatrix}$
-$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&\bm{0}\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&\mu I\end{bmatrix}$
+$DT(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}\ &(1-\alpha(1-\mu))I\end{bmatrix}$
+$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}$
 Now we are ready to prove the equivalence between the unstable fixed point of $T$ and the strict saddle point of (P).
@@
 $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0\iff\lambda_{\max}(I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>1.$
+If $\alpha$ is selected to satisfy $01-\frac{\alpha}{\beta}\lambda_{\max}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>$ 0 and by the strict saddle property $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ does not have 0 eigenvalue, we know $DT(x^{*},\epsilon^{*})$ is invertible.
+∎
 To apply Corollary 1, we next to show $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$.
@@
 Proposition 14.
-Suppose Assumption 1-3 hold.
+(Lipeomorphism) Suppose Assumption 1-3 hold.
 Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ defined in Proposition 10, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$.
-$T$ is a lipeomorphism when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
+Then $T$ is invertible, and $T$ and $T^{-1}$ is Lipschitz continuous (i.e., $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
 Proof.
-If $r^{\prime}(|0+|)<\infty$, $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|,\epsilon_{i}^{2})$ is Lipschitz continuous in $\mathcal{L}(F^{0})$.
+If $r^{\prime}(|0+|)<\infty$, $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|,\epsilon_{i})$ is Lipschitz continuous in $\mathcal{L}(F^{0})$.
 If $r^{\prime}(|0+|)=\infty$, by Proposition 10, $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$.
@@
 with
-$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i}^{2})&\text{ if }r^{\prime}(|0+|)<\infty\\ \min(w(\underline{x},0),w(x_{i},\epsilon_{i}))=\min(r^{\prime}(\underline{x}),r^{\prime}(|x_{i}|+\epsilon_{i}^{2}))&\text{ if }r^{\prime}(|0+|)=\infty.\end{cases}$
+$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i})&\text{ if }r^{\prime}(|0+|)<\infty\\ \min(w(\underline{x},0),w(x_{i},\epsilon_{i}))=\min(r^{\prime}(\underline{x}),r^{\prime}(|x_{i}|+\epsilon_{i}))&\text{ if }r^{\prime}(|0+|)=\infty.\end{cases}$
 Therefore, $u(\cdot)=\min(r^{\prime}(\underline{x}),r^{\prime}(\cdot))$ is also Lipschitz continuous in $\mathcal{L}(F^{0})$.
@@
 Next we prove $S$ is Lipschitz continuous. It follows that
-$S(\hat{x},\hat{\epsilon})-S(\tilde{x},\tilde{\epsilon})=$ $\ \mathbb{S}_{\lambda\hat{u}/\beta}(\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x}))-\mathbb{S}_{\lambda\tilde{u}/\beta}(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))$
+${\|}S^{x}(\hat{x},\hat{\epsilon})-S^{x}(\tilde{x},\tilde{\epsilon}){\|}=$ $\ {\|}\mathbb{S}_{\lambda\hat{u}/\beta}(\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x}))-\mathbb{S}_{\lambda\tilde{u}/\beta}(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x})){\|}$
 $\leq$ $\ \|\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x})-(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))\|+\|\hat{u}-\tilde{u}\|$
@@
 where the first inequality is by Lemma A.3 and the second is by (19) and (30).
-Therefore, $\|\alpha(\hat{x}-S(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
-Notice that $T=\begin{bmatrix}x\\ \epsilon\end{bmatrix}+\alpha(\begin{bmatrix}S(x,\epsilon)\\ \mu\epsilon\end{bmatrix}-\begin{bmatrix}x\\ \epsilon\end{bmatrix})$, Since $\begin{bmatrix}S(x,\epsilon)\\ \mu\epsilon\end{bmatrix}$ is Lipschitz with constant $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
-Applying [23, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is a lipeomorphism.
+Therefore, $\|\alpha(\hat{x}-S^{x}(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S^{x}(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
+Notice that $T=\begin{bmatrix}x\\ \epsilon\end{bmatrix}+\alpha(\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}-\begin{bmatrix}x\\ \epsilon\end{bmatrix})$, Since $\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}$ is Lipschitz with constant $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
+Applying [23, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is invertible and a lipeomorphism.
 ∎
@@
 (Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3.
-Let the condition of Proposition 10 holds.
+Let the condition of Proposition 14 holds and $\alpha<\beta/\rho$.
 Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10.
 Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
-of (P).
 Proof.
-It follows directly from Theorem 9 that map $T$ is local $C^{1}$-smooth around any critical points, and all critical points $x^{*}$ are unstable fixed points.
+It follows directly from Theorem 13 that map $T$ is local $C^{1}$-smooth around any stationary points, and all stationary points $x^{*}$ are unstable fixed points.
 By applying Proposition 14, $T$ is a lipeomorphism.
@@
 ∎
+We illustrate this property by applying DIRL1 to Example 1.
+We test $\alpha=0.1$ and $0.9$ with different initial points; first is $(3,3)$ representing an initial point far from the global minimizer and the second is $(0,0.043)$ very close to the saddle point.
+The trajectories are shown in Figure 3.
+In four cases, the algorithm always converges to the global minimizer $(0,1)$.
+(a) $\alpha=0.9$ and initial point $(3,3)$.
+(b) $\alpha=0.9$ and initial point $(0,0.043)$.
+(c) $\alpha=0.1$ and initial point $(3,3)$.
+(d) $\alpha=0.1$ and initial point $(0,0.043)$.
+Figure 3: Trajectories of DIL2.
+4 Damped Iterative Reweighted $\ell_{2}$ Algorithms
+The ${\rm IRL}_{2}$ algorithm is an effective method to solve problem (P).
+In this method, we write $h(\cdot)=r\circ\sqrt{(\cdot)}$, so that $s(x_{i})=h(x^{2}_{i})$.
+At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearize $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem,
+$y^{k}=S^{x}(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2},$
+where
+$w(x_{i},\epsilon_{i})=h^{\prime}(x_{i}^{2}+\epsilon_{i}^{2})=\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{2\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}.$
+In particular,
+$w(x_{i},0)=\begin{cases}\frac{r^{\prime}(|x_{i}|)}{2|x_{i}|}&\text{ if }x_{i}\neq 0\\ \infty&\text{ if }x_{i}=0.\end{cases}$
+The subproblem solution is given by $y^{k}=S^{x}(x^{k},\epsilon^{k})$, where the operator $S$ is defined as
+$S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}w(x_{i},\epsilon_{i})}(x_{i}-\frac{1}{\beta}\nabla_{i}f(x)).$ (31)
+If $w(x_{i},\epsilon_{i})=\infty$, we simply set $S_{i}^{x}(x,\epsilon)=0$. The damped iteratively reweighted $\ell_{2}$ algorithm, hereinafter named DIRL2 algorithm, is presented in Algorithm 10.
+1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
+2:Initialize: set $k=0$.
+3:repeat
+4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=h^{\prime}((x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2})$.
+5:     Compute new iterate:
+6:     $y^{k}=\arg\min\limits_{y\in\mathbb{R}^{n}}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2}.$
+7:     $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$.
+8:     $\epsilon^{k+1}=\mu\epsilon^{k}$.
+9:     Set $k=k+1$.
+10:until Convergence
+Algorithm 2 Damped Iteratively Reweighted $\ell_{2}$ Algorithm (${\rm DIRL}_{2}$)
+The fixed point iteration of DIRL2 is then given as $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=(1-\alpha)I+\alpha S$, where
+$T(x,\epsilon):=\begin{bmatrix}(1-\alpha)x+\alpha S^{x}(x,\epsilon)\\ (1-\alpha(1-\mu))\epsilon\end{bmatrix}.$ (32)
+We have the following result to connect the stationary point of (P) and the fixed point of $T$.
+Proposition 16.
+If $x^{*}$ is a stationary point of (P), then $x^{*}=S^{x}(x^{*},0)$ and $(x^{*},0)=T(x^{*},0)$.
+Conversely, for any $x^{*}$ satisfying $x^{*}=S^{x}(x^{*},0)$, if $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, then $x^{*}$ is a stationary point of (P).
+Proof.
+Suppose $x^{*}$ is a stationary point.
+It follows from (6) that for $i\in\mathcal{I}^{*}$, $x_{i}^{*}=x_{i}^{*}-\tfrac{1}{\beta}(\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*}))$ implying $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$.
+Therefore, $x_{i}^{*}=S_{i}(x^{*},0)$ since $w(x_{i}^{*},0)=\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}$.
+As for $i\in\mathcal{J}^{*}$, $w(0,0)=\infty$, implying that $0=S_{i}(x^{*},0)$.
+To summarize, $x^{*}=S^{x}(x^{*},0)$.
+Now consider $x^{*}=S^{x}(x^{*},0)$.
+If $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, we only have to verify (6) is satisfied by $x^{*}$.
+It follows from $x^{*}_{i}=S_{i}(x^{*},0),i\in\mathcal{I}^{*}$ that $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$, implying $\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*})=0$, completing the proof.
+∎
+The converse statement of Proposition 16 may not be true when $\mathcal{J}^{*}\neq\emptyset$ and $r^{\prime}(|0^{+}|)<\infty$.
+Since for any $x^{*}$ satisfying $x_{i}^{*}=S_{i}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it holds naturally true that $w(0,0)=\infty$ and $S_{i}(x^{*},0)=0$.
+However, there is no guarantee that $x^{*}$ satisfies (7).
+In those cases, $S$ may have two types of fixed points—those are stationary points and others are not.
+This represents a stark difference from the DIRL1 algorithm, where all the fixed points of the iteration map are stationary points.
+4.1 Convergence analysis
+In this subsection, we show that DIRL2 only converge to its fixed points which are the stationary points of (P).
+Notice that this is true in all cases even though it may have fixed points that are not the stationary points, as discussed in the previous subsection.
+Similar to the analysis of DIRL1 algorithm, the analysis of DIRL2 is also based on the monotonic decrease of the objective $F(x,\epsilon)$.
+We also need Assumption 3 and consequently (19) hold true for the analysis.
+However, it must be noticed that the definition of $F(x,\epsilon)$ is different from that for DIRL1 algorithm.
+Proposition 17.
+Suppose Assumption 1-3 hold.
+Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$.
+It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and the reduction satisfies
+$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
+Proof.
+Since $h(\cdot)=r\circ\sqrt{(\cdot)}$ is also concave by Assumption 1, it follows that for $i=1,...,n$,
+$h(x_{i}^{k+1})^{2}+(\epsilon_{i}^{k})^{2})\leq h(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2})+h^{\prime}((x_{i}^{k})^{2}+(\epsilon^{k}_{i})^{2})((x_{i}^{k+1})^{2}-(x_{i}^{k})^{2}),$
+which is equivalent to
+$r(\sqrt{(x_{i}^{k+1})^{2}+(\epsilon_{i}^{k})^{2}})\leq r(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})+w(x_{i}^{k},\epsilon^{k}_{i})((x_{i}^{k+1})^{2}-(x_{i}^{k})^{2}).$
+Following the same argument for (22)-(24), we have
+$\Delta F(x^{k+1},\epsilon^{k+1})\geq\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}.$ (33)
+For each subproblem, $y^{k}$ satisfies the optimal condition
+$\nabla f(x^{k})+2\lambda W(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})=0.$ (34)
+Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimal condition (25), the following holds
+$\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$ (35)
+$=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})((x_{i}^{k})^{2}-(x_{i}^{k+1})^{2})-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
+$\geq$ $\ \nabla f(x^{k})^{T}(\alpha x^{k}-\alpha y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(\alpha(x_{i}^{k})^{2}-\alpha(y_{i}^{k})^{2})-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
+$\geq$ $\ \alpha[\nabla f(x^{k})^{T}(x^{k}-y^{k})+2\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}(x_{i}^{k}-y_{i}^{k})]-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
+$=$ $\ \alpha[\nabla f(y^{k})+2\lambda W(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})]^{T}(x^{k}-y^{k})+\frac{\alpha\beta(2-\alpha)}{2}\|x^{k}-y^{k}\|^{2}$
+$=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2},$
+where the first inequality is by Lemma A.2 and the second is by the convexity of $(\cdot)^{2}$. Combine (24), (35) and get
+$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
+Replacing $k$ with $t$ and summing up from $t=0$ to $k-1$, we have
+$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
+It follows that $\{(x^{k},\epsilon^{k})\}\in\mathcal{L}(F^{0})$ since $F(x^{k},\epsilon^{k})\leq F(x^{0},\epsilon^{0})$.
+By Assumption 3, it is bounded and $\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})>-\infty$.
+Taking $k\rightarrow\infty$, we have
+$(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
+which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$.
+Moreover, it follows from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (21) holds for any $k\in\mathbb{N}$.
+∎
+With the monotonicity of $F$, we derive the global convergence of the ${\rm DIRL}_{1}$ method.
+Theorem 18.
+(Global convergence) Suppose Assumption 1 and 3 hold.
+Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{2}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$.
+Then $\mu^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
+Proof.
+Firstly, boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$.
+Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$.
+The optimality condition of the $k$th subproblem implies
+$-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}y_{i}^{k}.$ (36)
+As $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$, $y_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$, $\lim\limits_{k\in\mathcal{S}}\frac{y_{i}^{k}}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}=\text{sign}(x_{i}^{*})$ for $i\in\mathcal{I}^{*}$, and $\limsup\limits_{k\in\mathcal{S}}\frac{|y_{i}^{k}|}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}\leq 1$ for $i\in\mathcal{J}^{*}$.
+Therefore, taking $k\to\infty,k\in\mathcal{S}$ in both sides of (36), we have for $i\in\mathcal{I}^{*}$
+$\nabla_{i}f(x^{*})+2\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*})=0,$
+satisfying (6). As for $i\in\mathcal{J}^{*}$ and $r^{\prime}(|0^{+}|)<\infty$,
+$|\nabla_{i}f(x^{*})|=\lim_{k\in\mathcal{S}\atop k\to\infty}|\nabla_{i}f(x^{k})|\leq\lim_{k\in\mathcal{S}\atop k\to\infty}|\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}y_{i}^{k}|=\limsup\limits_{k\in\mathcal{S}\atop k\to\infty}\frac{r^{\prime}(|0^{+}|)|y_{i}^{k}|}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}\leq r^{\prime}(|0^{+}|),$
+satisfying (7) by Assumption 2. Overall, $x^{*}$ is stationary. ∎
+4.2 Avoidance of strictly saddle points
+In this subsection, we further show that DIRL2 algorithm does not converge to fixed points that are stationary points and strict saddle points.
+Unlike DIRL1, the map $S$ in DIRL2 does not enjoy the property of “active manifold identification” or “model identification”.
+There is no guarantee that the iteration map is $C^{1}$-smooth without this property.
+The following additional assumption is needed to guarantee the local smoothness.
+One can verify that most sparse non-convex penalty in Table 1 like LPN, TAN, FRA, EXP and LOG are satisfied.
+Assumption 4.
+$\lim\limits_{z\rightarrow 0}\frac{r^{\prime}(|z|)}{|z|}=\infty\quad\text{ and }\quad\lim\limits_{z\rightarrow 0}z\frac{r^{\prime\prime}(|z|)}{r^{\prime}(|z|)^{2}}\to 0.$
+To show the smoothness of $S$, we first calculate the Jacobian of $S$. Define
+$g(z):=\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|z|)}{|z|}}\ \text{ with }g(0):=0.$
+Then $S_{i}^{x}(x,\epsilon)=g(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ with $z=\sqrt{x^{2}+\epsilon^{2}}$. We first have the following result about $g$, $S$ and $T$.
+Theorem 19.
+(Local smoothness and Lipschitz continuity ) Suppose Assumption 1-4 hold. The following statements hold:
+  1. (i)
+    $g(z)$ is continuously differentiable on $\mathbb{R}$, so that it is Lipschitz continuous on $\mathcal{L}(F^{0})$.
+  2. (ii)
+    $S$ is continuously differentiable and Lipschitz continuous on $\mathcal{L}(F^{0})$ with constant $L_{S}$.
+  3. (iii)
+    $T$ is a lipeomorphism for $\alpha\in(0,\frac{1}{1+L_{S}})$.
+Proof.
+(i) It follows that
+$g^{\prime}(z)=\begin{cases}-\tfrac{\lambda}{\beta}\frac{r^{\prime\prime}(|z|)z-\text{sign}(z)r^{\prime}(|z|)}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}},&\text{\ for }z\neq 0,\\ \lim\limits_{\delta\to 0}\left(\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|\delta|)}{|\delta|}}-0\right)\frac{1}{\delta}=0,&\text{\ for }z=0.\end{cases}$ (37)
+It holds that
+$\lim\limits_{z\to 0}g^{\prime}(z)=$ $-\tfrac{\lambda}{\beta}\lim\limits_{z\to 0}\left(\frac{r^{\prime\prime}(|z|)z}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}}-\frac{\text{sign}(z)r^{\prime}(|z|)}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}}\right)$
+$=$ $-\lim\limits_{z\to 0}\left(\tfrac{\beta}{\lambda}\frac{r^{\prime\prime}(|z|)z}{{r^{\prime}(|z|)}^{2}}-\text{sign}(z)\frac{|z|}{r^{\prime}(|z|)}\right)=0=g^{\prime}(0)$
+by Assumption 4. Therefore, $g$ is continuously differentiable.
+(ii) It is only needed to show the smoothness of $S$ when $\mathcal{J}^{*}\neq\emptyset$.
+We simply calculate its partial derivative and verify it is continuous at any point $(x^{*},0)\in\mathbb{R}^{2n}$.
+First, consider $S_{i},i\in\mathcal{J}^{*}$. It follows from $S_{i}^{x}(x^{*},0)=0$ and $r^{\prime}(|0^{+}|)=\infty$ that
+$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$
+$=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta$
+$=$ $-\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}g^{\prime}(0)=0.$
+As for $j\neq i$, it follows that $\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\epsilon^{*})=0$ and $\frac{\partial}{\partial\epsilon_{j}}S_{i}(x^{*},\epsilon^{*})=0$.
+Generally, for any $i$ satisfying $z_{i}:=\sqrt{x_{i}^{2}+\delta_{i}^{2}}>0$, $S$ is smooth with respect to $x_{i}$ and $\epsilon_{i}$ with $z_{i}=\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}\neq 0$.
+$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=$ $\ g(z)(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x))+\tfrac{x_{i}}{z_{i}}g^{\prime}(z_{i})(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ (38)
+$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=$ $\ \tfrac{\epsilon_{i}}{z_{i}}g^{\prime}(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$ (39)
+$\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=$ $\ -g(z)\tfrac{1}{\beta}\nabla^{2}_{ij}f(x),\quad j\neq i$ (40)
+$\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=$ $\ 0,\quad j\neq i.$ (41)
+Now for any $i\in\mathcal{J}^{*}$, as $z_{i}\to 0$, it follows from $\frac{x_{i}}{z_{i}}\leq 1$ and $zr^{\prime\prime}(z)/(r^{\prime}(z))^{2}\to 0$ that
+$\lim_{z_{i}\to 0}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}),\ \lim_{z_{i}\to 0}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})$
+$\lim_{z_{i}\to 0}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}),\ \lim_{z_{i}\to 0}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})$
+Overall, we have shown that $S$ is continuously differentiable. The Lipschitz continuity of $S$ then trivially comes from (i) and the boundedness of $\mathcal{L}(F^{0})$.
+(iii) This is straightforward by applying [23, Lemma 2.13 and 2.14]. ∎
+Now we show that every strict saddle point of (P) is an unstable fixed point of $T$.
+Theorem 20.
+(Unstable Fixed Point) Suppose $x^{*}$ is a stationary point of (P).
+If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$.
+Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},0)$ is invertible.
+Proof.
+Consider the Jacobian of $S$ at a fixed point $(x^{*},0)$. Define $H\in\mathbb{R}^{|\mathcal{I}^{*}|\times n}$ as follows:
+$\forall i\in\mathcal{I}^{*},\ h_{ii}=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{(|x_{i}^{*}|+\tfrac{\lambda}{\beta}r^{\prime}(x^{*}_{i}))\tfrac{\lambda}{\beta}\left(|x^{*}_{i}|r^{\prime\prime}(|x^{*}_{i}|)-r^{\prime}(|x^{*}_{i}|)\right)}{\left(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(|x^{*}_{i}|)\right)^{2}},$
+$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{\tfrac{\lambda}{\beta}\left(r^{\prime\prime}(|x^{*}_{i}|)-\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}\right)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
+$=$ $\ \frac{1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*})-\tfrac{\lambda}{\beta}r^{\prime\prime}(|x^{*}_{i}|)+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
+$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x^{*}_{i}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
+$\forall j\neq i,\ h_{ij}:=$ $\ \frac{\partial}{\partial x_{j}}S_{i}(x^{*},0)=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{|x_{i}^{*}|}},$
+by (38) and (40).
+Therefore, $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$,
+$DS(x^{*},0)=$ $\begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&0\\ 0&0&0\\ 0&0&\mu I\end{bmatrix}$
+$DT(x^{*},0)=$ $\begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&0\\ 0&(1-\alpha)I&0\\ 0&0&(1-\alpha(1-\mu))I\end{bmatrix}.$
+Hence, the eigenvalues of $DS(x^{*},0)$ are the same as
+$(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}),$ (42)
+where $W^{*}:=W(x^{*},0)$ and $W^{*}_{ii}<\infty,i\in\mathcal{I}^{*}$.
+If $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$, it follows from Lemma [49, Proposition 6.1] that $\lambda_{\min}((I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$.
+Hence, $DT(x^{*},0)$ has an eigenvalue greater than 1, meaning $(x^{*},0)$ is the unstable fixed point of $T$.
+Moreover, since $\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$, it is clear that for $\alpha<\beta/\rho$, $DT(x^{*},0)$ is invertible.
+∎
+Similarly, we can derive the global escape property of ${\rm DIRL}_{2}$ algorithm.
+Theorem 21.
+(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{2}$ algorithm with Assumption 1-4.
+Let the condition of Proposition 19 holds and $\alpha<\beta/\rho$.
+Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10.
+Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
+Proof.
+It follows directly from Theorem 19 and 20 that 1) map $T$ is local $C^{1}$-smooth around any stationary points; 2) all stationary points $x^{*}$ are unstable fixed points; 3) $T$ is a lipeomorphism.
+Then by Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure.
+∎
+We illustrate this property by applying DIRL2 to Example 1.
+We test $\alpha=0.1$ and $0.9$ with different initial points; first is $(3,3)$ representing an initial point far from the global minimizer and the second is $(0,0.043)$ very close to the saddle point.
+The trajectories are shown in Figure 4.
+In four cases, the algorithm always converges to the global minimizer $(0,1)$.
+(a) $\alpha=0.9$ and initial point $(3,3)$.
+(b) $\alpha=0.9$ and initial point $(0,0.043)$.
+(c) $\alpha=0.1$ and initial point $(3,3)$.
+(d) $\alpha=0.1$ and initial point $(0,0.043)$.
+Figure 4: Trajectories of DIL2.
 Appendix
@@
 Lemma A.2.
+For $x,y\in\mathbb{R}$ and $\alpha\in(0,1]$, it holds that $x^{2}-((1-\alpha)x+\alpha y)^{2}\geq\alpha(x^{2}-y^{2})$.
+Proof.
+It follows that
+$x^{2}-((1-\alpha)x+\alpha y)^{2}$ $=x^{2}-(1-\alpha)^{2}x^{2}-\alpha^{2}y^{2}-2\alpha(1-\alpha)xy$
+$=(2\alpha-\alpha^{2})x^{2}-\alpha^{2}y^{2}-2\alpha(1-\alpha)xy$
+$=\alpha(x^{2}-y^{2})+(\alpha-\alpha^{2})x^{2}+(\alpha-\alpha^{2})y^{2}-2(\alpha-\alpha^{2})xy$
+$=\alpha(x^{2}-y^{2})+(\alpha-\alpha^{2})(x-y)^{2}$
+$\geq\alpha(x^{2}-y^{2}),$
+where the inequality is by $\alpha\in(0,1]$. ∎
+Lemma A.3.
 Given $\hat{w},\tilde{w}\in\mathbb{R}^{n}_{+}$ and $\hat{z},\tilde{z}\in\mathbb{R}^{n}$, it holds that
 $\|\mathbb{S}_{\hat{w}}(\hat{z})-\mathbb{S}_{\tilde{w}}(\tilde{z})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$
-Proof.
 (a) If $|\hat{z}_{i}|\leq\hat{w}_{i}$ and $|\tilde{y}_{i}|\leq\tilde{w}_{i}$, then $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})=0$ and $\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=0$.

### v4

@@
 Abstract
-In this paper, we consider a class of non-convex and non-smooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms.
+In this paper, we consider a class of nonconvex and nonsmooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms.
 We show the second-order optimality conditions only depend on the nonzeros of the stationary points.
-We propose two damped iterative reweighted algorithms including the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems.
+We propose two damped iterative reweighted algorithms, including the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems.
 For DIRL1, we show the reweighted $\ell_{1}$ subproblem has support identification property so that DIRL1 locally reverts to a gradient descent algorithm around a stationary point.
@@
 Therefore, the map of DIRL1 and DIRL2 and their inverse are Lipschitz continuous, and the strict saddle points are their unstable fixed points.
-By applying the stable manifold theorem, these algorithms are shown to converge only to local minimizers with randomly initialization when the strictly saddle point property is assumed.
+By applying the stable manifold theorem, these algorithms are shown to converge only to local minimizers with random initialization when the strictly saddle point property is assumed.
 keywords:
@@
 where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and the penalty/regularization term $\psi(x)$ takes the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$.
-The data fitting function $f$ is twice continuously differentiable and possibly nonconvex, and $r$ satisfies the following properties.
+The data fitting function $f$ is differentiable twice continuously and possibly nonconvex, and $r$ satisfies the following properties.
 Assumption 1.
-$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$, C2-smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
-This assumption indicates $r(t)>0$ for all $t>0$, but $r$ could be nonconvex or not Lipschitz around 0.
+$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$, C2 smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
+This assumption indicates $r(t)>0$ for all $t>0$, but $r$ could be nonconvex or not Lipschitz continuous around 0.
 The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance.
 This makes it well-suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [1, 2, 3, 4].
-Many widely used nonconvex regularization such as the smoothly clipped absolute deviation penalty (SCAD) [5] or the minimax concave penalty (MCP) [6], the $\ell_{p}$ quasi-norm penalty [7] satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability.
-We provide the explicit form of several commonly used nonconvex regularization in Table 1.
-It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizer with bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) non-Lipschitz continuous regularizer with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as $\ell_{p}$ quasi-norm (LPN).
+Many widely used nonconvex regularization such as the EXP penalty [38] or the LOG penalty [39], the $\ell_{p}$ quasi-norm penalty [7] satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability.
+We provide the explicit form of several commonly used nonconvex regularizations in Table 1.
+It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizer with bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) nonLipschitz continuous regularizer with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as the $\ell_{p}$ quasi-norm (LPN).
 Numerous studies focus on gradient-based optimization methods that are guaranteed to find first-order stationary points.
@@
 Therefore, it becomes a prevalent topic, particularly in high-dimensional optimization problems, and is also the central focus of this paper.
-In smooth optimization, many methods have been developed and analyzed to escape from saddle points, including first-order methods [8, 9, 10], momentum-based methods [11], second-order optimization algorithms [12], the algorithmic frameworks for constrained nonlinear problems [13], stochastic algorithms [14, 15], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [16, 17, 18, 19].
-In smooth settings, it is widely assumed that the strict saddle property holds, meaning all saddle points possess at least one strictly negative eigenvalue.
-This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases and it has been shown to be generally satisfied for smooth problems [20, 21].
+In smooth optimization, many methods have been developed and analyzed to escape saddle points, including first-order methods [8, 9, 10], momentum-based methods [11], second-order optimization algorithms [12], the algorithmic frameworks for constrained nonlinear problems [13], stochastic algorithms [14, 15], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [16, 17, 18, 19].
+In smooth settings, it is widely assumed that strict saddle property holds, which means that all saddle points have at least one strictly negative eigenvalue.
+This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases, and it has been shown to be generally satisfied for smooth problems [20, 21].
 In nonsmooth optimization, avoiding saddle points poses a greater challenge.
-At stationary points, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate the avoidance of saddle points.
+At stationary points, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate avoidance of saddle points.
 Developing nonsmooth optimization algorithms capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than in smooth optimization.
@@
 Recent research [23] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point.
-Additionally, several proximal methods, including proximal point method, proximal gradient method and proximal linear method [23], as well as a more efficient algorithm incorporating perturbation [24], have been proposed to converge to local minima in nonsmooth optimization settings.
-In these studies, the nonsmooth objective function $r(x)$ is assumed to be $\rho$-weakly convex meaning $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex.
-They assume an active-manifold around the limit point and then define the strict saddle points as those of the $f$ restricted to the manifold.
-The weakly $\rho$-convexity for $r$ plays a crucial role in constructing C2 differentiable Moreau envelope with the same stationary points and strict saddle points in proximal methods.
+Additionally, several proximal methods, including the proximal point method, the proximal gradient method, and the proximal linear method [23], as well as a gradient method incorporating random perturbation [24], have been proposed to converge to local minima in nonsmooth optimization settings.
+In these studies, the nonsmooth objective function $r(x)$ is assumed to be $\rho$ weakly convex, which means that $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex.
+They assumed an active manifold around the limit point and then defined the strict saddle points as those of the $f$ restricted to the manifold.
+The weakly $\rho$-convexity for $r$ plays a crucial role in constructing the C2 differentiable Moreau envelope with the same stationary points and strict saddle points in the proximal methods.
 As for the regularized problem (P), several first-order methods have been proposed.
 [25] introduces a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1.
-This method is subsequently extended to address group $\ell_{p,q}$ norm regularized problems in [26].
+This method is subsequently extended to address the $\ell_{p,q}$ group norm regularized problems of the group $\ell_{p,q}$ in [26].
 Furthermore, the Difference of Convex (DC) algorithm is enlisted [27] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions.
@@
 In particular, we focus on iteratively reweighted methods which approximate the regularization term by simpler weighted $\ell_{1}$ or $\ell_{2}$ norms [34, 35, 36, 30, 37, 29, 31, 33].
-In contrast to the proximal gradient algorithm, iteratively reweighted methods solve a sequence of weighted $\ell_{1}$ or $\ell_{2}$ regularized subproblems and are widely applicable to different non-convex regularization problems.
+In contrast to the proximal gradient algorithm, iteratively reweighted methods solve a sequence of weighted $\ell_{1}$ or $\ell_{2}$ regularized subproblems and are widely applicable to various nonconvex regularization problems.
 The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator.
-This method offers a favorable local property [31], known as the model identification, under the following additional assumption.
+This method offers a favorable local property [31], known as the model/support identification, under the following additional assumption.
 Assumption 2.
@@
 At the tail end of the iteration, the sign of the iterates remain unchanged while the nonzero components are kept uniformly bounded away from 0.
-In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{J}$ as follows after finite number of iterations.
+In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{I}$ as follows after a finite number of iterations.
 Definition 1.
-(Model Subspace) The model subspace defined by $\mathcal{J}\subset\{1,\ldots,n\}$ is
-$\mathcal{M}_{\mathcal{J}}:=\{u\in\mathbb{R}^{n}:\forall j\in\mathcal{J},u_{j}=0\}.$
-This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{J}}$.
-This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) on $\mathcal{M}_{\mathcal{J}}$, even though the underlying optimization problem is nonsmooth in $\mathbb{R}^{n}$.
-The iterative reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these non-convex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the solution is trivially given.
-Compared with ${\rm IRL}_{1}$, ${\rm IRL}_{2}$ does not enjoy the model subspace identification/active-manifold identity property.
+(Model Subspace) The model subspace defined by $\mathcal{I}\subset\{1,\ldots,n\}$ is
+:=$\mathcal{M}_{\mathcal{I}}:=\{u\in\mathbb{R}^{n}:\forall j\in\mathcal{I}^{c},u_{j}=0\}.$
+This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{I}}$.
+This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) in $\mathcal{M}_{\mathcal{I}}$, although the underlying optimization problem is nonsmooth in $\mathbb{R}^{n}$.
+The iterative reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these nonconvex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the solution is trivially given.
+Compared to ${\rm IRL}_{1}$, ${\rm IRL}_{2}$ does not enjoy the model subspace identification/active-manifold identity property.
 In fact, it can be often witnessed that this algorithm can generate iterates with elements asymptotically approaching zero.
@@
 (b) Contour of $F$ around $(0,0)$.
-Let $g(x;x^{k})=\nabla f(x^{k})^{T}(x-x^{k})+\tfrac{1}{2}\|x-x^{k}\|^{2}$. The iterations of iteratively reweighted algorithms are as follows [29]. We use the ${\rm IRL}_{1}$ algorithm with iteration
+Let $g(x;x^{k})=\nabla f(x^{k})^{T}(x-x^{k})+\tfrac{1}{2}\|x-x^{k}\|^{2}$.
+The iterations of the iteratively reweighted algorithms are as follows [29].
+We use the ${\rm IRL}_{1}$ algorithm with iteration
 $x^{k+1}=\arg\min_{x_{1},x_{2}}g(x;x^{k})+\tfrac{1}{2}\frac{1}{(|x_{1}^{k}|+0.1^{k})^{1/2}}|x_{1}|+\tfrac{1}{2}\frac{1}{(|x_{2}^{k}|+0.1^{k})^{1/2}}|x_{2}|$
@@
 Figure 2: Trajectories generated by iteratively reweighted methods. ${\rm IRL}_{1}$ eventually falls into the subspace (the $x_{2}$-axis), while ${\rm IRL}_{2}$ has nonzero elements for all iterates.
-The behavior of ${\rm IRL}_{2}$ brings another challenge of analyzing the avoidance of strict saddle point.
+The behavior of ${\rm IRL}_{2}$ brings another challenge in analyzing the avoidance of strict saddle point.
 Current analysis for nonsmooth optimization algorithms ubiquitously depends on the active-manifold identification property, which ensures the algorithm reverts to a smooth algorithm in the active manifold in a local neighbourhood of the stationary point.
@@
   1. (i)
-    We first show the local minimal and second-order optimality of $f$ are equivalent to those of $f$ restricted on $\mathcal{M}_{\mathcal{J}}$, and define the strict saddle points as those on $\mathcal{M}_{\mathcal{J}}$.
+    We first show the local minimal and second-order optimality of $f$ are equivalent to those of $f$ restricted on $\mathcal{M}_{\mathcal{I}}$, and define the strict saddle points as those on $\mathcal{M}_{\mathcal{I}}$.
   2. (ii)
-    We develop damped IRL1 algorithms individually for the two types of regularizers, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning the next iterate is the combination of a traditional IRL1 iterate and its previous iterate.
+    We develop damped IRL1 algorithms individually for the two types of regularizer, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning that the next iterate is the combination of a traditional IRL1 iterate and its previous iterate.
 The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ from becoming $\infty$ at each iteration.
-At the tail end of the iterations, we show this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{J}$ on the subspace $\mathcal{M}_{\mathcal{J}}$ and decrease the rest components to 0 at a linear speed.
+At the tail end of the iterations, we show that this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{I}$ on the subspace $\mathcal{M}_{\mathcal{I}}$ and decrease the rest components to 0 at a linear speed.
 We then apply the center stable manifold theorem and prove that the algorithm can avoid the convergence to strict saddle points.
   3. (iii)
-    We develop damped IRL2 algorithms for the nonconvex regularized problems and provide the global convergence.
+    We develop damped IRL2 algorithms for nonconvex regularized problems and provide global convergence.
 Under an additional assumption which is satisfied by most common nonconvex regularizers, we show that the algorithm is $C^{1}$ smooth and nondegenerate everywhere.
 We also proved the equivalence between the unstable fixed point of the iteration map and the strict saddle points of the regularization problem.
-Therefore, the avoidance of strict saddle points are derived.
-This is the first algorithm shown to have this property without using the active-manifold identification property nor the stochastic technique in nonsmooth optimization algorithms.
+Therefore, the avoidance of strict saddle points is derived.
+This is the first algorithm shown to have this property without using the active-manifold identification property or the stochastic technique in nonsmooth optimization algorithms.
 1.2 Outline
@@
 […]
-${\rm sign}(t)=\left\{\begin{aligned} &1,\ &t>0,\\ &-1,\ &t<0,\\ &0,\ &t=0.\end{aligned}\right.$ (1)
+We denote $\|\cdot\|_{p}$ as the quasi-norm $\ell_{p}$ with $p\in(0,1)$, that is, $\|x\|_{p}=(\sum_{i=1}^{n}|x_{i}|^{p})^{\frac{1}{p}}$.
+This does not define a proper norm because of the lack of subadditivity.
+For $x\in\mathbb{R}^{n}$, $x_{i}$ denotes the $i$-th element of $x$.
+The index set of nonzero and zero components of $x$ is defined as $\mathcal{I}(x)=\{i\mid x_{i}\neq 0\}$, and $\mathcal{J}(x)=\{i\mid x_{i}=0\}$, respectively.
+Moreover, $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ denotes $\mathcal{I}(x^{*})$ and $\mathcal{J}(x^{*})$ for short.
+${\rm diag}(x)$ is the diagonal matrix by placing elements of vector $x$ on the main diagonal.
+${\rm sign}(x):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the sign function defined by
+$[{\rm sign}(x)]_{i}=\left\{\begin{aligned} &1,\ &x_{i}>0,\\ &-1,\ &x_{i}<0,\\ &0,\ &x_{i}=0.\end{aligned}\right.$ (1)
 $DT(x)$ denote the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$.
 Given $x\in\mathbb{R}^{n}$ and $\mathcal{I}\subset\{1,\ldots,n\}$, denote $x_{\mathcal{I}^{*}}\subset\mathbb{R}^{\mathcal{I}^{*}}$ or $[x_{i}]_{i\in\mathcal{I}}$ as the subvector of the elements in $\mathcal{I}$; also, given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants.
-If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$.
+If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}f(x)=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$.
 Denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively.
@@
 Proof.
-We only need calculate the subgradients of $s$ at 00, i.e., $i\in\mathcal{J}(x)$. If $r^{\prime}(|0^{+}|)<\infty$, define for any $x\in\mathbb{R}$,
+We only need to calculate the subgradients of $s$ at 00, i.e., $i\in\mathcal{J}(x)$. If $r^{\prime}(|0^{+}|)<\infty$, define for any $x\in\mathbb{R}$,
 $\hat{s}(x):=\begin{cases}r(|x|)&\text{ if }x\geq 0,\\ r^{\prime}(|0^{+}|)x&\text{ otherwise}.\end{cases}$
@@
 By [42, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
-On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, meaning $r^{\prime}(|0^{+}|),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$.
+On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, which means $r^{\prime}(|0^{+}|)\in\hat{\partial}s(0),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$.
 It follows from [42, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$.
@@
 Hence, ^$\hat{\partial}s(0)=\partial s(0)=\mathbb{R}$.
-Overall, we have shown that (4) is true. It then follows from [42, Proposition 10.5] that (3) is true. ∎
+In general, we have shown that (4) is true. It then follows from [42, Proposition 10.5] that (3) is true. ∎
 We can appeal to the classical rule of Fermat [42, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a stationary point of (P), then
@@
 $\nabla_{i}f(x^{*})\in(-\lambda r^{\prime}(|0^{+}|),+\lambda r^{\prime}(|0^{+}|)),$ $i\in\mathcal{J}^{*}.$ (7)
-Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point satisfy the necessary optimality conditions (6) and (7) are referred as the stationary point.
+Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point that satisfies the necessary optimality conditions (6) and (7) is referred to as the stationary point.
 In the following, we demonstrate that the second-order optimality condition can be characterized by the nonzero components.
@@
 (i) is proved in [43, Theorem 3.3]. We only prove (ii).
-Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of non-zero $\{u^{k}\}$ satisfying
+Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of nonzero $\{u^{k}\}$ satisfying
 $F(x^{*}+u^{k})\leq F(x^{*})\quad{\rm and}\quad\lim\limits_{k\rightarrow\infty}u^{k}=0.$ (10)
@@
 It follows from $u^{k}=u_{\mathcal{I}^{*}}^{k}+u_{\mathcal{J}^{*}}^{k}$ that
-$F(x^{*}+u^{k})-F(x^{*})$
+$F(x^{*}+u^{k})-F(x^{*})$ (11)
 $=$ $\ f(x^{*}+u^{k})+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
@@
 $=$ $\ f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}^{k}r(|u_{i}^{k}|).$
-It is easy to know that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$ since $x^{*}$ is the stationary point of (P) and (8) holds. Therefore for sufficiently large $k$,
+From the classic optimality condition for smooth function and (9), it is easy to see that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$. Therefore for sufficiently large $k$,
 $F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})>0.$ (12)
 Moreover, by the Lagrangian mean value theorem, there exists sufficiently small $\epsilon^{k}\in(0,\min_{i}\{|u_{i}^{k}|\})$ such that
-$r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$
-The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}\|_{\infty})$; consequently,
+$r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$ (13)
+The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}\|_{\infty})$, which, together with (13), yields that
 $\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{1}\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{2}.$
-This, combined with (11), yields
+This, combined with (12) and (11), yields
 $F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|,$ (14)
@@
 $\lim_{u^{k}\to 0}\frac{f(x^{*}+u_{\mathcal{J}^{*}}^{k})-f(x^{*})-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\lim_{u^{k}\to 0}\eta^{k}-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=0,$
-implying $\limsup\limits_{k\to\infty}\eta^{k}=\limsup\limits_{k\to\infty}\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\|\nabla f(x^{*})\|_{\infty}$.
-By (7), $\limsup\limits_{k\to\infty}\eta^{k}=\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$.
+implying $\limsup\limits_{k\to\infty}\eta^{k}=\limsup\limits_{k\to\infty}\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}\leq\|\nabla f(x^{*})\|_{\infty}$.
+By (7), $\limsup\limits_{k\to\infty}\eta^{k}\leq\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$.
 Therefore, $\eta^{k}<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$.
@@
 In smooth settings, the strict saddle can be defined as the points with negative minimum eigenvalue of the Hessian matrix, which indicates a descent direction.
-For our case, Theorem 4 implies a saddle point can be defined using the nonzeros.
+For our case, Theorem 4 implies that a saddle point can be defined using nonzeros.
 Definition 3.
 A stationary point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$.
-If all stationary points of $F(x)$ are either strict saddle points or local minimums, $F$ satisfies the strict saddle property.
+If all stationary points of $F(x)$ are strict saddle points or local minimums, $F$ satisfies the strict saddle property.
 This definition is related to the “active strict saddle point” proposed in [23], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points.
 To address this, they define the active strict saddle point for nonsmooth problems as the saddle point on the active manifold containing the stationary point on which the function is partly smooth.
-Our definition also belongs to this kind, as $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold containing $x^{*}$.
-To show that the strict saddle property is mathematical generic, many works show that this property holds true for a full measure of linear perturbation [23, Theorem 2.9], [44, Theorem 2].
+Our definition also belongs to this kind, since $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold that contains $x^{*}$.
+To show that the strict saddle property is mathematically generic, many works show that this property holds for a full measure of linear perturbation [23, Theorem 2.9], [44, Theorem 2].
 Namely, the perturbed function has the strict saddle property for almost all $v\in\mathbb{R}^{n}$.
@@
 $\min\limits_{x\in\mathbb{R}^{n}}F_{v}(x):=F(x)-\langle v,x\rangle=f(x)+\lambda\psi(x)-\langle v,x\rangle.$ (15)
-To show the desired result, we need the assistance of the famous Sard’s theorem. Our proof of the strict saddle property mainly follows the proof of [45, Lemma 2.7] which is proved firstly in [46].
+To show the desired result, we need the assistance of the famous Sard’s theorem. Our proof of the strict saddle property mainly follows the proof of [45, Lemma 2.7], which is first proved in [46].
 Definition 4.
@@
 Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$ with $\mathcal{I}=\mathcal{I}(s)$.
-The Jacobian of $v$ at $x$ is then given by $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$.
+The Jacobian of $v$ at $x$ is then given by $Dv(x)=\begin{bmatrix}\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)&0\\ 0&I\end{bmatrix}$, and $Dv(x)$ has full rank if and only if $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$ has full rank.
 By Sard’s Theorem,
-$\{[\nabla_{\mathcal{I}}F(x);v_{\mathcal{I}^{c}}]\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$
-$\implies$ $\{v\mid v=v(x),\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$
-$\implies$ $\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}\}\text{ has measure 0 in }\mathbb{R}^{n}.$
+$\{v(x)\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n},$
+which is equivalent to
+$\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}(s)|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}.$
 Now we consider all the possible sign pattern vectors $s\in\{+1,-1,0\}^{n}$. There are at most $3^{n}$ different patterns. Therefore,
@@
 Definition 5.
-(Unstable fixed point) Let $x^{*}$ be a fixed point of a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
+(Unstable fixed point) Let $x^{*}$ be a fixed point on a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
 $x^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
 The Center Stable Manifold (CSM) Theorem plays a crucial role in deriving the avoidance of strict saddle points.
-The CSM theorem establishes a favorable property for the fixed point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, which converge to the fixed point with at least one eigenvalue of $DT$ larger than $1$ in magnitude, has Lebesgue measure zero.
+The CSM theorem establishes a favorable property for the fixed-point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, which converge to the fixed point with at least one eigenvalue of $DT$ larger than $1$ in magnitude, has Lebesgue measure zero.
 Theorem 7.
@@
 The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero.
-This requires the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero.
-Such result was established in [47, Theorem 2] for a global diffeomorphism map and then for a globally lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [23].
-Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown by using the fact that $\mathbb{R}^{n}$ is second countable no matter the number of stationary points are countable (see [23, Corollary 2.12] for a standard proof).
+This requires that the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero.
+This result was established in [47, Theorem 2] for a global diffeomorphism map and then for a global lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [23].
+Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown using the fact that $\mathbb{R}^{n}$ is second countable no matter if the number of stationary points is countable (see [23, Corollary 2.12] for a standard proof).
 We list their results below.
@@
 [47, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$.
-Let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ Then the set of initial points that converge to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
+Let $\mathcal{U}_{T}$ consist of all unstable fixed points $x$ of $T$ Then the set of initial points that converge to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
 Corollary 1.
@@
 To summarize, the analysis of escaping from strict saddle points under the strict saddle property generally consists of three steps:
-  1. (1).
+  1. (1)
     Interpret the iterative algorithm as a fixed point iteration, and build the equivalence of the unstable fixed points and the strict saddle points.
-  2. (2).
-    Show the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at fixed point), so that the CSM theorem guarantees the avoidance of converging to the strict saddle point.
-  3. (3a).
-    Show the fixed point iteration is globally a diffeomorphism.
-  4. (3b).
-    (Alternative to (3a)) Show the fixed point iteration is globally a lipeomorphism.
+  2. (2)
+    Show that the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at the fixed point), so that the CSM theorem guarantees the avoidance of converging to the strict saddle point.
+  3. (3a)
+    Show that the fixed-point iteration is globally a diffeomorphism.
+  4. (3b)
+    (Alternative to (3a)) Show that the fixed-point iteration is globally a lipeomorphism.
 3 Damped Iterative Reweighted $\ell_{1}$ Algorithms
@@
 where $w(x_{i})=r^{\prime}(|x_{i}|)$.
-To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a Lipeomorphism globally.
-However, this is not true in our case since $S$ is generally non-invertible.
-This algorithm may quickly converges to a local minimizers due to the rapid growth of the weight values, especially for the case when $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered.
+This algorithm may quickly converge to local minimizers due to the rapid growth of the weight values, especially for the case where $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered.
 To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
@@
 $S^{x}_{i}(x^{k},\epsilon^{k})=\begin{cases}x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ 0,&\text{otherwise }.\end{cases}$ (19)
-$S$ is not a bijection, as it maps multiple values to 00. Motivated by the technique used in [23], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
+To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a Lipeomorphism globally.
+However, this is not true in our case since $S$ is generally noninvertible.
+In particular, $S$ is not a bijection, as it maps multiple values to 00.
+Motivated by the technique used in [23], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
 $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S^{x}(x^{k},\epsilon^{k})\\ (1-\alpha(1-\mu))\epsilon^{k}\end{bmatrix}$ (20)
@@
 Therefore, $S_{i}^{x}(x^{*},0)=0=x_{i}^{*}$.
-Overall, we have shown that $x^{*}=S^{x}(x^{*},0)$ in all cases.
+In general, we have shown that $x^{*}=S^{x}(x^{*},0)$ in all cases.
 Conversely, suppose $x^{*}=S^{x}(x^{*},0)$.
@@
 Hence, $\nabla_{i}f(x^{*})\in[-\lambda r^{\prime}(|0^{+}|),-\lambda r^{\prime}(|0^{+}|)]$.
-Overall, $x^{*}$ satisfies (5) and is a stationary point of (P).
+In general, $x^{*}$ satisfies (5) and is a stationary point of (P).
 By Assumption 2, it satisfies (6) and (7).
@@
 3.1 Convergence analysis
-It is noteworthy that the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.
+It should be noted that the same conclusion for Algorithm 10 can be readily deduced by setting $\epsilon$ to zero. We first make the following assumptions for the algorithm.
 Assumption 3.
@@
 Therefore, we know $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above and this is also the case for $r^{\prime}(|0^{+}|)<\infty$.
-Overall, we can conclude
-$\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$ (20)
-on all stationary points in $\mathcal{L}(F^{0})$.
+Overall, we can conclude that for any stationary point $x^{*}\in\mathcal{L}(F^{0})$, there exists $\rho$ such that
+$\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho.$ (22)
 We first show the monotonicity of $F(x,\epsilon)$. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
@@
 Proposition 10.
-Suppose Assumption 1-3 hold.
+Suppose Assumptions 1-3 hold.
 Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$.
-It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and the reduction satisfies
+It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and that the reduction satisfies
 $F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
@@
 Proof.
-Since the concavity of $r(\cdot)$ on $\mathbb{R}_{++}$ gives $r(x_{1})\leq r(x_{2})+r^{\prime}(x_{2})(x_{1}-x_{2})$ for any $x_{1},x_{2}\in\mathbb{R}_{++}$, we can imply that for $i=1,...,n$,
+Since the concavity of $r(\cdot)$ on $\mathbb{R}_{++}$ gives $r(x_{1})\leq r(x_{2})+r^{\prime}(x_{2})(x_{1}-x_{2})$ for any $x_{1},x_{2}\in\mathbb{R}_{++}$, we have that for $i=1,...,n$,
 $r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+\epsilon_{i}^{k})+r^{\prime}(|x_{i}^{k}|+\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
@@
 $=$ $\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2},$
-where the first inequality is by the fact that $\epsilon^{k+1}\leq\epsilon^{k}$.
+where the first inequality is due to the fact that $\epsilon^{k+1}\leq\epsilon^{k}$.
 For each subproblem, $y^{k}$ satisfies the optimal condition
@@
 $(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
-which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$.
-Moreover, it follows from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (21) holds for any $k\in\mathbb{N}$.
-∎
+which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎
 With the monotonicity of $F$, we derive the global convergence of the ${\rm DIRL}_{1}$ method.
@@
 Proof.
-Firstly, boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$.
+Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$.
 Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$.
@@
 The map $S$ has the property of so-called “active manifold identification”, which means the iterates fall onto a smooth manifold after finite iterations.
-The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consists of the nonzeros at a limit point.
+The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [33, 32] and the manifold in this case is simply the subspace consisting of nonzeros at a limit point.
 Theorem 12.
@@
   2. (ii)
-    $S$ is $C^{1}$-smooth in $B$ and $S^{x}(x,\epsilon)=x-\frac{1}{\beta}g(x,\epsilon)$ with
+    $S$ is $C^{1}$ smooth in $B$ and $S^{x}(x,\epsilon)=x-\frac{1}{\beta}g(x,\epsilon)$ with
     $g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
@@
 Proof.
-(i) Since $S$ is a continuous map and $x^{*}=S^{x}(x^{*},\epsilon^{*})$, it holds that
+(i) Since $S$ is a continuous map and $x^{*}=S^{x}(x^{*},\epsilon^{*})$, we have
 $\text{sign}(x_{i})=\text{sign}(x^{*}_{i})\text{\ and\ }\text{sign}(S_{i}(x,\epsilon))=\text{sign}(S_{i}(x^{*},0))=\text{sign}(x^{*}_{i}),\quad i\in\mathcal{I}^{*}$
@@
 Therefore, there exists a neighborhood $B$ such that $|x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)|<\tfrac{1}{\beta}\lambda w(x_{i},\epsilon_{i})$, $\forall(x,\epsilon)\in B$, which is equivalent to $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))<0$ and $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))>0$, meaning $S_{i}(x,\epsilon)=0$.
-(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agree with following constrain problem
+(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agrees with following constrained problem
 $S^{x}(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (29)
@@
 $g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
-Therefore, $S$ is $C^{1}$-smooth in $B$
-(iii) It follows from (27) that the iteration $T$ is $C^{1}$-smooth in $B$ for any for any $(x,\epsilon)\in B$. ∎
-The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent for the variables in $\mathcal{I}^{*}$ locally around a stationary point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$.
+Therefore, $S$ is $C^{1}$ smooth in $B$.
+(iii) It follows from (29) that the iteration $T$ is $C^{1}$ smooth in $B$. ∎
+The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent method for the variables in $\mathcal{I}^{*}$ locally around a stationary point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$.
 Specifically, it can be expressed as
@@
 We then proceed to the saddle escape mechanism for DIRL1, which ensures that the algorithm does not converge to the strict saddle points defined in Definition 3 and, instead, converges to local minima.
-To apply the CSM theorem, we have to show the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
+To apply the CSM theorem, we have to show that the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
 Theorem 13.
@@
 ∎
-To apply Corollary 1, we next to show $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$.
+To apply Corollary 1, we next show that $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$.
 This is equivalent to show $T$ and its inverse $T^{-1}$ is Lipschitz continuous.
@@
 (Lipeomorphism) Suppose Assumption 1-3 hold.
-Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ defined in Proposition 10, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$.
-Then $T$ is invertible, and $T$ and $T^{-1}$ is Lipschitz continuous (i.e., $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
+Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ is defined in Proposition 10, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$.
+Then $T$ is invertible; moreover, $T$ and $T^{-1}$ is Lipschitz continuous (that is, $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
 Proof.
@@
 If $r^{\prime}(|0+|)=\infty$, by Proposition 10, $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$.
-When $r^{\prime}(|x_{i}|)\geq w(x_{i},\epsilon_{i})>C$, then it must be true that $S_{i}(x,\epsilon)=0$.
+When $r^{\prime}(|x_{i}|)\geq w(x_{i},\epsilon_{i})>C$, it must be true that $S_{i}(x,\epsilon)=0$.
 Therefore, for any $S_{i}(x)=0$ with $x\in\mathcal{L}(F^{0})$,
 $r^{\prime}(|x_{i}|)\leq C\text{ \ and \ }|x_{i}|\geq\underline{x}:=(r^{\prime})^{-1}(C)>0$ (31)
-by Assumption 3.
-This implies that we can equivalently write $S$ as
-$S_{i}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
+by Assumption 3. This implies that we can equivalently write $S$ as
+$S_{i}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}(x_{i},\epsilon_{i})/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
 with
@@
 A natural choice is $L_{r}=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$ and $L_{r}=r^{\prime\prime}(|\underline{x}|)$ if $r^{\prime}(|0^{+}|)=\infty$.
-Next we prove $S$ is Lipschitz continuous. It follows that
+Next we prove that $S$ is Lipschitz continuous. It follows that
 ${\|}S^{x}(\hat{x},\hat{\epsilon})-S^{x}(\tilde{x},\tilde{\epsilon}){\|}=$ $\ {\|}\mathbb{S}_{\lambda\hat{u}/\beta}(\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x}))-\mathbb{S}_{\lambda\tilde{u}/\beta}(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x})){\|}$
@@
 $\leq$ $\ (1+\tfrac{1}{\beta}L_{f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|,$
-where the first inequality is by Lemma A.3 and the second is by (19) and (30).
+where the first inequality is determined by Lemma A.3 and the second is by (21) and (32).
 Therefore, $\|\alpha(\hat{x}-S^{x}(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S^{x}(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
@@
 Theorem 15.
-(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3.
-Let the condition of Proposition 14 holds and $\alpha<\beta/\rho$.
+(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3 satisfied.
+Suppose $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$, $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$ and $\alpha<\beta/\rho$.
 Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 10.
@@
 Proof.
-It follows directly from Theorem 13 that map $T$ is local $C^{1}$-smooth around any stationary points, and all stationary points $x^{*}$ are unstable fixed points.
+It follows directly from Theorem 13 that the map $T$ is local $C^{1}$ smooth around any stationary points, and all stationary points $x^{*}$ are unstable fixed points.
 By applying Proposition 14, $T$ is a lipeomorphism.
-By Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure.
+By Corollary 1, the set of initial conditions attracted by such fixed points has a zero Lebesgue measure.
 ∎
 We illustrate this property by applying DIRL1 to Example 1.
-We test $\alpha=0.1$ and $0.9$ with different initial points; first is $(3,3)$ representing an initial point far from the global minimizer and the second is $(0,0.043)$ very close to the saddle point.
+We test $\alpha=0.1$ and $0.9$ with different initial points: $(3,3)$ representing an initial point far from the global minimizer and $(0,0.043)$ very close to the saddle point.
 The trajectories are shown in Figure 3.
@@
 (d) $\alpha=0.1$ and initial point $(0,0.043)$.
-Figure 3: Trajectories of DIL2.
+Figure 3: Trajectories of DIRL1.
 4 Damped Iterative Reweighted $\ell_{2}$ Algorithms
@@
 In this method, we write $h(\cdot)=r\circ\sqrt{(\cdot)}$, so that $s(x_{i})=h(x^{2}_{i})$.
-At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearize $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem,
+Since $\sqrt{\cdot}$ is concave on $\mathbb{R}_{+}$, $\sqrt{\lambda y_{1}+(1-\lambda)y_{2}}\geq\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}}$ for any $\lambda\in[0,1]$ and $y_{1},y_{2}\geq 0$.
+Since $r$ is increasing and concave, it follows that $h(\lambda y_{1}+(1-\lambda)y_{2})=r(\sqrt{\lambda y_{1}+(1-\lambda)y_{2}})\geq r(\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}})\geq\lambda r(\sqrt{y_{1}})+(1-\lambda)r(\sqrt{y_{2}})=\lambda h(y_{1})+(1-\lambda)h(y_{2})$, meaning $h$ is also concave on $\mathbb{R}_{+}$.
+At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearize $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem (for brievty, we still use $S$ and $T$ for the maps),
 $y^{k}=S^{x}(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2},$
@@
 $w(x_{i},0)=\begin{cases}\frac{r^{\prime}(|x_{i}|)}{2|x_{i}|}&\text{ if }x_{i}\neq 0\\ \infty&\text{ if }x_{i}=0.\end{cases}$
-The subproblem solution is given by $y^{k}=S^{x}(x^{k},\epsilon^{k})$, where the operator $S$ is defined as
-$S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}w(x_{i},\epsilon_{i})}(x_{i}-\frac{1}{\beta}\nabla_{i}f(x)).$ (31)
-If $w(x_{i},\epsilon_{i})=\infty$, we simply set $S_{i}^{x}(x,\epsilon)=0$. The damped iteratively reweighted $\ell_{2}$ algorithm, hereinafter named DIRL2 algorithm, is presented in Algorithm 10.
+The solution of the subproblem is given by
+$y^{k}_{i}=S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}w(x_{i}^{k},\epsilon_{i}^{k})}(x_{i}^{k}-\frac{1}{\beta}\nabla_{i}f(x^{k})).$ (33)
+If $w(x_{i}^{k},\epsilon_{i}^{k})=\infty$, we simply set $S_{i}^{x}(x^{k},\epsilon^{k})=0$.
+The damped iteratively reweighted $\ell_{2}$ algorithm, hereinafter named DIRL2 algorithm, is presented in Algorithm 10.
 1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
@@
 Algorithm 2 Damped Iteratively Reweighted $\ell_{2}$ Algorithm (${\rm DIRL}_{2}$)
-The fixed point iteration of DIRL2 is then given as $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=(1-\alpha)I+\alpha S$, where
+The fixed point iteration of DIRL2 is then given as $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})$ with $T=(1-\alpha)I+\alpha S$ and
 $T(x,\epsilon):=\begin{bmatrix}(1-\alpha)x+\alpha S^{x}(x,\epsilon)\\ (1-\alpha(1-\mu))\epsilon\end{bmatrix}.$ (34)
@@
 As for $i\in\mathcal{J}^{*}$, $w(0,0)=\infty$, implying that $0=S_{i}(x^{*},0)$.
-To summarize, $x^{*}=S^{x}(x^{*},0)$.
+In summary, $x^{*}=S^{x}(x^{*},0)$.
 Now consider $x^{*}=S^{x}(x^{*},0)$.
@@
 The converse statement of Proposition 16 may not be true when $\mathcal{J}^{*}\neq\emptyset$ and $r^{\prime}(|0^{+}|)<\infty$.
-Since for any $x^{*}$ satisfying $x_{i}^{*}=S_{i}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it holds naturally true that $w(0,0)=\infty$ and $S_{i}(x^{*},0)=0$.
+Since for any $x^{*}$ that satisfies $x_{i}^{*}=S_{i}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it is naturally true that $w(0,0)=\infty$ and $S_{i}(x^{*},0)=0$.
 However, there is no guarantee that $x^{*}$ satisfies (7).
 In those cases, $S$ may have two types of fixed points — those are stationary points and others are not.
-This represents a stark difference from the DIRL1 algorithm, where all the fixed points of the iteration map are stationary points.
+This represents a stark difference from the DIRL1 algorithm, where all fixed points of the iteration map are stationary points.
 4.1 Convergence analysis
-In this subsection, we show that DIRL2 only converge to its fixed points which are the stationary points of (P).
-Notice that this is true in all cases even though it may have fixed points that are not the stationary points, as discussed in the previous subsection.
-Similar to the analysis of DIRL1 algorithm, the analysis of DIRL2 is also based on the monotonic decrease of the objective $F(x,\epsilon)$.
-We also need Assumption 3 and consequently (19) hold true for the analysis.
-However, it must be noticed that the definition of $F(x,\epsilon)$ is different from that for DIRL1 algorithm.
+In this subsection, we show that DIRL2 only converges to its fixed points which are the stationary points of (P) under the assumption that $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$.
+Similarly to the analysis of the DIRL1 algorithm, the analysis of DIRL2 is also based on the monotonic decrease in the objective $F(x,\epsilon)$.
+We also need Assumption 3 and consequently (21) to hold for the analysis.
+However, it should be noted that the definition of $F(x,\epsilon)$ is different from that of the DIRL1 algorithm.
 Proposition 17.
@@
 Proof.
-Since $h(\cdot)=r\circ\sqrt{(\cdot)}$ is also concave by Assumption 1, it follows that for $i=1,...,n$,
+It follows from the concavity of $h$ that for $i=1,...,n$,
 $h(x_{i}^{k+1})^{2}+(\epsilon_{i}^{k})^{2})\leq h(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2})+h^{\prime}((x_{i}^{k})^{2}+(\epsilon^{k}_{i})^{2})((x_{i}^{k+1})^{2}-(x_{i}^{k})^{2}),$
@@
 $=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2},$
-where the first inequality is by Lemma A.2 and the second is by the convexity of $(\cdot)^{2}$. Combine (24), (35) and get
+where the first inequality is by Lemma A.2 and the second is by the convexity of $(\cdot)^{2}$. Combining (26) and (37), we get
 $\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
@@
 $(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
-which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$.
-Moreover, it follows from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (21) holds for any $k\in\mathbb{N}$.
-∎
+which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$, there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎
 With the monotonicity of $F$, we derive the global convergence of the ${\rm DIRL}_{1}$ method.
 Theorem 18.
-(Global convergence) Suppose Assumption 1 and 3 hold.
+(Global convergence) Suppose Assumption 13 hold, and $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$.
 Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{2}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$.
-Then $\mu^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
+Then any $x^{*}\in\Omega$ is first-order optimal for (P).
 Proof.
-Firstly, boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$.
+Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$.
 Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$.
 The optimality condition of the $k$th subproblem implies
-$-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}y_{i}^{k}.$ (36)
-As $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$, $y_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$, $\lim\limits_{k\in\mathcal{S}}\frac{y_{i}^{k}}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}=\text{sign}(x_{i}^{*})$ for $i\in\mathcal{I}^{*}$, and $\limsup\limits_{k\in\mathcal{S}}\frac{|y_{i}^{k}|}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}\leq 1$ for $i\in\mathcal{J}^{*}$.
-Therefore, taking $k\to\infty,k\in\mathcal{S}$ in both sides of (36), we have for $i\in\mathcal{I}^{*}$
-$\nabla_{i}f(x^{*})+2\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*})=0,$
-satisfying (6). As for $i\in\mathcal{J}^{*}$ and $r^{\prime}(|0^{+}|)<\infty$,
-$|\nabla_{i}f(x^{*})|=\lim_{k\in\mathcal{S}\atop k\to\infty}|\nabla_{i}f(x^{k})|\leq\lim_{k\in\mathcal{S}\atop k\to\infty}|\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}y_{i}^{k}|=\limsup\limits_{k\in\mathcal{S}\atop k\to\infty}\frac{r^{\prime}(|0^{+}|)|y_{i}^{k}|}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}\leq r^{\prime}(|0^{+}|),$
-satisfying (7) by Assumption 2. Overall, $x^{*}$ is stationary. ∎
+$-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})}{\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}}}y_{i}^{k}.$ (38)
+As $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$, $y_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$, $\lim\limits_{k\in\mathcal{S}}\frac{y_{i}^{k}}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}=\text{sign}(x_{i}^{*})$ for $i\in\mathcal{I}^{*}$.
+For $i\in\mathcal{J}^{*}$, there is nothing to prove since $|\nabla_{i}f(x^{*})|\leq r^{\prime}(|0^{+}|)$ naturally holds. In general, $x^{*}$ is stationary. ∎
 4.2 Avoidance of strictly saddle points
-In this subsection, we further show that DIRL2 algorithm does not converge to fixed points that are stationary points and strict saddle points.
+In this subsection, we further show that the DIRL2 algorithm does not converge to fixed points that are stationary points and strict saddle points.
 Unlike DIRL1, the map $S$ in DIRL2 does not enjoy the property of “active manifold identification” or “model identification”.
@@
 The following additional assumption is needed to guarantee the local smoothness.
-One can verify that most sparse non-convex penalty in Table 1 like LPN, TAN, FRA, EXP and LOG are satisfied.
+One can verify that most sparse nonconvex penalty in Table 1 like LPN, TAN, FRA, EXP and LOG are satisfied.
 Assumption 4.
-$\lim\limits_{z\rightarrow 0}\frac{r^{\prime}(|z|)}{|z|}=\infty\quad\text{ and }\quad\lim\limits_{z\rightarrow 0}z\frac{r^{\prime\prime}(|z|)}{r^{\prime}(|z|)^{2}}\to 0.$
+$\lim\limits_{z\rightarrow 0}r^{\prime}(|z|)=\infty\ \text{ and }\ \lim\limits_{z\rightarrow 0}z\frac{r^{\prime\prime}(|z|)}{r^{\prime}(|z|)^{2}}\to 0.$
 To show the smoothness of $S$, we first calculate the Jacobian of $S$. Define
 $g(z):=\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|z|)}{|z|}}\ \text{ with }g(0):=0.$
-Then $S_{i}^{x}(x,\epsilon)=g(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ with $z=\sqrt{x^{2}+\epsilon^{2}}$. We first have the following result about $g$, $S$ and $T$.
-Theorem 19.
-(Local smoothness and Lipschitz continuity ) Suppose Assumption 1-4 hold. The following statements hold:
-  1. (i)
-    $g(z)$ is continuously differentiable on $\mathbb{R}$, so that it is Lipschitz continuous on $\mathcal{L}(F^{0})$.
-  2. (ii)
-    $S$ is continuously differentiable and Lipschitz continuous on $\mathcal{L}(F^{0})$ with constant $L_{S}$.
-  3. (iii)
-    $T$ is a lipeomorphism for $\alpha\in(0,\frac{1}{1+L_{S}})$.
-Proof.
-(i) It follows that
+We can write $S_{i}^{x}(x,\epsilon)=g(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ with $z=\sqrt{x^{2}+\epsilon^{2}}$. It follows that
 $g^{\prime}(z)=\begin{cases}-\tfrac{\lambda}{\beta}\frac{r^{\prime\prime}(|z|)z-\text{sign}(z)r^{\prime}(|z|)}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}},&\text{\ for }z\neq 0,\\ \lim\limits_{\delta\to 0}\left(\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|\delta|)}{|\delta|}}-0\right)\frac{1}{\delta}=0,&\text{\ for }z=0.\end{cases}$ (39)
@@
 $=$ $-\lim\limits_{z\to 0}\left(\tfrac{\beta}{\lambda}\frac{r^{\prime\prime}(|z|)z}{{r^{\prime}(|z|)}^{2}}-\text{sign}(z)\frac{|z|}{r^{\prime}(|z|)}\right)=0=g^{\prime}(0)$
-by Assumption 4. Therefore, $g$ is continuously differentiable.
-(ii) It is only needed to show the smoothness of $S$ when $\mathcal{J}^{*}\neq\emptyset$.
-We simply calculate its partial derivative and verify it is continuous at any point $(x^{*},0)\in\mathbb{R}^{2n}$.
-First, consider $S_{i},i\in\mathcal{J}^{*}$. It follows from $S_{i}^{x}(x^{*},0)=0$ and $r^{\prime}(|0^{+}|)=\infty$ that
-$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$
+by Assumption 4. Therefore, $g$ is continuously differentiable in $\mathbb{R}$.
+We have the following result for $S$ and $T$.
+Theorem 19.
+(Local smoothness and Lipschitz continuity) Suppose Assumption 1-4 hold. The following statements hold:
+  1. (i)
+    $S$ is $C^{1}$ smooth and Lipschitz continuous on $\mathcal{L}(F^{0})$ with constant $L_{S}$; moreover
+    $DS(x^{*},\bm{0})=\ \begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&,\bm{0}\\ 0&0&0\\ 0&0&\mu I\end{bmatrix},$ (40)
+    where $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}W_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$.
+  2. (ii)
+    $T$ is a lipeomorphism for $\alpha\in(0,\frac{1}{1+L_{S}})$ with
+    $DT(x^{*},\bm{0})=\ \begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&\bm{0}\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}.$
+Proof.
+(i) We simply calculate its partial derivative and verify that it is continuous at any point $(x^{*},\epsilon^{*})\in\mathbb{R}^{2n}$.
+Consider $i$ satisfying $z_{i}^{*}:=\sqrt{(x_{i}^{*})^{2}+(\epsilon_{i}^{*})^{2}}>0$. In this case, $S$ is smooth with respect to $x_{i}$ and $\epsilon_{i}$, and
+$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ g(z^{*}_{i})(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))+\tfrac{x_{i}^{*}}{z_{i}^{*}}g^{\prime}(z_{i}^{*})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (41)
+$\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ -g(z^{*}_{i})\tfrac{1}{\beta}\nabla^{2}_{ij}f(x^{*}),\quad j\neq i;$ (42)
+$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ \tfrac{\epsilon_{i}^{*}}{z_{i}^{*}}g^{\prime}(z^{*}_{i})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (43)
+$\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ 0,\quad j\neq i.$ (44)
+Consider $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$. It follows from $S_{i}^{x}(x^{*},\bm{0})=0$ and $r^{\prime}(|0^{+}|)=\infty$ that
+$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$ (45)
 $=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta$
-$=$ $-\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}g^{\prime}(0)=0.$
-As for $j\neq i$, it follows that $\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\epsilon^{*})=0$ and $\frac{\partial}{\partial\epsilon_{j}}S_{i}(x^{*},\epsilon^{*})=0$.
-Generally, for any $i$ satisfying $z_{i}:=\sqrt{x_{i}^{2}+\delta_{i}^{2}}>0$, $S$ is smooth with respect to $x_{i}$ and $\epsilon_{i}$ with $z_{i}=\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}\neq 0$.
-$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=$ $\ g(z)(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x))+\tfrac{x_{i}}{z_{i}}g^{\prime}(z_{i})(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ (38)
-$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=$ $\ \tfrac{\epsilon_{i}}{z_{i}}g^{\prime}(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$ (39)
-$\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=$ $\ -g(z)\tfrac{1}{\beta}\nabla^{2}_{ij}f(x),\quad j\neq i$ (40)
-$\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=$ $\ 0,\quad j\neq i.$ (41)
-Now for any $i\in\mathcal{J}^{*}$, as $z_{i}\to 0$, it follows from $\frac{x_{i}}{z_{i}}\leq 1$ and $zr^{\prime\prime}(z)/(r^{\prime}(z))^{2}\to 0$ that
-$\lim_{z_{i}\to 0}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}),\ \lim_{z_{i}\to 0}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})$
-$\lim_{z_{i}\to 0}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}),\ \lim_{z_{i}\to 0}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})$
-Overall, we have shown that $S$ is continuously differentiable. The Lipschitz continuity of $S$ then trivially comes from (i) and the boundedness of $\mathcal{L}(F^{0})$.
-(iii) This is straightforward by applying [23, Lemma 2.13 and 2.14]. ∎
+$=$ $-\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}g^{\prime}(0)=0,$
+$\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{j})]/\delta$ (46)
+$=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta=0,\quad j\neq i;$
+$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(0,\delta)}[-\tfrac{1}{\beta}\nabla_{i}f(x^{*})]/\delta$ (47)
+$=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta=0;$
+[…]
+Now we verify that the Jacobian is continuous at $(x^{*},\epsilon^{*})$.
+It suffices to show this for any $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$; indeed, it holds that It follows from (41), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g(0)=0,g^{\prime}(0)=0$ that
+$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}).$
+It follows from (42) and $g(0)=0$ that
+$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*}).$
+It follows from (43), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g^{\prime}(0)=0$ that
+$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
+It follows from (44) that
+$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
+Overall, we have shown that $S$ is continuously differentiable.
+To derive the matrix form of the Jacobian, it follows from (41) that for $i\in\mathcal{I}^{*}$,
+$h_{ii}:=$ $\ \frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})$
+$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(x^{*}_{i}))\tfrac{\lambda}{\beta}\left(|x^{*}_{i}|r^{\prime\prime}(|x^{*}_{i}|)-r^{\prime}(|x^{*}_{i}|)\right)}{\left(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(|x^{*}_{i}|)\right)^{2}},$
+$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{\tfrac{\lambda}{\beta}\left(r^{\prime\prime}(|x^{*}_{i}|)-\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}\right)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x_{i}^{*}|}}$
+$=$ $\ \frac{1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*})-\tfrac{\lambda}{\beta}r^{\prime\prime}(|x_{i}^{*}|)+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
+$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x_{i}^{*}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
+It follows from (42) that for $j\neq i$,
+$h_{ij}:=\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}.$
+We then combine this with (43)–(48) to derive (40) of the Jacobian $DS(x^{*},\bm{0})$. The Lipschitz continuity of $S$ trivially leads to the boundedness of $\mathcal{L}(F^{0})$.
+(ii) This is straightforward by applying [23, Lemma 2.13 and 2.14]. The Jacobian $DT(x^{*},\bm{0})$ is obtained directly by (i). ∎
 Now we show that every strict saddle point of (P) is an unstable fixed point of $T$.
 Theorem 20.
-(Unstable Fixed Point) Suppose $x^{*}$ is a stationary point of (P).
-If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$.
-Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},0)$ is invertible.
+(Unstable Fixed Point) Suppose Assumption 1-4 hold.
+Let $x^{*}$ be a stationary point of (P).
+If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=\bm{0}$ is an unstable fixed point of $T$.
+Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible.
 Proof.
-Consider the Jacobian of $S$ at a fixed point $(x^{*},0)$. Define $H\in\mathbb{R}^{|\mathcal{I}^{*}|\times n}$ as follows:
-$\forall i\in\mathcal{I}^{*},\ h_{ii}=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{(|x_{i}^{*}|+\tfrac{\lambda}{\beta}r^{\prime}(x^{*}_{i}))\tfrac{\lambda}{\beta}\left(|x^{*}_{i}|r^{\prime\prime}(|x^{*}_{i}|)-r^{\prime}(|x^{*}_{i}|)\right)}{\left(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(|x^{*}_{i}|)\right)^{2}},$
-$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{\tfrac{\lambda}{\beta}\left(r^{\prime\prime}(|x^{*}_{i}|)-\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}\right)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
-$=$ $\ \frac{1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*})-\tfrac{\lambda}{\beta}r^{\prime\prime}(|x^{*}_{i}|)+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
-$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x^{*}_{i}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
-$\forall j\neq i,\ h_{ij}:=$ $\ \frac{\partial}{\partial x_{j}}S_{i}(x^{*},0)=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{|x_{i}^{*}|}},$
-by (38) and (40).
-Therefore, $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$,
-$DS(x^{*},0)=$ $\begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&0\\ 0&0&0\\ 0&0&\mu I\end{bmatrix}$
-$DT(x^{*},0)=$ $\begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&0\\ 0&(1-\alpha)I&0\\ 0&0&(1-\alpha(1-\mu))I\end{bmatrix}.$
-Hence, the eigenvalues of $DS(x^{*},0)$ are the same as
+It follows from Theorem 19(ii) that we only have to check the eigenvalues of
 $(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}),$ (49)
-where $W^{*}:=W(x^{*},0)$ and $W^{*}_{ii}<\infty,i\in\mathcal{I}^{*}$.
+where we have $W^{*}_{ii}=w(x_{i}^{*},0)<\infty,i\in\mathcal{I}^{*}$.
 If $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$, it follows from Lemma [49, Proposition 6.1] that $\lambda_{\min}((I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$.
-Hence, $DT(x^{*},0)$ has an eigenvalue greater than 1, meaning $(x^{*},0)$ is the unstable fixed point of $T$.
-Moreover, since $\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$, it is clear that for $\alpha<\beta/\rho$, $DT(x^{*},0)$ is invertible.
+Hence, $DT(x^{*},\bm{0})$ has an eigenvalue greater than 1, meaning $(x^{*},\bm{0})$ is the unstable fixed point of $T$.
+Moreover, since $\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$, it is clear that for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible.
 ∎
-Similarly, we can derive the global escape property of ${\rm DIRL}_{2}$ algorithm.
+Now we can derive the global escape property of ${\rm DIRL}_{2}$ algorithm.
 Theorem 21.
@@
 Proof.
-It follows directly from Theorem 19 and 20 that 1) map $T$ is local $C^{1}$-smooth around any stationary points; 2) all stationary points $x^{*}$ are unstable fixed points; 3) $T$ is a lipeomorphism.
+It follows directly from Theorem 19 and 20 that: (i) map $T$ is local $C^{1}$-smooth around any stationary points; (ii) all stationary points $x^{*}$ are unstable fixed points; (iii) $T$ is a lipeomorphism.
 Then by Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure.
@@
 (d) $\alpha=0.1$ and initial point $(0,0.043)$.
-Figure 4: Trajectories of DIL2.
+Figure 4: Trajectories of DIRL2.
 Appendix
@@
 $\|\mathbb{S}_{\hat{w}}(\hat{z})-\mathbb{S}_{\tilde{w}}(\tilde{z})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$
+Proof.
 (a) If $|\hat{z}_{i}|\leq\hat{w}_{i}$ and $|\tilde{y}_{i}|\leq\tilde{w}_{i}$, then $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})=0$ and $\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=0$.

### v5

@@
 Abstract
-In this paper, we consider a class of nonconvex and nonsmooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms.
-We show the second-order optimality conditions only depend on the nonzeros of the stationary points.
-We propose two damped iterative reweighted algorithms, including the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems.
-For DIRL1, we show the reweighted $\ell_{1}$ subproblem has support identification property so that DIRL1 locally reverts to a gradient descent algorithm around a stationary point.
-For DIRL2, we show the solution map of the reweighted $\ell_{2}$ subproblem is differentiable and Lipschitz continuous everywhere.
-Therefore, the map of DIRL1 and DIRL2 and their inverse are Lipschitz continuous, and the strict saddle points are their unstable fixed points.
-By applying the stable manifold theorem, these algorithms are shown to converge only to local minimizers with random initialization when the strictly saddle point property is assumed.
+This paper considers a class of nonconvex and nonsmooth sparse optimization problems, encompassing most existing nonconvex sparsity-inducing terms.
+We show that their second-order optimality conditions depend only on the nonzeros of the stationary points.
+We propose two damped iteratively reweighted algorithms, which are the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems.
+For DIRL1, we show that the reweighted $\ell_{1}$ subproblem has the support identification property so that DIRL1 locally reverts to a gradient descent algorithm around a stationary point.
+For DIRL2, we show that the solution map of the reweighted $\ell_{2}$ subproblem is differentiable and Lipschitz continuous everywhere.
+Therefore, the solution maps of DIRL1 and DIRL2 and their inverses are Lipschitz continuous, and the strict saddle points are their unstable fixed points.
+By applying the stable manifold theorem, these algorithms starting from almost every initial point are shown to converge to local minima when the strict saddle point property is assumed.
 keywords:
-strict saddle, iterative reweighted $\ell_{1}$ algorithm, nonconvex regularization, center stable manifold theorem
+strict saddle property, iterative reweighted $\ell_{1}$ algorithm, nonconvex regularization, center stable manifold theorem
+{AMS}
+49J50, 49M20, 49M37, 65K05, 65K10, 90C26, 90C30
 1 Introduction
-In this paper, we consider the following nonsmooth regularization problem,
-$\min\limits_{x\in\mathbb{R}^{n}}F(x):=f(x)+\lambda\psi(x),$ (P)
-where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and the penalty/regularization term $\psi(x)$ takes the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$.
-The data fitting function $f$ is differentiable twice continuously and possibly nonconvex, and $r$ satisfies the following properties.
+In this paper, we consider the following nonconvex and nonsmooth regularization problem,
+(P) $\min\limits_{x\in\mathbb{R}^{n}}F(x):=f(x)+\lambda\psi(x),$
+where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a twice continuously differentiable function and $\psi(x)$ is a penalty/regularization term taking the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$ and $r$ satisfying the following properties.
 Assumption 1.
-$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$, C2 smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
+$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$ and C2 smooth on $(0,+\infty)$, $\lim\limits_{t\rightarrow 0^{+}}r^{\prime}(t)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
 This assumption indicates $r(t)>0$ for all $t>0$, but $r$ could be nonconvex or not Lipschitz continuous around 0.
-The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance.
-This makes it well-suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [1, 2, 3, 4].
-Many widely used nonconvex regularization such as the EXP penalty [38] or the LOG penalty [39], the $\ell_{p}$ quasi-norm penalty [7] satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability.
+The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance of the model.
+This makes it well suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [17, 31, 35, 47].
+Many widely used nonconvex regularizations, such as the EXP penalty [5], the LOG penalty [26], and the $\ell_{p}$ quasi-norm penalty ($10<p<$ 1) [18], satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability.
 We provide the explicit form of several commonly used nonconvex regularizations in Table 1.
-It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizer with bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) nonLipschitz continuous regularizer with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as the $\ell_{p}$ quasi-norm (LPN).
+It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizers with a bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) nonLipschitz continuous regularizers with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as the $\ell_{p}$ quasi-norm (LPN).
 Numerous studies focus on gradient-based optimization methods that are guaranteed to find first-order stationary points.
-However, the nonconvexity of $F$ can lead to multiple local minima and saddle points, which can cause algorithms to converge to stationary points rather than the (local) optimum.
-Additionally, the presence of saddle points can significantly slow down the convergence of optimization algorithms, especially gradient-based methods.
-At a saddle point, the gradient provides misleading information about the descent directions.
+However, the nonconvexity of $F$ can lead to multiple local minima and saddle points, which can cause algorithms to converge to a stationary point rather than a local minimum.
+In addition, the presence of saddle points can significantly slow the convergence of optimization algorithms, especially gradient-based methods.
+In this situation the gradient provides misleading information about the descent directions at a saddle point.
 As a result, avoiding saddle points can help optimization algorithms more effectively converge to better solutions.
 Therefore, it becomes a prevalent topic, particularly in high-dimensional optimization problems, and is also the central focus of this paper.
-In smooth optimization, many methods have been developed and analyzed to escape saddle points, including first-order methods [8, 9, 10], momentum-based methods [11], second-order optimization algorithms [12], the algorithmic frameworks for constrained nonlinear problems [13], stochastic algorithms [14, 15], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [16, 17, 18, 19].
+In smooth optimization, many methods have been developed and analyzed to escape saddle points, including first-order methods [32, 25, 39], momentum-based methods [45], second-order algorithms [33], the algorithmic frameworks for constrained nonlinear problems [30], stochastic algorithms [19, 13], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [21, 22, 27, 12].
 In smooth settings, it is widely assumed that strict saddle property holds, which means that all saddle points have at least one strictly negative eigenvalue.
@@
 At stationary points, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate avoidance of saddle points.
-Developing nonsmooth optimization algorithms capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than in smooth optimization.
-An example of this complexity is demonstrated in the analysis of the Douglas splitting algorithm for nonsmooth problems, as discussed in [22].
-Recent research [23] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point.
-Additionally, several proximal methods, including the proximal point method, the proximal gradient method, and the proximal linear method [23], as well as a gradient method incorporating random perturbation [24], have been proposed to converge to local minima in nonsmooth optimization settings.
-In these studies, the nonsmooth objective function $r(x)$ is assumed to be $\rho$ weakly convex, which means that $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex.
+Developing nonsmooth optimization algorithms being capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than that in smooth optimization.
+An example of this complexity was demonstrated in the analysis of the Douglas splitting algorithm for nonsmooth problems, as discussed in [2].
+Recent research [15] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming that $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point.
+Furthermore, several proximal algorithms with damped methods have been proposed to converge to local minima in nonsmooth optimization settings, including the proximal point algorithm, the proximal gradient algorithm, and the damped proximal linear algorithm [15].
+Furthermore, an inexact analog of a stochastically perturbed gradient method applied to the Moreau envelope has also been introduced [14].
+In these studies, the nonsmooth function $r(x)$ is assumed to be $\rho$ weakly convex, which means that $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex.
 They assumed an active manifold around the limit point and then defined the strict saddle points as those of the $f$ restricted to the manifold.
-The weakly $\rho$-convexity for $r$ plays a crucial role in constructing the C2 differentiable Moreau envelope with the same stationary points and strict saddle points in the proximal methods.
+The weak $\rho$-convexity of $r$ plays a crucial role in constructing a $C^{2}$ differentiable Moreau envelope with the same stationary points and strict saddle points in the proximal methods.
 As for the regularized problem (P), several first-order methods have been proposed.
-[25] introduces a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1.
-This method is subsequently extended to address the $\ell_{p,q}$ group norm regularized problems of the group $\ell_{p,q}$ in [26].
-Furthermore, the Difference of Convex (DC) algorithm is enlisted [27] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions.
-It is important to note, however, that the DC algorithm generally assumes $r^{\prime}(|0^{+}|)<\infty$ to accommodate the DC decomposition, which does not hold true for the $\ell_{p}$ regularized problem.
-Another avenue of research involves iteratively reweighted algorithms, which approximate the nonconvex regularizers by iteratively reweighted $\ell_{1}$ or $\ell_{2}$ regularizers.
-Initially applied to solve the unconstrained $\ell_{p}$ regularized problem in [28] without convergence analysis, this approach was later employed to address general constrained nonconvex regularization problems [29].
-Subsequent work in [30] established its global convergence and demonstrated its locally stable support property [31].
+[46] introduced a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1.
+This method was subsequently extended to address the $\ell_{p,q}$ group norm regularized problems with $0<q<1\leq p$ in [20].
+Furthermore, the difference of convex (DC) algorithm was enlisted [1] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions.
+It is important to note, however, that the DC algorithm generally assumes $r^{\prime}(|0^{+}|)<\infty$ to accommodate the DC decomposition, which does not hold for the $\ell_{p}$ regularized problem.
+Another avenue of research involves iteratively reweighted algorithms, which approximate the nonconvex regularizers by iteratively reweighted $\ell_{1}$ or $\ell_{2}$ regularization subproblems.
+Initially applied to solve the unconstrained $\ell_{p}$ regularized problem in [7] without convergence analysis, this approach was later employed to address general constrained nonconvex regularization problems in [44].
+Subsequent work in [28] established its global convergence and demonstrates its locally stable support property [43].
 Furthermore, convergence rate analysis under the Kurdyka-Łojasiewicz property [42] and its extrapolated version was proposed in [41].
-While these first-order methods primarily focus on identifying first-order stationary points, none of them guarantee the avoidance of saddle points.
+While these first-order methods primarily focus on identifying first-order stationary points, none of them guarantee the avoidance of strict saddle points.
 1.1 The main results
-In this paper, our main purpose is to develop and analyze first-order methods that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weakly convexity.
+In this paper, our main purpose is to develop and analyze first-order methods that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weak convexity.
 In particular, we focus on iteratively reweighted methods which approximate the regularization term by simpler weighted $\ell_{1}$ or $\ell_{2}$ norms [8, 10, 23, 28, 38, 44, 43, 41].
-In contrast to the proximal gradient algorithm, iteratively reweighted methods solve a sequence of weighted $\ell_{1}$ or $\ell_{2}$ regularized subproblems and are widely applicable to various nonconvex regularization problems.
-The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator.
+In contrast to the proximal gradient algorithm, the iteratively reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) solves a sequence of weighted $\ell_{1}$ subproblems, which have an explicit solution characterized by the soft-threshold operator.
+Therefore, it is widely applicable to various nonconvex regularization problems.
 This method offers a favorable local property [43], known as the model/support identification, under the following additional assumption.
 Assumption 2.
-For any stationary point $x^{*}$, $0\in{\rm rint}\partial F(x^{*})$ (where $\partial F$ denotes the regular subdifferential and ${\rm rint}$ denotes the relative interior).
-This kind of assumption is also widely made to guarantee that the existence of an active manifold around a stationary point as assumed by [23].
-At the tail end of the iteration, the sign of the iterates remain unchanged while the nonzero components are kept uniformly bounded away from 0.
+For any stationary point $x^{*}$, $0\in{\rm rint}\partial F(x^{*})$ (where $\partial F$ denotes the regular subdifferential and ${\rm rint}$ denotes the relative interior.)
+This kind of assumption is also needed to guarantee the existence of an active manifold around a stationary point as assumed by [15].
+At the tail end of the ${\rm IRL}_{1}$ iteration, the sign of the iterates remains unchanged while the nonzero components are kept uniformly bounded away from 0.
 In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{I}$ as follows after a finite number of iterations.
-Definition 1.
-(Model Subspace) The model subspace defined by $\mathcal{I}\subset\{1,\ldots,n\}$ is
-:=$\mathcal{M}_{\mathcal{I}}:=\{u\in\mathbb{R}^{n}:\forall j\in\mathcal{I}^{c},u_{j}=0\}.$
-This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{I}}$.
+Definition 1.1.
+The model subspace defined by $\mathcal{I}\subset\{1,\ldots,n\}$ is
+:=$\mathcal{M}_{\mathcal{I}}:=\{u\in\mathbb{R}^{n}:u_{j}=0,\forall j\in\mathcal{I}^{c}\}.$
+This means that every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{I}}$.
 This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) in $\mathcal{M}_{\mathcal{I}}$, although the underlying optimization problem is nonsmooth in $\mathbb{R}^{n}$.
-The iterative reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these nonconvex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the solution is trivially given.
+The iteratively reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these nonconvex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the analytic solution of each subproblem is trivially given.
 Compared to ${\rm IRL}_{1}$, ${\rm IRL}_{2}$ does not enjoy the model subspace identification/active-manifold identity property.
 In fact, it can be often witnessed that this algorithm can generate iterates with elements asymptotically approaching zero.
-To show the different behaviors of the two algorithms, we provide the following example.
-Example 1.
-Consider the $\ell_{1/2}^{1/2}$ regularization problem
-$\min_{x\in\mathbb{R}^{2}}F(x)=f(x)+\|x\|_{1/2}^{1/2},\ \text{ with }f(x)=x_{1}^{2}+(x_{2}-5/4)^{2}.$
-The global minimizer is $(0,1)$, the local minimizer111(0,0)00(0,0)( 0 , 0 ) is always a dummy local minimizer of $\ell_{p}$ regularization problems is $(0,0)$ and the saddle point minimizer is $(0,\frac{3-2\sqrt{2}}{4})\approx(0,0.043)$ which is also a local maximizer on the $x_{2}$-axis.
-(a) Contour of $F$.
-(b) Contour of $F$ around $(0,0)$.
-Let $g(x;x^{k})=\nabla f(x^{k})^{T}(x-x^{k})+\tfrac{1}{2}\|x-x^{k}\|^{2}$.
-The iterations of the iteratively reweighted algorithms are as follows [29].
-We use the ${\rm IRL}_{1}$ algorithm with iteration
-$x^{k+1}=\arg\min_{x_{1},x_{2}}g(x;x^{k})+\tfrac{1}{2}\frac{1}{(|x_{1}^{k}|+0.1^{k})^{1/2}}|x_{1}|+\tfrac{1}{2}\frac{1}{(|x_{2}^{k}|+0.1^{k})^{1/2}}|x_{2}|$
-and the ${\rm IRL}_{2}$ algorithm with iteration
-$x^{k+1}=\arg\min_{x_{1},x_{2}}g(x;x^{k})+\tfrac{1}{4}\frac{1}{((x_{1}^{k})^{2}+0.1^{k})^{3/4}}x_{1}^{2}+\tfrac{1}{4}\frac{1}{((x_{2}^{k})^{2}+0.1^{k})^{3/4}}x_{2}^{2}.$
-The trajectories of the two algorithms are depicted in Figure 2.
-(a) Trajectory of ${\rm IRL}_{1}$
-(b) Trajectory of ${\rm IRL}_{2}$
-Figure 2: Trajectories generated by iteratively reweighted methods. ${\rm IRL}_{1}$ eventually falls into the subspace (the $x_{2}$-axis), while ${\rm IRL}_{2}$ has nonzero elements for all iterates.
-The behavior of ${\rm IRL}_{2}$ brings another challenge in analyzing the avoidance of strict saddle point.
-Current analysis for nonsmooth optimization algorithms ubiquitously depends on the active-manifold identification property, which ensures the algorithm reverts to a smooth algorithm in the active manifold in a local neighbourhood of the stationary point.
-However, there is no answer how a nonsmooth first-order algorithm like ${\rm IRL}_{2}$ would behave in the local region of a strict saddle point.
-We now summarize our contribution as follows:
+The behavior of ${\rm IRL}_{2}$ brings another challenge in analyzing the avoidance of strict saddle points.
+Current analysis for nonsmooth optimization algorithms ubiquitously depends on the active-manifold identification property, which ensures that the algorithm reverts to a smooth algorithm on the active manifold in a local neighborhood of the stationary point.
+However, there is no answer to how a nonsmooth first-order algorithm like ${\rm IRL}_{2}$ would behave in the local region of a strict saddle point.
+We now summarize our contributions as follows:
   1. (i)
-    We first show the local minimal and second-order optimality of $f$ are equivalent to those of $f$ restricted on $\mathcal{M}_{\mathcal{I}}$, and define the strict saddle points as those on $\mathcal{M}_{\mathcal{I}}$.
+    We show that the local minimum and second-order optimality of $F$ are equivalent to those of $F$ restricted in $\mathcal{M}_{\mathcal{I}}$, and we define the strict saddle points of $F$ as those in $\mathcal{M}_{\mathcal{I}}$.
   2. (ii)
-    We develop damped IRL1 algorithms individually for the two types of regularizer, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning that the next iterate is the combination of a traditional IRL1 iterate and its previous iterate.
+    We develop damped IRL1 algorithms individually for the two types of regularizers, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning that the next iterate is the combination of a traditional IRL1 iterate and its previous iterate.
 The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ from becoming $\infty$ at each iteration.
-At the tail end of the iterations, we show that this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{I}$ on the subspace $\mathcal{M}_{\mathcal{I}}$ and decrease the rest components to 0 at a linear speed.
-We then apply the center stable manifold theorem and prove that the algorithm can avoid the convergence to strict saddle points.
+At the end of the iterations, we show that this algorithm reverts to a damped gradient descent algorithm in the subspace $\mathcal{M}_{\mathcal{I}}$ and reduces the rest components to 0 at a linear speed.
+We then apply the center stable manifold theorem and prove that the algorithm can avoid convergence to strict saddle points.
   3. (iii)
-    We develop damped IRL2 algorithms for nonconvex regularized problems and provide global convergence.
-Under an additional assumption which is satisfied by most common nonconvex regularizers, we show that the algorithm is $C^{1}$ smooth and nondegenerate everywhere.
-We also proved the equivalence between the unstable fixed point of the iteration map and the strict saddle points of the regularization problem.
+    We develop damped IRL2 algorithms for nonconvex and nonsmooth regularized problems and provide global convergence.
+Under an additional assumption which is satisfied by the $\ell_{p}$ regularization, we show that the algorithm is $C^{1}$ smooth and nondegenerate everywhere.
+We also prove the equivalence between a unstable fixed point of the iteration map and a strict saddle point of the nonconvex regularized problem.
 Therefore, the avoidance of strict saddle points is derived.
-This is the first algorithm shown to have this property without using the active-manifold identification property or the stochastic technique in nonsmooth optimization algorithms.
+This is the first algorithm shown to have this property without using the active-manifold identification property or the stochastic technique in nonsmooth optimization setting.
 1.2 Outline
@@
 This does not define a proper norm because of the lack of subadditivity.
-For $x\in\mathbb{R}^{n}$, $x_{i}$ denotes the $i$-th element of $x$.
-The index set of nonzero and zero components of $x$ is defined as $\mathcal{I}(x)=\{i\mid x_{i}\neq 0\}$, and $\mathcal{J}(x)=\{i\mid x_{i}=0\}$, respectively.
+The index sets of nonzero and zero components of $x$ are defined as $\mathcal{I}(x)=\{i\mid x_{i}\neq 0\}$, and $\mathcal{J}(x)=\{i\mid x_{i}=0\}$, respectively.
 Moreover, $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ denotes $\mathcal{I}(x^{*})$ and $\mathcal{J}(x^{*})$ for short.
 ${\rm diag}(x)$ is the diagonal matrix by placing elements of vector $x$ on the main diagonal.
-${\rm sign}(x):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the sign function defined by
-$[{\rm sign}(x)]_{i}=\left\{\begin{aligned} &1,\ &x_{i}>0,\\ &-1,\ &x_{i}<0,\\ &0,\ &x_{i}=0.\end{aligned}\right.$ (1)
-$DT(x)$ denote the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$.
-Given $x\in\mathbb{R}^{n}$ and $\mathcal{I}\subset\{1,\ldots,n\}$, denote $x_{\mathcal{I}^{*}}\subset\mathbb{R}^{\mathcal{I}^{*}}$ or $[x_{i}]_{i\in\mathcal{I}}$ as the subvector of the elements in $\mathcal{I}$; also, given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants.
+${\rm sign}(x):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the sign function defined by $[{\rm sign}(x)]_{i}=1$ if $x_{i}>0$, $[{\rm sign}(x)]_{i}=-1$ if $x_{i}<0$ and $[{\rm sign}(x)]_{i}=0$ if $x_{i}=0$.
+$DT(x)$ denotes the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$.
+Given $x\in\mathbb{R}^{n}$ and $\mathcal{I}\subset\{1,\ldots,n\}$, denote $x_{\mathcal{I}}\subset\mathbb{R}^{\mathcal{I}}$ or $[x_{i}]_{i\in\mathcal{I}}$ as the subvector of the elements in $\mathcal{I}$; also, given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$ with respect to $\mathcal{I}$, respectively.
 If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}f(x)=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$.
-Denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively.
-Given a symmetric matrix $H\in\mathbb{R}^{n\times n}$, let $\lambda_{\min}(H)$ be the smallest eigenvalue of $H$.
-Let $x\in\mathbb{R}^{n}$ and $w\in\mathbb{R}_{+}^{n}$, the soft-thresholding operator $\mathbb{S}_{w}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with the vector of thresholds $w$ is defined $[\mathbb{S}_{w}(x)]_{i}=\text{sign}(x_{i})\max(|x_{i}|-w_{i},0)$.
+Denote $f_{|\mathcal{I}}$ as the restriction of $f$ in $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$; therefore, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}};0_{\mathcal{J}}]).$ Let $x\in\mathbb{R}^{n}$ and $w\in\mathbb{R}_{+}^{n}$, the soft-thresholding operator $\mathbb{S}_{w}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with $w$ defined as $[\mathbb{S}_{w}(x)]_{i}=\text{sign}(x_{i})\max(|x_{i}|-w_{i},0)$.
 In particular, if $w_{i}=\infty$, it follows that $[\mathbb{S}_{w}(x)]_{i}=0$.
-Regularizer $\psi(x)$ $r^{\prime}(|x_{i}|)$ $r^{\prime\prime}(|x_{i}|)$
-EXP [38] $\sum\limits_{i=1}^{n}(1-e^{-p|x_{i}|})$ $pe^{-p|x_{i}|}$ $-\frac{1}{p^{2}}e^{-\frac{|x_{i}|}{p}}{p}$
-LOG [39] $\sum\limits_{i=1}^{n}\log(1+p|x_{i}|)$ $\frac{p}{1+p|x_{i}|}$ $-\frac{1}{(|x_{i}|+p)^{2}}$
-FRA [40] $\sum\limits_{i=1}^{n}\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{p}{(|x_{i}|+p)^{2}}$ $-\frac{2p}{(|x_{i}|+p)^{2}}$
-LPN [7] $\sum\limits_{i=1}^{n}|x_{i}|^{p}$ $p|x_{i}|^{p-1}$ $p(p-1)|x_{i}|^{p-2}$
-TAN [41] $\sum\limits_{i=1}^{n}\arctan\frac{|x_{i}|}{p}$ $\frac{p}{(|x_{i}|^{2}+p^{2})^{2}}$ $-\frac{2p|x_{i}|}{(|x_{i}|^{2}+p^{2})^{2}}$
+Regularizer $\psi(x)$ $r(|x_{i}|)$ $r^{\prime}(|x_{i}|)$ $r^{\prime\prime}(|x_{i}|)$
+EXP [5] $\sum\limits_{i=1}^{n}(1-e^{-p|x_{i}|})$ $1-e^{-p|x_{i}|}$ $pe^{-p|x_{i}|}$ $-p^{2}e^{-p|x_{i}|}$
+LOG [26] $\sum\limits_{i=1}^{n}\log(1+p|x_{i}|)$ $\log(1+p|x_{i}|)$ $\frac{p}{1+p|x_{i}|}$ $-\frac{p^{2}}{(1+p|x_{i}|)^{2}}$
+FRA [16] $\sum\limits_{i=1}^{n}\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{p}{(|x_{i}|+p)^{2}}$ $-\frac{2p}{(|x_{i}|+p)^{3}}$
+LPN [18] $\sum\limits_{i=1}^{n}|x_{i}|^{p}$ $|x_{i}|^{p}$ $p|x_{i}|^{p-1}$ $p(p-1)|x_{i}|^{p-2}$
+TAN [6] $\sum\limits_{i=1}^{n}\arctan(\frac{|x_{i}|}{p})$ $\arctan(\frac{|x_{i}|}{p})$ $\frac{p}{|x_{i}|^{2}+p^{2}}$ $-\frac{2p|x_{i}|}{(|x_{i}|^{2}+p^{2})^{2}}$
 Table 1: Concrete examples of nonconvex regularization
@@
 2.1 Optimality conditions
-In this subsection, we introduce the first and the second-order optimality condition of problem (P). First, let us recall the definition of subgradient and the first-order optimality condition.
-Definition 2.
+We discuss the first- and second-order optimality conditions of problem (P) in this subsection. First, let us recall the definition of subgradients and the first-order optimality condition.
+Definition 2.1.
 […]
   1. (1)
-    $v$ is a regular subgradient of $f$ at $\bar{x}$, written $v\in\hat{\partial}f(\bar{x})$ if
-    ^$\hat{\partial}f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$ (2)
+    $v$ is a regular subgradient of $f$ at $\bar{x}$, if $v\in\hat{\partial}f(\bar{x})$ where
+    (1) ^$\hat{\partial}f(\bar{x})=\{v\mid f(\bar{x})+\langle v,x-\bar{x}\rangle+o(\|x-\bar{x}\|)\leq f(x),\ {\rm as}\ x\rightarrow\bar{x}\}.$
   2. (2)
-    $v$ is a (general) subgradient of $f$ at $\bar{x}$, written $v\in\partial f(\bar{x})$, if there are sequences $x^{v}\rightarrow\bar{x}$ and $v^{v}\in\hat{\partial}f(x^{v})$ with $x^{v}\rightarrow v$.
+[…]
 It is trivial to see $s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$. The subgradients of $s$ are given in the following proposition.
-Proposition 2.
+Proposition 2.2.
 It holds that $\partial F(x)=\hat{\partial}F(x)=\nabla f(x)+\lambda\partial\psi(x)$ where
-$\partial\psi(x)=\hat{\partial}\psi(x)=\partial\sum\limits_{i=1}^{n}s(x_{i})=\partial s(x_{1})\times...\times\partial s(x_{n})$ (3)
+(2) $\partial\psi(x)=\hat{\partial}\psi(x)=\partial\sum\limits_{i=1}^{n}s(x_{i})=\partial s(x_{1})\times...\times\partial s(x_{n})$
 and
-$\partial s(x_{i})=\hat{\partial}s(x_{i})=\begin{cases}\{\text{sign}(x_{i})r^{\prime}(|x_{i}|)\},&i\in\mathcal{I}(x),\\ [-r^{\prime}(|0^{+}|),+r^{\prime}(|0^{+}|)],&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)<\infty,\\ \mathbb{R},&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)=\infty.\end{cases}$ (4)
-Proof.
+(3) $\partial s(x_{i})=\hat{\partial}s(x_{i})=\begin{cases}\{\text{sign}(x_{i})r^{\prime}(|x_{i}|)\},&i\in\mathcal{I}(x),\\ [-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)],&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)<\infty,\\ \mathbb{R},&i\in\mathcal{J}(x)\text{ and }r^{\prime}(|0^{+}|)=\infty.\end{cases}$
+Proof 2.3.
 We only need to calculate the subgradients of $s$ at 00, i.e., $i\in\mathcal{J}(x)$. If $r^{\prime}(|0^{+}|)<\infty$, define for any $x\in\mathbb{R}$,
@@
 Obviously, $\hat{s}(\cdot)$ is smooth on $\mathbb{R}$ with $\hat{s}^{\prime}(0)=r^{\prime}(|0^{+}|)>0$ and $s(x)=\max\{\hat{s}(x),-\hat{s}(x)\}$.
-By [42, Exercise 8.31], we have $\partial s(x_{i})=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
-On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ hold true for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, which means $r^{\prime}(|0^{+}|)\in\hat{\partial}s(0),-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$.
-It follows from [42, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$.
+By [34, Exercise 8.31], we have $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
+On the other hand, $\liminf\limits_{x\rightarrow 0,x\neq 0}\frac{s(x)-s(0)-v(x-0)}{|x-0|}\geq 0$ holds for $v=s^{\prime}(0^{+})=r^{\prime}(|0^{+}|)$ and $v=s^{\prime}(0^{-})=-r^{\prime}(|0^{+}|)$, which means $r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$ and $-r^{\prime}(|0^{+}|)\in\hat{\partial}s(0)$.
+It follows from [34, Theorem 8.6] that $\partial s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]\subset\hat{\partial}s(0)\subset\partial s(0)$.
 Hence, $\partial s(0)=\hat{\partial}s(0)=[-r^{\prime}(|0^{+}|),r^{\prime}(|0^{+}|)]$.
@@
 Hence, ^$\hat{\partial}s(0)=\partial s(0)=\mathbb{R}$.
-In general, we have shown that (4) is true. It then follows from [42, Proposition 10.5] that (3) is true. ∎
-We can appeal to the classical rule of Fermat [42, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a stationary point of (P), then
-$0\in\partial F(x^{*})=\nabla f(x^{*})+\lambda\partial\psi(x^{*})$ (5)
+In general, we have shown that (3) is true. It then follows from [34, Proposition 10.5] that (2) is true.
+We can appeal to the classical rule of Fermat [34, Theorem 10.1] to characterize a local minimum. If $x^{*}$ is a local minimum of (P), then
+(4) $0\in\partial F(x^{*})=\nabla f(x^{*})+\lambda\partial\psi(x^{*})$
 This condition, combined with Assumption 2 and the explicit form of the subgradients, yields the following results.
-Proposition 3.
-If $x^{*}$ is a local minimum for (P), then the following holds
-$\nabla_{i}f(x^{*})=-\text{sign}(x_{i}^{*})\lambda r^{\prime}(|x^{*}_{i}|),$ $i\in\mathcal{I}^{*},$ (6)
-$\nabla_{i}f(x^{*})\in(-\lambda r^{\prime}(|0^{+}|),+\lambda r^{\prime}(|0^{+}|)),$ $i\in\mathcal{J}^{*}.$ (7)
-Notice that if $r^{\prime}(|0^{+}|)=\infty$, (7) holds naturally true. Any point that satisfies the necessary optimality conditions (6) and (7) is referred to as the stationary point.
-In the following, we demonstrate that the second-order optimality condition can be characterized by the nonzero components.
-Theorem 4.
+Proposition 2.4.
+If $x^{*}$ is a local minimum for (P), then the following hold
+(5) $\nabla_{i}f(x^{*})=-\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x^{*}_{i}|),$ $i\in\mathcal{I}^{*},$
+(6) $\nabla_{i}f(x^{*})\in(-\lambda r^{\prime}(|0^{+}|),+\lambda r^{\prime}(|0^{+}|)),$ $i\in\mathcal{J}^{*}.$
+Notice that if $r^{\prime}(|0^{+}|)=\infty$, (6) holds naturally true. Any point that satisfies the necessary optimality conditions (5) and (6) is referred to as a stationary point.
+In the following, we demonstrate that the second-order optimality condition can be characterized by nonzero components.
+Theorem 2.5.
   1. (i)
-    (necessary) If $x^{*}$ is a local minimum of $F$, then
-    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (8)
+    (necessity) If $x^{*}$ is a local minimum of $F$, then
+    (7) $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}\geq$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$
     or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))\geq 0$.
   2. (ii)
-    (sufficient) A stationary point $x^{*}$ of $F$ is a strict local minimum if
-    $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}},$ (9)
+    (sufficiency) A stationary point $x^{*}$ of $F$ is a strict local minimum if
+    (8) $d^{T}\nabla^{2}f(x^{*})d+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r^{\prime\prime}(|x^{*}_{i}|)d_{i}^{2}>$ $\ 0,\quad\forall d\in\mathcal{M}_{\mathcal{I}^{*}}\setminus\{0\},$
     or, equivalently, $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>0$.
-Proof.
-(i) is proved in [43, Theorem 3.3]. We only prove (ii).
-Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of nonzero $\{u^{k}\}$ satisfying
-$F(x^{*}+u^{k})\leq F(x^{*})\quad{\rm and}\quad\lim\limits_{k\rightarrow\infty}u^{k}=0.$ (10)
+Proof 2.6.
+(i) is proved in [8, Theorem 3.3]. We only prove (ii).
+Assume by contradiction that $x^{*}$ is not a strict local minimum. Then there exists a sequence of nonzero vectors $\{u^{k}\}$ satisfying
+(9) $F(x^{*}+u^{k})\leq F(x^{*})\quad{\rm and}\quad\lim\limits_{k\rightarrow\infty}u^{k}=0.$
 Consider $F_{|\mathcal{I}^{*}}$, which is locally smooth around $x^{*}_{\mathcal{I}^{*}}$ and $x^{*}_{\mathcal{I}^{*}}$ is stationary for $F_{|\mathcal{I}^{*}}$.
-It follows from $u^{k}=u_{\mathcal{I}^{*}}^{k}+u_{\mathcal{J}^{*}}^{k}$ that
-$F(x^{*}+u^{k})-F(x^{*})$ (11)
+It follows from $u^{k}=\begin{bmatrix}u_{\mathcal{I}^{*}}^{k};u_{\mathcal{J}^{*}}^{k}\end{bmatrix}$ that
+(10) $F(x^{*}+u^{k})-F(x^{*})$
 $=$ $\ f(x^{*}+u^{k})+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
 $=$ $\ f(x^{*}+u^{k})+\lambda\sum\limits_{i\in\mathcal{I}^{*}}r(|x^{*}_{i}+u_{i}^{k}|)+f(x^{*}+u_{\mathcal{I}^{*}}^{k})$
-$-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
-$=$ $\ f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})+F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}^{k}r(|u_{i}^{k}|).$
-From the classic optimality condition for smooth function and (9), it is easy to see that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$. Therefore for sufficiently large $k$,
-$F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})>0.$ (12)
-Moreover, by the Lagrangian mean value theorem, there exists sufficiently small $\epsilon^{k}\in(0,\min_{i}\{|u_{i}^{k}|\})$ such that
-$r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$ (13)
-The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}\|_{\infty})$, which, together with (13), yields that
+$-f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}])+\lambda\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})$
+$=$ $\ f(x^{*}+u^{k})-f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}])+F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})+\lambda\sum\limits_{i\in\mathcal{J}^{*}}^{k}r(|u_{i}^{k}|).$
+From the classic optimality condition for a smooth function and (8), it is easy to see that $x^{*}$ is a strict local minimum of $F_{|\mathcal{I}^{*}}(x)$. Therefore for sufficiently large $k$,
+(11) $F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}}+u_{\mathcal{I}^{*}}^{k})-F_{|\mathcal{I}^{*}}(x^{*}_{\mathcal{I}^{*}})>0.$
+Moreover, by the Lagrangian mean value theorem,
+(12) $r(|u_{i}^{k}|)=r(|u_{i}^{k}|)-r(0)=r^{\prime}(\hat{u}_{i}^{k})|u^{k}_{i}|,\text{ \ with }\hat{u}_{i}^{k}\in(0,|u_{i}^{k}|).$
+The concavity (Assumption 1) of $r$ implies that $r^{\prime}(\hat{u}_{i}^{k})\geq r^{\prime}(\|u^{k}_{\mathcal{J}^{*}}\|_{\infty})$, which, together with (12), yields that
 $\sum\limits_{i\in\mathcal{J}^{*}}r(|u_{i}^{k}|)\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{1}\geq\sum\limits_{i\in\mathcal{J}^{*}}r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})\|u^{k}_{\mathcal{J}^{*}}\|_{2}.$
 This, combined with (11) and (10), yields
-$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|,$ (14)
-where $\eta^{k}:=\frac{f(x^{*}+u^{k})-f(x^{*}+u_{\mathcal{I}^{*}}^{k})}{\|u_{\mathcal{J}^{*}}^{k}\|}$. By the definition of gradient,
-$\lim_{u^{k}\to 0}\frac{f(x^{*}+u_{\mathcal{J}^{*}}^{k})-f(x^{*})-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=\lim_{u^{k}\to 0}\eta^{k}-\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}=0,$
-implying $\limsup\limits_{k\to\infty}\eta^{k}=\limsup\limits_{k\to\infty}\nabla_{\mathcal{J}^{*}}f(x^{*})^{T}\frac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}\leq\|\nabla f(x^{*})\|_{\infty}$.
-By (7), $\limsup\limits_{k\to\infty}\eta^{k}\leq\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$.
-Therefore, $\eta^{k}<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$.
+(13) $F(x^{*}+u^{k})-F(x^{*})>(\lambda r^{\prime}(\|\hat{u}^{k}_{\mathcal{J}^{*}}\|_{\infty})+\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|,$
+where
+$\eta^{k}:=\frac{f(x^{*}+u^{k})-f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}])}{\|u_{\mathcal{J}^{*}}^{k}\|}=\nabla f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}]+t[\boldsymbol{0};u_{\mathcal{J}^{*}}^{k}])^{T}[\boldsymbol{0};u_{\mathcal{J}^{*}}^{k})]/\|u_{\mathcal{J}^{*}}^{k}\|,$
+and $t\in[0,1]$, implying
+$\limsup\limits_{k\to\infty}|\eta^{k}|\leq\limsup\limits_{k\to\infty}\|\nabla f(x^{*}+[u_{\mathcal{I}^{*}}^{k};\boldsymbol{0}]+t[\boldsymbol{0};u_{\mathcal{J}^{*}}^{k}])\|_{\infty}\|\tfrac{u_{\mathcal{J}^{*}}^{k}}{\|u_{\mathcal{J}^{*}}^{k}\|}\|_{1}=\|\nabla f(x^{*})\|_{\infty}.$
+It follows from $\|\nabla f(x^{*})\|_{\infty}<\lambda r^{\prime}(|0^{+}|)$ by (6) that $|\eta^{k}|<\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})$ for sufficiently large $k$, as $\|\hat{u}^{k}\|_{\infty}\to 0$.
 It follows from (13) that for sufficiently large $k$,
-$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})-\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|>0,$
-contradicting (10). ∎
+$F(x^{*}+u^{k})-F(x^{*})\geq(\lambda r^{\prime}(\|\hat{u}^{k}\|_{\infty})+\eta^{k})\|u^{k}_{\mathcal{J}^{*}}\|>0,$
+contradicting (9).
 2.2 Strict Saddle Points
-In this section, we define the strict saddle point of the problem (P).
-In smooth settings, the strict saddle can be defined as the points with negative minimum eigenvalue of the Hessian matrix, which indicates a descent direction.
-For our case, Theorem 4 implies that a saddle point can be defined using nonzeros.
-Definition 3.
-A stationary point $x^{*}$ is a strict saddle point of $F(x)$ if $d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$ for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$.
-If all stationary points of $F(x)$ are strict saddle points or local minimums, $F$ satisfies the strict saddle property.
+In this subsection, we define the strict saddle point of the problem (P).
+In smooth settings, the strict saddle point can be defined by virtue of a negative minimum eigenvalue of the corresponding Hessian, indicating a descent direction.
+For our case, Theorem 2.5 implies that a saddle point can be defined using nonzeros.
+Definition 2.7.
+A stationary point $x^{*}$ of $F(x)$ is a strict saddle point if
+$d^{T}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})d<0$
+for $\forall d\in\mathbb{R}^{|\mathcal{I}^{*}|}\setminus\{0\}$. If all stationary points of $F(x)$ are strict saddle points or local minima, then we say that $F$ satisfies the strict saddle property.
 This definition is related to the “active strict saddle point” proposed in [15], in which it is shown that the presence of negative curvature alone in nonsmooth settings does not guarantee escape from saddle points.
@@
 Our definition also belongs to this kind, since $\mathcal{M}_{\mathcal{I}^{*}}$ is an active manifold that contains $x^{*}$.
-To show that the strict saddle property is mathematically generic, many works show that this property holds for a full measure of linear perturbation [23, Theorem 2.9], [44, Theorem 2].
-Namely, the perturbed function has the strict saddle property for almost all $v\in\mathbb{R}^{n}$.
-In our case, we also have the same results by considering the following perturbed nonsmooth regularization problem for $v\in\mathbb{R}^{n}$,
-$\min\limits_{x\in\mathbb{R}^{n}}F_{v}(x):=F(x)-\langle v,x\rangle=f(x)+\lambda\psi(x)-\langle v,x\rangle.$ (15)
-To show the desired result, we need the assistance of the famous Sard’s theorem. Our proof of the strict saddle property mainly follows the proof of [45, Lemma 2.7], which is first proved in [46].
-Definition 4.
-Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, if the derivative at $x$ is singular (Jacobian is not full rank), then $x$ is called a critical point of $f$.
-Theorem 5.
-(Sard’s Theorem) Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, and let $C$ be the set of stationary points of $f$ in $\mathcal{M}$.
-Then $f(C)$ has measure zero in $\mathcal{N}$.
-The genericity of the strict saddle property is stated below.
-Theorem 6.
-For a full Lebesgue measure set of perturbations $v\in\mathbb{R}^{n}$, the perturbation function $F_{v}(x):=F(x)-\langle v,x\rangle$ satisfies
-$\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d\neq 0,$ (16)
-for all stationary points $x$ of $F_{v}$ with $\mathcal{I}=\mathcal{I}(x)$.
-Proof.
+To show that the strict saddle property is mathematically generic, many works show that this property holds for a full measure of linear perturbation [15, Theorem 2.9], [4, Theorem 2].
+Namely, the perturbed function has the strict saddle property for almost all linear perturbations.
+In our case, we also have the same results by considering the following linearly perturbed nonsmooth regularization problem for $v\in\mathbb{R}^{n}$,
+(14) $\min\limits_{x\in\mathbb{R}^{n}}F_{v}(x):=F(x)-\langle v,x\rangle=f(x)+\lambda\psi(x)-\langle v,x\rangle.$
+To show the desired result, we need the assistance of the famous Sard’s theorem [matsumoto2002introduction].
+Our proof of the strict saddle property mainly follows the proof of [40, Lemma 2.7], which is first proved in [milnor2025lectures].
+Definition 2.8.
+Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds. If the derivative at $x$ is singular (Jacobian is not of full rank), then $x$ is called a critical point of $f$.
+Theorem 2.9.
+(Sard’s Theorem) Let $f:\mathcal{M}\rightarrow\mathcal{N}$ be a smooth map of manifolds, and $C$ be the set of critical points of $f$ in $\mathcal{M}$. Then $f(C)$ has measure zero in $\mathcal{N}$.
+The genericity of the strict saddle property is stated in the following.
+Theorem 2.10.
+Let $\mathcal{I}$ stand for $\mathcal{I}(x)$. For a full Lebesgue measure set of perturbations $v\in\mathbb{R}^{n}$, the perturbation function $F_{v}(x):=F(x)-\langle v,x\rangle$ satisfies
+(15) $\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d\neq 0,$
+for all stationary points $x$ of $F_{v}$.
+Proof 2.11.
 Let $s\in\{+1,-1,0\}^{n}$ be a vector of sign pattern and
@@
 be a manifold.
-Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$ with $\mathcal{I}=\mathcal{I}(s)$.
+Consider the behavior of $F_{v}$ in $\mathcal{M}(s)$, where $\mathcal{I}=\mathcal{I}(x)=\mathcal{I}(s)$ for all $x\in\mathcal{M}(s)$.
+Let $v:\mathcal{M}(s)\rightarrow\mathbb{R}^{n}$ be a smooth map: $[v(x)]_{\mathcal{I}}=\nabla_{\mathcal{I}}F(x)$ and $[v(x)]_{\mathcal{I}^{c}}=x$.
 The Jacobian of $v$ at $x$ is then given by $Dv(x)=\begin{bmatrix}\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)&0\\ 0&I\end{bmatrix}$, and $Dv(x)$ has full rank if and only if $\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)$ has full rank.
-By Sard’s Theorem,
-$\{v(x)\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n},$
-which is equivalent to
-$\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}(s)|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}.$
+By Sard’s Theorem, $\{v(x)\mid\nabla_{\mathcal{I}\mathcal{I}}^{2}F(x)\text{ is singular},x\in\mathcal{M}(s)\}\text{ has measure 0 in }\mathbb{R}^{n}$, which implies its subset
+$\mathcal{V}(s):=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}(s)|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathcal{M}(s)\}$
+has measure 0 in $\mathbb{R}^{n}$.
 Now we consider all the possible sign pattern vectors $s\in\{+1,-1,0\}^{n}$. There are at most $3^{n}$ different patterns. Therefore,
 $\bigcup_{s\in\{+1,-1,0\}^{n}}\mathcal{V}(s)=\{v\mid 0\in\partial F_{v}(x),\inf\limits_{d\in\mathbb{R}^{|\mathcal{I}|}\setminus\{0\}}d^{T}\nabla^{2}_{\mathcal{I}\mathcal{I}}F(x)d=0,x\in\mathbb{R}^{n}\}$
-has measure 0 in $\mathbb{R}^{n}$. ∎
-Center Stable Manifold Theorem
-As many other works in smooth setting like [47], we interpret our algorithm as the fixed point iteration of a nonlinear map $T$.
-Likewise, we define the fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as the unstable fixed points, as described below.
-Definition 5.
+has measure 0 in $\mathbb{R}^{n}$.
+2.3 Center Stable Manifold Theorem
+As with many other works in smooth settings like [24], we will interpret our algorithm as the fixed point iteration of a nonlinear map $T$.
+Likewise, we define fixed points with at least one eigenvalue of $DT$ larger than $1$ in magnitude as unstable fixed points, as described below.
+Definition 2.12.
 (Unstable fixed point) Let $x^{*}$ be a fixed point on a $C^{1}$ map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
 $x^{*}$ is called an unstable fixed point of $T$ if $DT(x^{*})$ has at least one eigenvalue of magnitude strictly larger than 1.
-The Center Stable Manifold (CSM) Theorem plays a crucial role in deriving the avoidance of strict saddle points.
-The CSM theorem establishes a favorable property for the fixed-point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, which converge to the fixed point with at least one eigenvalue of $DT$ larger than $1$ in magnitude, has Lebesgue measure zero.
-Theorem 7.
+The Center Stable Manifold Theorem (CSM) plays a crucial role in the avoidance of strict saddle points.
+The CSM theorem establishes a favorable property for the fixed-point iteration of a $C^{1}$ local diffeomorphism map: the set of initial points around a fixed point, from which an algorithm converges to the fixed point with at least one eigenvalue of $DT$ greater than $1$ in magnitude, has Lebesgue measure zero.
+Theorem 2.13.
 (The Center Stable Manifold Theorem) Let 00 be a fixed point for the $C^{1}$ local diffeomorphism $T:U\rightarrow\mathbb{R}^{n}$ where $U$ is a neighborhood of zero in $\mathbb{R}^{n}$.
-Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant splitting of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one, and greater than one.
+Let $E^{s}\oplus E^{c}\oplus E^{u}$ be the invariant split of $\mathbb{R}^{n}$ into the generalized eigenspaces of $DT(0)$ corresponding to eigenvalues of absolute value less than one, equal to one and greater than one.
 There exists a $C^{1}$ embedded disk $W_{loc}^{cs}$ that is tangent to $E_{s}\oplus E_{c}$ at 0 and a neighborhood $B$ around 0 such that $T(W_{loc}^{cs})\cap B\subset W_{loc}^{cs}$.
 In addition, if $T^{k}(x)\in B$ for all $k\geq 0$, then $x\in W_{loc}^{cs}$.
-If $\nabla T(0)$ is invertible, then $T$ is locally a diffeomorphism around 0.
-The CSM theorem only guarantees the avoidance of strict saddle point in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points that converge to any unstable fixed point has measure zero.
+If $DT(0)$ is invertible, then $T$ is locally a diffeomorphism around 0.
+The CSM theorem only guarantees the avoidance of strict saddle points in the neighborhood for local diffeomorphism, and this local result needs to be extended to the global setting to guarantee that the set of almost all initial points from which an algorithm converges to any unstable fixed point has a measure zero.
 This requires that the map has the so-called 0-property, meaning the preimage of a measure zero set under this map still has measure zero.
-This result was established in [47, Theorem 2] for a global diffeomorphism map and then for a global lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [23].
-Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown using the fact that $\mathbb{R}^{n}$ is second countable no matter if the number of stationary points is countable (see [23, Corollary 2.12] for a standard proof).
+This result was established in [24, Theorem 2] for a global diffeomorphism map and then for a global lipeomorphism (A map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is called a lipeomorphism if it is a bijection and both $T$ and $T^{-1}$ are Lipschitz continuous) in [15].
+Moreover, to extend this property from a particular stationary point $x^{*}$ to the set of all stationary points, this can be shown using the fact that $\mathbb{R}^{n}$ is second countable no matter if the number of stationary points is countable (see [15, Corollary 2.12] for a standard proof).
 We list their results below.
-Theorem 8.
-[47, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$.
-Let $\mathcal{U}_{T}$ consist of all unstable fixed points $x$ of $T$ Then the set of initial points that converge to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
-Corollary 1.
-[23, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $\nabla T(x)$ is invertible.
+Theorem 2.14.
+[24, Theorem 2] Let $g$ be a $C^{1}$ mapping from $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $det(DT(x))\neq 0$ for all $x\in\mathbb{R}^{n}$.
+Let $\mathcal{U}_{T}$ consist of all unstable fixed points $x$ of $T$ Then the set of initial points from which an algorithm converges to an unstable fixed point has measure zero, $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
+Corollary 2.15.
+[15, Corollary 12] Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a lipeomorphism and let $\mathcal{U}_{T}$ consists of all unstable fixed points $x$ of $T$ at which the Jacobian $DT(x)$ is invertible.
 Then, the set of initial points attracted by such fixed points $\mathscr{U}=\{x\in\mathbb{R}^{n}:\lim\limits_{k\to\infty}T^{k}(x)\in\mathcal{U}_{T}\}$ has zero Lebesgue measure.
@@
   1. (1)
-    Interpret the iterative algorithm as a fixed point iteration, and build the equivalence of the unstable fixed points and the strict saddle points.
+    Interpret the iterative algorithm as a fixed point iteration and build the equivalence of the unstable fixed points and the strict saddle points.
   2. (2)
-    Show that the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at the fixed point), so that the CSM theorem guarantees the avoidance of converging to the strict saddle point.
+    Show that the fixed point iteration is locally a $C^{1}$ local diffeomorphism around a strict saddle point (e.g., a $C^{1}$ smooth map with invertible Jacobian at the fixed point) so that the CSM theorem guarantees the avoidance of converging to a strict saddle point.
   3. (3a)
@@
 The ${\rm IRL}_{1}$ algorithm is an effective method to solve problem (P). At each iteration, the ${\rm IRL}_{1}$ algorithm solves the following weighted $\ell_{1}$ regularized subproblem,
-$x^{k+1}=S(x^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i})|y_{i}|,$
-where $w(x_{i})=r^{\prime}(|x_{i}|)$.
-This algorithm may quickly converge to local minimizers due to the rapid growth of the weight values, especially for the case where $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$ can be quickly encountered.
+$x^{k+1}=S^{x}(x^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}(y-x^{k})+\tfrac{\beta}{2}\|y-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}\tilde{w}^{k}_{i}|y_{i}|,$
+where weight $\tilde{w}_{i}^{k}=r^{\prime}(|x_{i}^{k}|)$.
+This algorithm may quickly converge to local minima due to the rapid growth of the weight values, especially for the case where $r^{\prime}(|0^{+}|)=\infty$ and $w(0)=\infty$.
 To address this issue, a common approach is to add a perturbation $\epsilon_{i}>0$ to each $r$ and drive $\epsilon_{i}$ to 0 during the iteration, resulting in the following approximation $F(x,\epsilon)$ of $F(x)$
-$\min\limits_{x,\epsilon\in\mathbb{R}^{n}}F(x,\epsilon):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}).$ (17)
-Here we also treat $\epsilon$ as variables, so that the algorithm updates $x$ and $\epsilon$ by $(x^{k+1},\epsilon^{k+1})=S(x^{k},\epsilon^{k})$
-$\begin{cases}x^{k+1}=S^{x}(x^{k},\epsilon^{k})=\arg\min\limits_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum\limits_{i=1}^{n}w_{i}^{k}|y_{i}|,\\ \epsilon^{k+1}=S^{\epsilon}(x^{k},\epsilon^{k})=\mu\epsilon^{k}\end{cases}$
-with $w_{i}^{k}=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and $\mu\in(0,1)$.
-The subproblem solution can be given explicitly using the solft-thresholding operator,
-$S^{x}(x^{k},\epsilon)=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$ (18)
+(16) $\min\limits_{x\in\mathbb{R}^{n},\epsilon\in\mathbb{R}^{n}_{+}}F(x,\epsilon):=f(x)+\lambda\sum\limits_{i=1}^{n}r(|x_{i}|+\epsilon_{i}).$
+Here we treat $\epsilon$ also as variables, so that the algorithm updates $x$ and $\epsilon$ by
+$(x^{k+1};\epsilon^{k+1})=S(x^{k},\epsilon^{k}),$
+where $S=[S^{x};S^{\epsilon}]$ and
+$x^{k+1}$ $=S^{x}(x^{k},\epsilon^{k})=\arg\min\limits_{y}G(y;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum\limits_{i=1}^{n}w_{i}^{k}|y_{i}|,$
+$\epsilon^{k+1}$ $=S^{\epsilon}(x^{k},\epsilon^{k})=\mu\epsilon^{k}$
+with $w_{i}^{k}:=w(x_{i}^{k},\epsilon_{i}^{k})=r^{\prime}(|x_{i}^{k}|+\epsilon_{i}^{k})$ and $\mu\in(0,1)$.
+The subproblem solution can be given using the soft-thresholding operator,
+(17) $S^{x}(x^{k},\epsilon)=\mathbb{S}_{\lambda w^{k}/\beta}(x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})),$
 or explicitly,
-$S^{x}_{i}(x^{k},\epsilon^{k})=\begin{cases}x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ 0,&\text{otherwise }.\end{cases}$ (19)
-To leverage the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a Lipeomorphism globally.
-However, this is not true in our case since $S$ is generally noninvertible.
+(18) $S^{x}_{i}(x^{k},\epsilon^{k})=\begin{cases}x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})+\lambda w^{k}_{i})>0,\\ x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i}),&\text{if }x_{i}^{k}-\tfrac{1}{\beta}(\nabla_{i}f(x^{k})-\lambda w^{k}_{i})<0,\\ 0,&\text{otherwise }.\end{cases}$
+To take advantage of the CSM theorem, the map $S$ needs to be a local diffeomorphism at a stationary point and a lipeomorphism globally.
+However, this is not true in our case, since $S$ is generally noninvertible.
 In particular, $S$ is not a bijection, as it maps multiple values to 00.
 Motivated by the technique used in [15], we propose to use the damped operator: $T=(1-\alpha)I+\alpha S$, which can be written explicitly as
-$(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S^{x}(x^{k},\epsilon^{k})\\ (1-\alpha(1-\mu))\epsilon^{k}\end{bmatrix}$ (20)
+(19) $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})=\begin{bmatrix}(1-\alpha)x^{k}+\alpha S^{x}(x^{k},\epsilon^{k})\\ (1-\alpha(1-\mu))\epsilon^{k}\end{bmatrix}$
 with $\epsilon_{0}\in\mathbb{R}^{n}_{++}$ and $\alpha\in(0,1)$.
@@
 We have the following result to connect the optimality of (P) and the fixed point of our iteration.
-Proposition 9.
+Proposition 3.1.
 $x^{*}$ is a stationary point of (P) if and only if $x^{*}=S^{x}(x^{*},0)$. In addition, $x^{*}$ is a stationary point of (P) if and only if $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$.
-Proof.
+Proof 3.2.
 Suppose $x^{*}$ is a stationary point.
 It follows from (5) that if $x_{i}^{*}>0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)]=x_{i}^{*}>0$, implying $S_{i}^{x}(x^{*},0)=x_{i}^{*}$.
-This result also holds true by the same argument if $x_{i}^{*}<0$.
+This result also holds by the same argument if $x_{i}^{*}<0$.
 As for $x_{i}^{*}=0$, $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|0^{+}|)]<0$ by (6).
-Same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$.
+The same argument also yields $x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda w(x_{i}^{*},0)]=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|0^{+}|)]>0$.
 Therefore, $S_{i}^{x}(x^{*},0)=0=x_{i}^{*}$.
@@
 If $x_{i}^{*}>0$, it must be true that $x_{i}^{*}=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)]=x_{i}^{*}-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]>0$, implying $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)]$.
-Same argument yields $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$.
+The same argument yields $0=x_{i}^{*}-S_{i}^{x}(x^{*},0)=-\frac{1}{\beta}[\nabla_{i}f(x^{*})-\lambda r^{\prime}(|x_{i}^{*}|)]$ if $x_{i}^{*}<0$.
 Therefore, $\nabla_{i}f(x^{*})+\lambda\text{sign}(x_{i}^{*})r^{\prime}(|x_{i}^{*}|)=0$ for $i\in\mathcal{I}^{*}$.
@@
 By Assumption 2, it satisfies (5) and (6).
-The rest part of the statement of this proposition is obviously true by noticing that $(x^{*},\epsilon^{*})=T(x^{*},\epsilon^{*})$ implies that $\epsilon^{*}=0$. ∎
-1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
+The rest part of the statement of this proposition is obviously true by noticing that $(x^{*};\epsilon^{*})=T(x^{*},\epsilon^{*})$ implies that $\epsilon^{*}=0$.
+1:  Input $\beta>\alpha L_{\nabla f}/2$, $\alpha\in(0,1)$, $x^{0}$, $\epsilon^{0}>0$ and $\mu\in(0,1)$.
 2:  Initialize: set $k=0$.
@@
 Assumption 3.
-Given the initial point $(x^{0}$, $\epsilon^{0})$ and $F^{0}=F(x^{0},\epsilon^{0})$, the level set $\mathcal{L}(F^{0}):=\{(x,\epsilon)\mid F(x,\epsilon)\leq F^{0}\}$ is bounded.
-This assumption means $f$ is Lipschitz continuous and Lipschitz differentiable on $\mathcal{L}(F^{0})$ with constant $L_{f}$, i.e., $\forall x,y\in\mathbb{R}^{n}$ (we use the same constant for $f$ and $\nabla f$ for simplicity).
-$|f(x)-f(y)|\leq L_{f}\|x-y\|\text{ \ and \ }\|\nabla f(x)-\nabla f(y)\|\leq L_{f}\|x-y\|.$ (21)
-This assumption also implies that if $r^{\prime}(|0^{+}|)=\infty$, then for any stationary point $x^{*}$, $|x_{i}^{*}|\geq(r^{\prime})^{-1}(L_{f}/\lambda),\forall i\in\mathcal{I}^{*}$ by (6).
-Therefore, we know $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above and this is also the case for $r^{\prime}(|0^{+}|)<\infty$.
-Overall, we can conclude that for any stationary point $x^{*}\in\mathcal{L}(F^{0})$, there exists $\rho$ such that
-$\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho.$ (22)
-We first show the monotonicity of $F(x,\epsilon)$. Define the reductions of $G$ and $F$ caused by $x^{k+1}$ as
-$\Delta F(x^{k+1},\epsilon^{k+1})=$ $\ F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k+1}),$
+Given the initial point $(x^{0}$, $\epsilon^{0})$ and $F^{0}=F(x^{0},\epsilon^{0})$, the level set $\mathcal{L}(F^{0}):=\{x\mid F(x)\leq F^{0}\}$ is bounded.
+This assumption means that there exists $C>0$ such that
+(20) $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C,\quad\forall x\in\mathcal{L}(F^{0}).$
+It also means $f$ is Lipschitz continuous and differentiable with a Lipschitz gradient on $\mathcal{L}(F^{0})$.
+Hence, there exist constants $L_{f}$ and $L_{\nabla f}$, such that $\forall x,y\in\mathcal{L}(F^{0})$,
+(21) $|f(x)-f(y)|\leq L_{f}\|x-y\|\text{ and }\|\nabla f(x)-\nabla f(y)\|\leq L_{\nabla f}\|x-y\|.$
+Moreover, (21) implies that if $r^{\prime}(|0^{+}|)=\infty$, then for any stationary point $x^{*}\in\mathcal{L}(F^{0})$, $|x_{i}^{*}|\geq(r^{\prime})^{-1}(L_{f}/\lambda),\forall i\in\mathcal{I}^{*}$ by (5) and $r^{\prime\prime}(|x_{i}^{*}|)>0$ is bounded above.
+As for $r^{\prime}(|0^{+}|)<\infty$, it holds naturally that $r^{\prime\prime}(|x_{i}^{*}|)$ is bounded above for $i\in\mathcal{I}^{*}$.
+Overall, we can conclude that for any stationary point $x^{*}\in\mathcal{L}(F^{0})$, there exists $\rho>0$ such that
+(22) $\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho.$
+We first show the monotonicity of $F(x,\epsilon)$. Considering the model $G$ at $(x^{k},\epsilon^{k})$
+$G(x;x^{k},\epsilon^{k}):=\nabla f(x^{k})^{T}x+\frac{\beta}{2}\|x^{k}-x\|^{2}+\lambda\sum_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})|x_{i}|$
+with minimum $y^{k}$, its reduction caused by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ is defined as
 $\Delta G(x^{k+1};x^{k},\epsilon^{k})=\ G(x^{k};x^{k},\epsilon^{k})-G(x^{k+1};x^{k},\epsilon^{k}).$
-Proposition 10.
+The reduction of $F$ caused by $(x^{k+1},\epsilon^{k+1})$ with $\epsilon^{k+1}=(1-\alpha(1-\mu))\epsilon^{k}$ is defined as
+$\Delta F(x^{k+1},\epsilon^{k+1})=$ $\ F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k+1}).$
+Proposition 3.3.
 Suppose Assumptions 1-3 hold.
-Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$.
+Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 1 with $\alpha\in(0,1)$ and $\beta\geq\frac{\alpha L_{\nabla f}}{2}$.
 It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and that the reduction satisfies
-$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
-Moreover, $\lim\limits_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$ and there exists $C>0$ such that
-$\|x^{k}-\tfrac{1}{\beta}\nabla f(x^{k})\|_{\infty}\leq C,$ (23)
-for any $k\in\mathbb{N}$.
-Proof.
+$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2},$
+which means $\lim\limits_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$. Moreover, $\{x^{k}\}\subset\mathcal{L}(F^{0})$ and (20) holds for any $x=x^{k},k\in\mathbb{N}$.
+Proof 3.4.
 Since the concavity of $r(\cdot)$ on $\mathbb{R}_{++}$ gives $r(x_{1})\leq r(x_{2})+r^{\prime}(x_{2})(x_{1}-x_{2})$ for any $x_{1},x_{2}\in\mathbb{R}_{++}$, we have that for $i=1,...,n$,
-$r(|x_{i}^{k+1}|+(\epsilon_{i}^{k})^{2})\leq r(|x_{i}^{k}|+\epsilon_{i}^{k})+r^{\prime}(|x_{i}^{k}|+\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
+$r(|x_{i}^{k+1}|+\epsilon_{i}^{k})\leq r(|x_{i}^{k}|+\epsilon_{i}^{k})+r^{\prime}(|x_{i}^{k}|+\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
 Summing up the above inequality over $i$ yields
-$\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+\epsilon_{i}^{k})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+\epsilon_{i}^{k})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$ (24)
+(23) $\sum\limits_{i=1}^{n}r(|x_{i}^{k+1}|+\epsilon_{i}^{k})-\sum\limits_{i=1}^{n}r(|x_{i}^{k}|+\epsilon_{i}^{k})\leq\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon^{k}_{i})(|x_{i}^{k+1}|-|x_{i}^{k}|).$
 On the other hand, it follows that
-$f(x^{k})-f(x^{k+1})\geq\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}.$ (25)
+(24) $f(x^{k})-f(x^{k+1})\geq\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{\nabla f}}{2}\|x^{k}-x^{k+1}\|^{2}.$
 Combining (23) and (24), we have
-$\Delta F(x^{k+1},\epsilon^{k+1})\geq F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k})$ (26)
-$\geq$ $\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)$
-$=$ $\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2},$
+(25) $\Delta F(x^{k+1},\epsilon^{k+1})\geq F(x^{k},\epsilon^{k})-F(x^{k+1},\epsilon^{k})$
+$\geq$ $\nabla f(x^{k})^{T}(x^{k}-x^{k+1})-\frac{L_{\nabla f}}{2}\|x^{k}-x^{k+1}\|^{2}+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)$
+$=$ $\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{\nabla f}}{2}\|x^{k}-x^{k+1}\|^{2},$
 where the first inequality is due to the fact that $\epsilon^{k+1}\leq\epsilon^{k}$.
-For each subproblem, $y^{k}$ satisfies the optimal condition
-$\nabla f(x^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})=0.$ (27)
+For each subproblem, $y^{k}$ satisfies the optimality condition
+(26) $\nabla f(x^{k})+\lambda W(x^{k},\epsilon^{k})\xi^{k}+\beta(y^{k}-x^{k})=\boldsymbol{0}.$
 with $\xi^{k}\in\partial\|y^{k}\|_{1}$ and $W(x^{k},\epsilon^{k}):={\rm diag}(w(x_{1}^{k},\epsilon_{1}^{k}),...,w(x_{n}^{k},\epsilon_{n}^{k}))$.
-Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimal condition (27), the following holds
-$\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$ (28)
+Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimality condition (26), the following holds
+(27) $\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$
 $=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(|x_{i}^{k}|-|x_{i}^{k+1}|)-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
@@
 $=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2},$
-where the first inequality is by $\alpha\in(0,1)$ and the second is by the convexity of ||$\cdot$|| ⋅ |. Combine (26), (37) and get
-$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
+where the first inequality is by $\alpha\in(0,1)$ and the second is by the convexity of ||$\cdot$|| ⋅ |. Combining (25) and (27), we get
+$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
 Replacing $k$ with $t$ and summing up from $t=0$ to $k-1$, we have
-$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
+$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
 It follows that $\{(x^{k},\epsilon^{k})\}\in\mathcal{L}(F^{0})$ since $F(x^{k},\epsilon^{k})\leq F(x^{0},\epsilon^{0})$.
@@
 Taking $k\rightarrow\infty$, we have
-$(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
-which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$ that there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎
+$(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
+which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$.
+Moreover, $F(x^{k})\leq F(x^{k},\epsilon^{k})\leq F(x^{0},\epsilon^{0})$, meaning $\{x^{k}\}\subset\mathcal{L}(F^{0})$. Therefore, (20) holds for all $x=x^{k},k\in\mathbb{N}$.
 With the monotonicity of $F$, we derive the global convergence of the ${\rm DIRL}_{1}$ method.
-Theorem 11.
-(Global convergence) Suppose Assumption 1 and 3 hold.
-Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$.
-Then $\epsilon^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is first-order optimal for (P).
-Proof.
+Theorem 3.5.
+(Global convergence) Suppose Assumptions 1 and 3 hold.
+Let the sequence $\{(x^{k},\epsilon^{k})\}$ be generated by ${\rm DIRL}_{1}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$.
+Then $\epsilon^{*}=0$ and $\Omega\neq\emptyset$ and any $x^{*}\in\Omega$ is a stationary point of (P).
+Proof 3.6.
 Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$.
-Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$.
+Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$.
 This also implies $\text{sign}(y_{i}^{k})=\text{sign}(x_{i}^{*}),i\in\mathcal{I}^{*}$ for sufficiently large $k\in\mathcal{S}$.
-The optimality condition of the $k$th subproblem implies
-$-\beta(y_{i}^{k}-x_{i}^{k})\in\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k})\partial|y_{i}^{k}|.$
-If $w(x_{i}^{*},0)<+\infty$, it follows from [48, Proposition 2.1.5], [42, Theorem 5.19] and the outer semi-continuity of $\partial r$ that there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that
-$0=\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)\xi^{*}_{i}$
-as $y_{i}^{k}-x_{i}^{k}\overset{\mathcal{S}}{\to}0$, $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$.
-Therefore, for $w(x_{i}^{*},0)=r^{\prime}(|x_{i}^{*}|)<\infty,i\in\mathcal{I}^{*}$, (6) is satisfied; for $w(x_{i}^{*},0)=r^{\prime}(|0^{+}|)<\infty,i\in\mathcal{J}^{*}$, (7) is satisfied by Assumption 2.
-Therefore, $x^{*}$ is stationary.
-∎
-3.2 Avoidance of strict saddle point
-The map $S$ has the property of so-called “active manifold identification”, which means the iterates fall onto a smooth manifold after finite iterations.
+The optimality condition of the $k$th subproblem (26) implies $-\beta(y_{i}^{k}-x_{i}^{k})\in\nabla_{i}f(x^{k})+\lambda w(x_{i}^{k},\epsilon_{i}^{k})\partial|y_{i}^{k}|.$ If $w(x_{i}^{*},0)<+\infty$, it follows from the outer semicontinuity of $\partial r$ that there exists $\xi^{*}_{i}\in\partial|x^{*}_{i}|$ such that $0=\nabla_{i}f(x^{*})+\lambda w(x_{i}^{*},0)\xi^{*}_{i}$ as $y_{i}^{k}-x_{i}^{k}\overset{\mathcal{S}}{\to}0$, $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$.
+Therefore, for $w(x_{i}^{*},0)=r^{\prime}(|x_{i}^{*}|)<\infty,i\in\mathcal{I}^{*}$, (5) is satisfied; for $w(x_{i}^{*},0)=r^{\prime}(|0^{+}|)=\infty,i\in\mathcal{J}^{*}$, (6) is satisfied by Assumption 2.
+Therefore, $x^{*}$ is a stationary point.
+3.2 Avoidance of strict saddle points
+The map $S^{x}$ has the property of so-called “active manifold identification”, which means that the iterates fall onto a smooth manifold after finite iterations.
 The iteratively reweighted $\ell_{1}$ algorithms were proved to have this property by [41, 42] and the manifold in this case is simply the subspace consisting of nonzeros at a limit point.
-Theorem 12.
-(Model Identification and Local Smoothness) Under Assumption 1-3, for any stationary point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
+Theorem 3.7.
+(Model Identification and Local Smoothness) Under Assumptions 1-3, for any stationary point $(x^{*},\epsilon^{*})$, there exists a sufficiently small neighborhood $B$ of $(x^{*},\epsilon^{*})$, so that the following hold:
   1. (i)
@@
     $S$ is $C^{1}$ smooth in $B$ and $S^{x}(x,\epsilon)=x-\frac{1}{\beta}g(x,\epsilon)$ with
-    $g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
+    (28) $g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
   3. (iii)
     Map $T$ is $C^{1}$-smooth in $B$.
-Proof.
+Proof 3.8.
 (i) Since $S$ is a continuous map and $x^{*}=S^{x}(x^{*},\epsilon^{*})$, we have
@@
 Therefore, there exists a neighborhood $B$ such that $|x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)|<\tfrac{1}{\beta}\lambda w(x_{i},\epsilon_{i})$, $\forall(x,\epsilon)\in B$, which is equivalent to $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))<0$ and $x_{i}-\tfrac{1}{\beta}(\nabla_{i}f(x)-\lambda w(x_{i},\epsilon_{i}))>0$, meaning $S_{i}(x,\epsilon)=0$.
-(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agrees with following constrained problem
-$S^{x}(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$ (29)
+(ii) It follows from (i) that there exists a neighborhood $B$ around $(x^{*},\epsilon^{*})$, so that the subproblem at $(x^{k},\epsilon^{k})\in B$ agrees with the following constrained problem,
+(29) $S^{x}(x^{k},\epsilon^{k})=$ $\ \arg\min\limits_{x\in x^{*}+\mathcal{M}_{\mathcal{I}^{*}}}\{\nabla f(x^{k})^{T}x+\tfrac{\beta}{2}\|x-x^{k}\|^{2}+\lambda\sum_{i=1}^{n}w^{k}_{i}\text{sign}(x_{i}^{*})x_{i}\}$
 $=$ $\ \arg\min\limits_{x\in\mathbb{R}^{n}}\{g(x^{k},\epsilon^{k})^{T}x+\frac{\beta}{2}\|x-x^{k}\|^{2}\},$
-where each component of $g:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ is defined by
-$g_{i}(x,\epsilon)=\left\{\begin{aligned} &\beta x_{i},&i\in\mathcal{J}^{*},\\ &\nabla_{i}f(x)+\lambda r^{\prime}(x_{i}\text{sign}(x_{i}^{*})+\epsilon_{i}^{2})\text{sign}(x_{i}^{*}),&i\in\mathcal{I}^{*}.\end{aligned}\right.$
-Therefore, $S$ is $C^{1}$ smooth in $B$.
-(iii) It follows from (29) that the iteration $T$ is $C^{1}$ smooth in $B$. ∎
+where each component of $g:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ is defined by (28). Therefore, $S$ is $C^{1}$ smooth in $B$.
+(iii) It follows from (29) that the iteration $T$ is $C^{1}$ smooth in $B$.
 The above theorem shows the local behavior of ${\rm DIRL}_{1}$: it is equivalent to the gradient descent method for the variables in $\mathcal{I}^{*}$ locally around a stationary point in $x^{*}+\mathcal{M}_{\mathcal{I}^{*}}$.
 Specifically, it can be expressed as
-$T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &(1-\alpha(1-\mu))\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$ (30)
-around $(x^{*},\epsilon^{*})$.
-We then proceed to the saddle escape mechanism for DIRL1, which ensures that the algorithm does not converge to the strict saddle points defined in Definition 3 and, instead, converges to local minima.
+(30) $T_{i}(x,\epsilon)=\left\{\begin{aligned} &(1-\alpha)x_{i},&&i\in\mathcal{J}^{*},\\ &x_{i}-\frac{\alpha}{\beta}\nabla_{x_{i}}F(x,\epsilon),&&i\in\mathcal{I}^{*},\\ &(1-\alpha(1-\mu))\epsilon_{i},&&i\in\{n+1,\ldots,2n\}\end{aligned}\right.$
+around $(x^{*};\epsilon^{*})$.
+We then proceed to the saddle escaping mechanism for DIRL1, ensuring that the algorithm does not converge to the strict saddle points defined in Definition 2.7.
 To apply the CSM theorem, we have to show that the unstable fixed point of $T$ corresponds to the strict saddle point of (P).
-Theorem 13.
+Theorem 3.9.
 (Unstable Fixed Point) Suppose $x^{*}$ is a stationary point of (P).
 $x^{*}$ is a strict saddle point of (P) if and only if the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=0$ is an unstable fixed point of $T$.
-Moreover, if $\alpha<\frac{\beta}{\rho}$, $DT(x^{*},\epsilon^{*})$ is invertible.
-Proof.
-First of all, we calculate the Jabobian of $T$ in the neighborhood $B$ described in Theorem 12.
-It holds that $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \bm{0}&\beta I\end{bmatrix}$.
+Moreover, if $\alpha<\frac{\beta}{\rho}$ where $\rho$ is defined by (22), $DT(x^{*},\epsilon^{*})$ is invertible.
+Proof 3.10.
+First of all, we calculate the Jacobian of $T$ in the neighborhood $B$ described in Theorem 3.7.
+It holds that $D_{x}g(x^{*},\epsilon^{*})=\begin{bmatrix}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})\\ \boldsymbol{0}&\beta I\end{bmatrix}$.
 Therefore, the Jacobian of $T$ at $(x^{*},\epsilon^{*})$ is given by
-$DT(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}\ &(1-\alpha(1-\mu))I\end{bmatrix}$
-$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&D_{\epsilon}g(x^{*},\epsilon^{*})\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}$
+$DT(x^{*},\epsilon^{*})=$ $\ \begin{bmatrix}(1-\alpha)I+\alpha(I-\tfrac{1}{\beta}D_{x}g(x^{*},\epsilon^{*}))\ &D_{\epsilon}g(x^{*},\epsilon^{*})\\ \boldsymbol{0}\ &(1-\alpha(1-\mu))I\end{bmatrix}$
+$=$ $\ \begin{bmatrix}I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})&I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{J}^{*}}f(x^{*})&D_{\epsilon}g(x^{*},\epsilon^{*})\\ \boldsymbol{0}&(1-\alpha)I&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&(1-\alpha(1-\mu))I\end{bmatrix}$
 Now we are ready to prove the equivalence between the unstable fixed point of $T$ and the strict saddle point of (P).
-Recall Definition 5 and Definition 3 and notice $\mu\in(0,1)$ and $\alpha\in(0,1)$.
+Recall Definition 2.12 and Definition 2.7 and notice $\mu\in(0,1)$ and $\alpha\in(0,1)$.
 It suffices to show that $I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has an eigenvalue greater than 1 if and only if $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ has a negative eigenvalue.
@@
 $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0\iff\lambda_{\max}(I-\frac{\alpha}{\beta}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>1.$
-If $\alpha$ is selected to satisfy $01-\frac{\alpha}{\beta}\lambda_{\max}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))>$ 0 and by the strict saddle property $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ does not have 0 eigenvalue, we know $DT(x^{*},\epsilon^{*})$ is invertible.
-∎
-To apply Corollary 1, we next show that $T$ is a lipeomorphism, even if it might not be differentiable globally on $\mathbb{R}^{2n}$.
-This is equivalent to show $T$ and its inverse $T^{-1}$ is Lipschitz continuous.
-Proposition 14.
-(Lipeomorphism) Suppose Assumption 1-3 hold.
-Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C))$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ is defined in Proposition 10, and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$.
-Then $T$ is invertible; moreover, $T$ and $T^{-1}$ is Lipschitz continuous (that is, $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$ and $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
-Proof.
-If $r^{\prime}(|0+|)<\infty$, $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|,\epsilon_{i})$ is Lipschitz continuous in $\mathcal{L}(F^{0})$.
-If $r^{\prime}(|0+|)=\infty$, by Proposition 10, $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$.
-When $r^{\prime}(|x_{i}|)\geq w(x_{i},\epsilon_{i})>C$, it must be true that $S_{i}(x,\epsilon)=0$.
-Therefore, for any $S_{i}(x)=0$ with $x\in\mathcal{L}(F^{0})$,
-$r^{\prime}(|x_{i}|)\leq C\text{ \ and \ }|x_{i}|\geq\underline{x}:=(r^{\prime})^{-1}(C)>0$ (31)
-by Assumption 3. This implies that we can equivalently write $S$ as
-$S_{i}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}(x_{i},\epsilon_{i})/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
-with
-$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i})&\text{ if }r^{\prime}(|0+|)<\infty\\ \min(w(\underline{x},0),w(x_{i},\epsilon_{i}))=\min(r^{\prime}(\underline{x}),r^{\prime}(|x_{i}|+\epsilon_{i}))&\text{ if }r^{\prime}(|0+|)=\infty.\end{cases}$
-Therefore, $u(\cdot)=\min(r^{\prime}(\underline{x}),r^{\prime}(\cdot))$ is also Lipschitz continuous in $\mathcal{L}(F^{0})$.
-Let $L_{r}$ be the Lipschitz constant, meaning for any $\hat{x},\tilde{x}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$
-$\|u(\hat{x},\hat{\epsilon})-u(\tilde{x},\tilde{\epsilon})\|\leq L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|.$ (32)
+Moreover, if $\alpha$ is selected to satisfy $01-\frac{\alpha}{\beta}\rho>$ 0 and by the strict saddle property $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ does not have a 0 eigenvalue, we know by (22) that $DT(x^{*},\epsilon^{*})$ is invertible.
+To apply Corollary 2.15, we next show that $T$ is a lipeomorphism, even if it may not be differentiable globally in $\mathbb{R}^{2n}$.
+This is equivalent to showing that $T$ and its inverse $T^{-1}$ are Lipschitz continuous.
+Proposition 3.11.
+(Lipeomorphism) Suppose Assumptions 1-3 hold.
+Let $L_{r}:=r^{\prime\prime}(|(r^{\prime})^{-1}(C)|)$ if $r^{\prime}(|0^{+}|)=\infty$, where $C$ is defined in (20), and $L_{r}:=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$.
+Then $T$ is invertible; moreover, $T$ and $T^{-1}$ are Lipschitz continuous (that is, $T$ is a lipeomorphism) when $10<\alpha<$ 1, $\beta>\alpha L_{\nabla f}/2$ and $\alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r}+\mu)<1$.
+Proof 3.12.
+If $r^{\prime}(|0^{+}|)<\infty$, $w(x_{i},\epsilon_{i})\leq r^{\prime}(|0^{+}|)$ by Assumption 1 and $w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i})$ is Lipschitz continuous for $x\in\mathcal{L}(F^{0})$ and $\epsilon_{i}\in\mathbb{R}_{+}$ by Assumption 2.
+If $r^{\prime}(|0^{+}|)=\infty$, by (20), $\|x-\tfrac{1}{\beta}\nabla f(x)\|_{\infty}\leq C$, $\forall x\in\mathcal{L}(F^{0})$. When $w(x_{i},\epsilon_{i})>C/\lambda$, by the definition of $S^{x}$,
+$0\in\nabla_{i}f(x)+\beta(y_{i}-x_{i})+\lambda w(x_{i},\epsilon_{i})\partial|y_{i}|.$
+It is true that $S_{i}^{x}(x,\epsilon)=0$. Therefore, for any $x\in\mathcal{L}(F^{0})$ with $S_{i}^{x}(x,\epsilon)\neq 0$,
+(31) $r^{\prime}(|x_{i}|+\epsilon_{i})\leq C/\lambda\text{ \ and \ }|x_{i}|+\epsilon_{i}\geq\underline{x}:=(r^{\prime})^{-1}(C/\lambda)>0$
+by Assumption 1.
+It implies that we can equivalently write $S$ as
+$S_{i}^{x}(x,\epsilon)=\mathbb{S}_{\lambda u_{i}(x_{i},\epsilon_{i})/\beta}(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x)),$
+where we define $u:\mathbb{R}^{2n}\to\mathbb{R}^{n}$ with
+$u_{i}(x_{i},\epsilon_{i})=\begin{cases}w(x_{i},\epsilon_{i})=r^{\prime}(|x_{i}|+\epsilon_{i})&\text{ if }r^{\prime}(|0^{+}|)<\infty,\\ \min(C/\lambda,w(x_{i},\epsilon_{i}))=\min(C/\lambda,r^{\prime}(|x_{i}|+\epsilon_{i}))&\text{ if }r^{\prime}(|0^{+}|)=\infty.\end{cases}$
+Therefore, $u(\cdot)$ is also Lipschitz continuous for $x\in\mathcal{L}(F^{0})$ and $\epsilon\in\mathbb{R}^{n}_{+}$.
+Let $L_{r}$ be its Lipschitz constant, meaning for any $\hat{x},\tilde{x}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$
+(32) $\|u(\hat{x},\hat{\epsilon})-u(\tilde{x},\tilde{\epsilon})\|\leq L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|.$
 A natural choice is $L_{r}=r^{\prime\prime}(|0^{+}|)$ if $r^{\prime}(|0^{+}|)<\infty$ and $L_{r}=r^{\prime\prime}(|\underline{x}|)$ if $r^{\prime}(|0^{+}|)=\infty$.
-Next we prove that $S$ is Lipschitz continuous. It follows that
+Next, we prove that $S$ is Lipschitz continuous. It follows that
 ${\|}S^{x}(\hat{x},\hat{\epsilon})-S^{x}(\tilde{x},\tilde{\epsilon}){\|}=$ $\ {\|}\mathbb{S}_{\lambda\hat{u}/\beta}(\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x}))-\mathbb{S}_{\lambda\tilde{u}/\beta}(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x})){\|}$
-$\leq$ $\ \|\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x})-(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))\|+\|\hat{u}-\tilde{u}\|$
+$\leq$ $\ \|\hat{x}-\tfrac{1}{\beta}\nabla f(\hat{x})-(\tilde{x}-\tfrac{1}{\beta}\nabla f(\tilde{x}))\|+\tfrac{\lambda}{\beta}\|\hat{u}-\tilde{u}\|$
 $\leq$ $\ \|\hat{x}-\tilde{x}\|+\tfrac{1}{\beta}\|\nabla f(\hat{x})-\nabla f(\tilde{x}))\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
-$\leq$ $\ (1+\tfrac{1}{\beta}L_{f})\|\hat{x}-\tilde{x}\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
-$\leq$ $\ (1+\tfrac{1}{\beta}L_{f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|,$
+$\leq$ $\ (1+\tfrac{1}{\beta}L_{\nabla f})\|\hat{x}-\tilde{x}\|+\tfrac{\lambda}{\beta}L_{r}\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|$
+$\leq$ $\ (1+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|,$
 where the first inequality is determined by Lemma A.1 and the second is by (21) and (32).
-Therefore, $\|\alpha(\hat{x}-S^{x}(\hat{x},\hat{\epsilon}))-\alpha(\tilde{x}-S^{x}(\tilde{x},\tilde{\epsilon}))\|\leq\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))$ holds for any $\hat{y},\tilde{y}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$.
-Notice that $T=\begin{bmatrix}x\\ \epsilon\end{bmatrix}+\alpha(\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}-\begin{bmatrix}x\\ \epsilon\end{bmatrix})$, Since $\begin{bmatrix}S^{x}(x,\epsilon)\\ \mu\epsilon\end{bmatrix}$ is Lipschitz with constant $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$.
-Applying [23, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is invertible and a lipeomorphism.
-∎
-Now we are ready to apply Corollary 1 to derive the following global escape theorem.
-Theorem 15.
-(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumption 1-3 satisfied.
-Suppose $10<\alpha<$ 1, $\beta>\alpha L_{f}/2$, $\alpha(2+\frac{1}{\beta}(L_{f}+\lambda L_{r}))<1$ and $\alpha<\beta/\rho$.
+Therefore, for any $\hat{x},\tilde{x}\in\mathcal{L}(F^{0})$ and $\hat{\epsilon},\tilde{\epsilon}\in\mathbb{R}^{n}_{+}$,
+$\|\alpha(I-S)(\hat{x},\hat{\epsilon})-\alpha(I-S)(\tilde{x},\tilde{\epsilon})\|$
+$\leq$ $\ \alpha\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|+\alpha\|S^{x}(\hat{x},\hat{\epsilon})-S^{x}(\tilde{x},\tilde{\epsilon})\|+\alpha\|S^{\epsilon}(\hat{x},\hat{\epsilon})-S^{\epsilon}(\tilde{x},\tilde{\epsilon})\|$
+$\leq$ $\ \alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r})\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|+\alpha\mu\|\hat{\epsilon}-\tilde{\epsilon}\|$
+$\leq$ $\ \alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r}+\mu)\|(\hat{x};\hat{\epsilon})-(\tilde{x};\tilde{\epsilon})\|.$
+Notice that $T=I+\alpha(S-I)$.
+Since $\alpha(I-S)$ is Lipschitz with constant $\alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r}+\mu)<1$.
+Applying [15, Lemma 2.13], we directly derive that $T=I+\alpha(S-I)$ is invertible and a lipeomorphism.
+Now we are ready to apply Corollary 2.15 to derive the following global escape theorem.
+Theorem 3.13.
+(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{1}$ algorithm with Assumptions 1-3 being satisfied.
+Suppose $10<\alpha<$ 1, $\beta>\alpha L_{\nabla f}/2$, $\alpha(2+\tfrac{1}{\beta}L_{\nabla f}+\tfrac{\lambda}{\beta}L_{r}+\mu)<1$ and $\alpha<\beta/\rho$.
 Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 2.
 Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
-Proof.
-It follows directly from Theorem 13 that the map $T$ is local $C^{1}$ smooth around any stationary points, and all stationary points $x^{*}$ are unstable fixed points.
-By applying Proposition 14, $T$ is a lipeomorphism.
-By Corollary 1, the set of initial conditions attracted by such fixed points has a zero Lebesgue measure.
-∎
-We illustrate this property by applying DIRL1 to Example 1.
-We test $\alpha=0.1$ and $0.9$ with different initial points: $(3,3)$ representing an initial point far from the global minimizer and $(0,0.043)$ very close to the saddle point.
-The trajectories are shown in Figure 3.
-In four cases, the algorithm always converges to the global minimizer $(0,1)$.
-(a) $\alpha=0.9$ and initial point $(3,3)$.
-(b) $\alpha=0.9$ and initial point $(0,0.043)$.
-(c) $\alpha=0.1$ and initial point $(3,3)$.
-(d) $\alpha=0.1$ and initial point $(0,0.043)$.
-Figure 3: Trajectories of DIRL1.
+Proof 3.14.
+It follows directly from Theorem 3.9 that the map $T$ is local $C^{1}$ smooth around any stationary point, and that all stationary points are unstable fixed points.
+By applying Proposition 3.11, $T$ is a lipeomorphism.
+By Corollary 2.15, the set of initial conditions attracted by such fixed points has a zero Lebesgue measure.
 4 Damped Iterative Reweighted $\ell_{2}$ Algorithms
@@
 In this method, we write $h(\cdot)=r\circ\sqrt{(\cdot)}$, so that $s(x_{i})=h(x^{2}_{i})$.
-Since $\sqrt{\cdot}$ is concave on $\mathbb{R}_{+}$, $\sqrt{\lambda y_{1}+(1-\lambda)y_{2}}\geq\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}}$ for any $\lambda\in[0,1]$ and $y_{1},y_{2}\geq 0$.
-Since $r$ is increasing and concave, it follows that $h(\lambda y_{1}+(1-\lambda)y_{2})=r(\sqrt{\lambda y_{1}+(1-\lambda)y_{2}})\geq r(\lambda\sqrt{y_{1}}+(1-\lambda)\sqrt{y_{2}})\geq\lambda r(\sqrt{y_{1}})+(1-\lambda)r(\sqrt{y_{2}})=\lambda h(y_{1})+(1-\lambda)h(y_{2})$, meaning $h$ is also concave on $\mathbb{R}_{+}$.
-At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearize $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem (for brievty, we still use $S$ and $T$ for the maps),
-$y^{k}=S^{x}(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2},$
+Since $r$ is increasing and concave, it follows that $h$ is also concave on $\mathbb{R}_{+}$.
+At each iteration, after adding perturbation $\epsilon_{i}^{2}$ to each $h$ and linearizing $h$ at $(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}$, the ${\rm IRL}_{2}$ algorithm solves the following weighted $\ell_{2}$ regularized subproblem (for brevity, we still use $S$ and $T$ for the maps),
+$y^{k}=S^{x}(x^{k},\epsilon^{k})=\arg\min_{y}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}u(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2},$
 where
-$w(x_{i},\epsilon_{i})=h^{\prime}(x_{i}^{2}+\epsilon_{i}^{2})=\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{2\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}.$
-In particular,
-$w(x_{i},0)=\begin{cases}\frac{r^{\prime}(|x_{i}|)}{2|x_{i}|}&\text{ if }x_{i}\neq 0\\ \infty&\text{ if }x_{i}=0.\end{cases}$
+$u(x_{i},\epsilon_{i})=h^{\prime}(x_{i}^{2}+\epsilon_{i}^{2})=\frac{r^{\prime}(\sqrt{x_{i}^{2}+\epsilon_{i}^{2}})}{2\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}\ \text{ and }\ u(x_{i},0)=\begin{cases}\frac{r^{\prime}(|x_{i}|)}{2|x_{i}|}&\text{ if }x_{i}\neq 0\\ \infty&\text{ if }x_{i}=0.\end{cases}$
 The solution of the subproblem is given by
-$y^{k}_{i}=S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}w(x_{i}^{k},\epsilon_{i}^{k})}(x_{i}^{k}-\frac{1}{\beta}\nabla_{i}f(x^{k})).$ (33)
-If $w(x_{i}^{k},\epsilon_{i}^{k})=\infty$, we simply set $S_{i}^{x}(x^{k},\epsilon^{k})=0$.
+(33) $y^{k}_{i}=S_{i}^{x}(x,\epsilon):=\frac{1}{1+\frac{2\lambda}{\beta}u(x_{i}^{k},\epsilon_{i}^{k})}(x_{i}^{k}-\tfrac{1}{\beta}\nabla_{i}f(x^{k})).$
+If $u(x_{i}^{k},\epsilon_{i}^{k})=\infty$, we simply set $S_{i}^{x}(x^{k},\epsilon^{k})=0$.
 The damped iteratively reweighted $\ell_{2}$ algorithm, hereinafter named DIRL2 algorithm, is presented in Algorithm 2.
-1:Input $\beta>\alpha L_{f}/2$, $\alpha\in(0,1)$, $x_{0}$, and $\epsilon_{0}>0$.
+1:  Input $\beta>\alpha L_{\nabla f}/2$, $\alpha\in(0,1)$, $x_{0}$, $\epsilon_{0}>0$, and $\mu\in(0,1)$.
 2:  Initialize: set $k=0$.
 3:  repeat
-4:     Reweighting: $w(x_{i}^{k},\epsilon_{i}^{k})=h^{\prime}((x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2})$.
+4:     Reweighting: $u(x_{i}^{k},\epsilon_{i}^{k})=h^{\prime}((x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2})$.
 5:     Compute new iterate:
-6:     $y^{k}=\arg\min\limits_{y\in\mathbb{R}^{n}}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}w(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2}.$
+6:     $y^{k}=\arg\min\limits_{y\in\mathbb{R}^{n}}G(y;x^{k}):=\nabla f(x^{k})^{T}y+\tfrac{\beta}{2}\|x^{k}-y\|^{2}+\lambda\sum_{i=1}^{n}u(x^{k}_{i},\epsilon_{i}^{k})y_{i}^{2}.$
 7:     $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$.
@@
 Algorithm 2 Damped Iteratively Reweighted $\ell_{2}$ Algorithm (${\rm DIRL}_{2}$)
-The fixed point iteration of DIRL2 is then given as $(x^{k+1},\epsilon^{k+1})=T(x^{k},\epsilon^{k})$ with $T=(1-\alpha)I+\alpha S$ and
-$T(x,\epsilon):=\begin{bmatrix}(1-\alpha)x+\alpha S^{x}(x,\epsilon)\\ (1-\alpha(1-\mu))\epsilon\end{bmatrix}.$ (34)
-We have the following result to connect the stationary point of (P) and the fixed point of $T$.
-Proposition 16.
-If $x^{*}$ is a stationary point of (P), then $x^{*}=S^{x}(x^{*},0)$ and $(x^{*},0)=T(x^{*},0)$.
+The fixed point iteration of DIRL2 is then given as $(x^{k+1};\epsilon^{k+1})=T(x^{k},\epsilon^{k})$ with $T=(1-\alpha)I+\alpha S$ and
+(34) $T(x,\epsilon):=\begin{bmatrix}(1-\alpha)x+\alpha S^{x}(x,\epsilon)\\ (1-\alpha(1-\mu))\epsilon\end{bmatrix}.$
+We have the following result to connect the stationary point of (P) with the fixed point of $T$.
+Proposition 4.1.
+If $x^{*}$ is a stationary point of (P), then $x^{*}=S^{x}(x^{*},0)$ and $(x^{*};0)=T(x^{*},0)$.
 Conversely, for any $x^{*}$ satisfying $x^{*}=S^{x}(x^{*},0)$, if $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, then $x^{*}$ is a stationary point of (P).
-Proof.
+Proof 4.2.
 Suppose $x^{*}$ is a stationary point.
 It follows from (5) that for $i\in\mathcal{I}^{*}$, $x_{i}^{*}=x_{i}^{*}-\tfrac{1}{\beta}(\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*}))$ implying $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$.
-Therefore, $x_{i}^{*}=S_{i}(x^{*},0)$ since $w(x_{i}^{*},0)=\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}$.
+Therefore, $x_{i}^{*}=S_{i}^{x}(x^{*},0)$ since $u(x_{i}^{*},0)=\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}$.
 As for $i\in\mathcal{J}^{*}$, $w(0,0)=\infty$, implying that $0=S_{i}(x^{*},0)$.
@@
 Now consider $x^{*}=S^{x}(x^{*},0)$.
-If $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, we only have to verify (6) is satisfied by $x^{*}$.
-It follows from $x^{*}_{i}=S_{i}(x^{*},0),i\in\mathcal{I}^{*}$ that $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$, implying $\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*})=0$, completing the proof.
-∎
-The converse statement of Proposition 16 may not be true when $\mathcal{J}^{*}\neq\emptyset$ and $r^{\prime}(|0^{+}|)<\infty$.
-Since for any $x^{*}$ that satisfies $x_{i}^{*}=S_{i}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it is naturally true that $w(0,0)=\infty$ and $S_{i}(x^{*},0)=0$.
+If $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$, it suffices to verify (5) is satisfied by $x^{*}$.
+It follows from $x^{*}_{i}=S_{i}^{x}(x^{*},0),i\in\mathcal{I}^{*}$ that $x_{i}^{*}\left(1+\tfrac{2\lambda}{\beta}\tfrac{r^{\prime}(|x_{i}^{*}|)}{2|x_{i}^{*}|}\right)=x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*})$, implying $\nabla_{i}f(x^{*})+\lambda r^{\prime}(|x_{i}^{*}|)\text{sign}(x_{i}^{*})=0$, completing the proof.
+Remark 4.3.
+The converse statement of Proposition 4.1 may not be true when $\mathcal{J}^{*}\neq\emptyset$ and $r^{\prime}(|0^{+}|)<\infty$.
+Since for any $x^{*}$ that satisfies $x_{i}^{*}=S_{i}^{x}(x^{*},0)$, $i\in\mathcal{I}^{*}$, it is naturally true that $u(0,0)=\infty$ and $S_{i}^{x}(x^{*},0)=0$.
 However, there is no guarantee that $x^{*}$ satisfies (6).
@@
 However, it should be noted that the definition of $F(x,\epsilon)$ is different from that of the DIRL1 algorithm.
-Proposition 17.
-Suppose Assumption 1-3 hold.
-Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 10 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{f}}{2}$.
+Proposition 4.4.
+Suppose Assumptions 1-3 hold.
+Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by Algorithm 2 with $\alpha\in(0,1)$ and $\frac{\beta}{\alpha}\geq\frac{L_{\nabla f}}{2}$.
 It holds that {$F(x^{k},\epsilon^{k})$} is monotonically decreasing and the reduction satisfies
-$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
-Proof.
+$F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
+Moreover, $\lim\limits_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$ and $\{x^{k}\}\subset\mathcal{L}(F^{0})$.
+Proof 4.5.
 It follows from the concavity of $h$ that for $i=1,...,n$,
@@
 which is equivalent to
-$r(\sqrt{(x_{i}^{k+1})^{2}+(\epsilon_{i}^{k})^{2}})\leq r(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})+w(x_{i}^{k},\epsilon^{k}_{i})((x_{i}^{k+1})^{2}-(x_{i}^{k})^{2}).$
+$r(\sqrt{(x_{i}^{k+1})^{2}+(\epsilon_{i}^{k})^{2}})\leq r(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})+u(x_{i}^{k},\epsilon^{k}_{i})((x_{i}^{k+1})^{2}-(x_{i}^{k})^{2}).$
 Following the same argument for (23)-(25), we have
-$\Delta F(x^{k+1},\epsilon^{k+1})\geq\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{f}}{2}\|x^{k}-x^{k+1}\|^{2}.$ (35)
-For each subproblem, $y^{k}$ satisfies the optimal condition
-$\nabla f(x^{k})+2\lambda W(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})=0.$ (36)
-Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimal condition (27), the following holds
-$\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$ (37)
-$=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})((x_{i}^{k})^{2}-(x_{i}^{k+1})^{2})-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
-$\geq$ $\ \nabla f(x^{k})^{T}(\alpha x^{k}-\alpha y^{k})+\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})(\alpha(x_{i}^{k})^{2}-\alpha(y_{i}^{k})^{2})-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
-$\geq$ $\ \alpha[\nabla f(x^{k})^{T}(x^{k}-y^{k})+2\lambda\sum\limits_{i=1}^{n}w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}(x_{i}^{k}-y_{i}^{k})]-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
-$=$ $\ \alpha[\nabla f(y^{k})+2\lambda W(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})]^{T}(x^{k}-y^{k})+\frac{\alpha\beta(2-\alpha)}{2}\|x^{k}-y^{k}\|^{2}$
-$=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2},$
-where the first inequality is by Lemma A.2 and the second is by the convexity of $(\cdot)^{2}$. Combining (26) and (37), we get
-$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
+(35) $\Delta F(x^{k+1},\epsilon^{k+1})\geq\Delta G(x^{k+1};x^{k},\epsilon^{k})+\frac{\beta-L_{\nabla f}}{2}\|x^{k}-x^{k+1}\|^{2}.$
+For each subproblem, $y^{k}$ satisfies the optimality condition
+(36) $\nabla f(x^{k})+2\lambda U(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})=0.$
+where $U(x^{k},\epsilon^{k}):={\rm diag}(u(x_{1}^{k},\epsilon_{1}^{k}),...,u(x_{n}^{k},\epsilon_{n}^{k}))$.
+Consequently, by $x^{k+1}=(1-\alpha)x^{k}+\alpha y^{k}$ and the optimality condition (26), the following holds
+(37) $\ \Delta G(x^{k+1};x^{k},\epsilon^{k})$
+$=$ $\ \nabla f(x^{k})^{T}(x^{k}-x^{k+1})+\lambda\sum\limits_{i=1}^{n}u(x_{i}^{k},\epsilon_{i}^{k})((x_{i}^{k})^{2}-(x_{i}^{k+1})^{2})-\frac{\beta}{2}\|x^{k}-x^{k+1}\|^{2}$
+$\geq$ $\ \nabla f(x^{k})^{T}(\alpha x^{k}-\alpha y^{k})+\lambda\sum\limits_{i=1}^{n}u(x_{i}^{k},\epsilon_{i}^{k})(\alpha(x_{i}^{k})^{2}-\alpha(y_{i}^{k})^{2})-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
+$\geq$ $\ \alpha[\nabla f(x^{k})^{T}(x^{k}-y^{k})+2\lambda\sum\limits_{i=1}^{n}u(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}(x_{i}^{k}-y_{i}^{k})]-\frac{\beta\alpha^{2}}{2}\|x^{k}-y^{k}\|^{2}$
+$=$ $\ \alpha[\nabla f(y^{k})+2\lambda U(x^{k},\epsilon^{k})y^{k}+\beta(y^{k}-x^{k})]^{T}(x^{k}-y^{k})+\frac{\alpha\beta(2-\alpha)}{2}\|x^{k}-y^{k}\|^{2}$
+$=$ $\ (\frac{\beta}{\alpha}-\frac{\beta}{2})\|x^{k}-x^{k+1}\|^{2}.$
+Here the first inequality is by $(1-\alpha)(x_{i}^{k})^{2}-\alpha(y_{i}^{k})^{2}\geq(\alpha x_{i}^{k}+(1-\alpha)y_{i}^{k})^{2}$ and the second is by the convexity of $(\cdot)^{2}$.
+Combining (25) and (37), we get
+$\Delta F(x^{k+1},\epsilon^{k+1})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\|x^{k}-x^{k+1}\|^{2}.$
 Replacing $k$ with $t$ and summing up from $t=0$ to $k-1$, we have
-$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
-It follows that $\{(x^{k},\epsilon^{k})\}\in\mathcal{L}(F^{0})$ since $F(x^{k},\epsilon^{k})\leq F(x^{0},\epsilon^{0})$.
+$\sum\limits_{t=0}^{k-1}[F(x^{t},\epsilon^{t})-F(x^{t+1},\epsilon^{t+1})]=F(x^{0},\epsilon^{0})-F(x^{k},\epsilon^{k})\geq(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{k-1}\|x^{t}-x^{t+1}\|^{2}.$
+It follows that $F(x^{k})\leq F(x^{k},\epsilon^{k})\leq F^{0}$, which implies that $\{x^{k}\}\in\mathcal{L}(F^{0})$.
 By Assumption 3, it is bounded and $\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})>-\infty$.
 Taking $k\rightarrow\infty$, we have
-$(\frac{\beta}{\alpha}-\frac{L_{f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
-which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|^{2}=0$. Moreover, from the boundedness of $\{x^{k}\}$, there must exist $C>0$ such that (23) holds for any $k\in\mathbb{N}$. ∎
-With the monotonicity of $F$, we derive the global convergence of the ${\rm DIRL}_{1}$ method.
-Theorem 18.
-(Global convergence) Suppose Assumption 13 hold, and $r^{\prime}(|0^{+}|)=\infty$ or $\mathcal{J}^{*}=\emptyset$.
+$(\frac{\beta}{\alpha}-\frac{L_{\nabla f}}{2})\sum\limits_{t=0}^{\infty}\|x^{t}-x^{t+1}\|^{2}\leq F(x^{0},\epsilon^{0})-\mathop{\lim\inf}\limits_{k\rightarrow\infty}F(x^{k},\epsilon^{k})<\infty.$
+which implies $\lim\limits_{k\rightarrow\infty}\|x^{k}-x^{k+1}\|=0$.
+With the monotonicity of $F$, we derive the global convergence for the ${\rm DIRL}_{2}$ method.
+Theorem 4.6.
+(Global convergence) Suppose that Assumptions 13 hold, and $\mathcal{J}^{*}=\emptyset$ or $r^{\prime}(|0^{+}|)=\infty$.
 Let $\{(x^{k},\epsilon^{k})\}$ being generated by ${\rm DIRL}_{2}$ and $\Omega$ be the set of the limit points of $\{x^{k}\}$.
-Then any $x^{*}\in\Omega$ is first-order optimal for (P).
-Proof.
+Then any $x^{*}\in\Omega$ is a stationary point for (P).
+Proof 4.7.
 Firstly, the boundedness of $\mathcal{L}(F^{0})$ from Assumption 3 implies $\Omega\neq\emptyset$.
-Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{S}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$.
+Then let $x^{*}$ be a limit point of $\{x^{k}\}$ with subsequence $\{x^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ and $\{y^{k}\}_{\mathcal{S}}\rightarrow x^{*}$ since $y^{k}=\frac{1}{\alpha}(x^{k}-x^{k+1})+x^{k}$.
 The optimality condition of the $k$th subproblem implies
-$-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda w(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})}{\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}}}y_{i}^{k}.$ (38)
+(38) $-\beta(y_{i}^{k}-x_{i}^{k})=\nabla_{i}f(x^{k})+2\lambda u(x_{i}^{k},\epsilon_{i}^{k})y_{i}^{k}=\nabla_{i}f(x^{k})+\frac{r^{\prime}(\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}})}{\sqrt{(x_{i}^{k})^{2}+(\epsilon_{i}^{k})^{2}}}y_{i}^{k}.$
 As $x_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$, $y_{i}^{k}\overset{\mathcal{S}}{\to}x_{i}^{*}$ and $\epsilon_{i}^{k}\to 0$, $\lim\limits_{k\in\mathcal{S}}\frac{y_{i}^{k}}{\sqrt{x_{i}^{2}+\epsilon_{i}^{2}}}=\text{sign}(x_{i}^{*})$ for $i\in\mathcal{I}^{*}$.
-For $i\in\mathcal{J}^{*}$, there is nothing to prove since $|\nabla_{i}f(x^{*})|\leq r^{\prime}(|0^{+}|)$ naturally holds. In general, $x^{*}$ is stationary. ∎
+Therefore, (5) is satisfied.
+If $\mathcal{J}^{*}=\emptyset$, there is nothing to prove.
+If not, for $i\in\mathcal{J}^{*}$, there is nothing to prove since $|\nabla_{i}f(x^{*})|\leq r^{\prime}(|0^{+}|)$ naturally holds.
+In general, $x^{*}$ is a stationary point.
 4.2 Avoidance of strictly saddle points
@@
 There is no guarantee that the iteration map is $C^{1}$-smooth without this property.
-The following additional assumption is needed to guarantee the local smoothness.
-One can verify that most sparse nonconvex penalty in Table 1 like LPN, TAN, FRA, EXP and LOG are satisfied.
+The following additional assumption is needed to guarantee local smoothness, which is satisfied by the $\ell_{p}$ regularization.
 Assumption 4.
-$\lim\limits_{z\rightarrow 0}r^{\prime}(|z|)=\infty\ \text{ and }\ \lim\limits_{z\rightarrow 0}z\frac{r^{\prime\prime}(|z|)}{r^{\prime}(|z|)^{2}}\to 0.$
-To show the smoothness of $S$, we first calculate the Jacobian of $S$. Define
-$g(z):=\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|z|)}{|z|}}\ \text{ with }g(0):=0.$
+$\lim\limits_{z\rightarrow 0^{+}}r^{\prime}(z)=\infty\ \text{ and }\ \lim\limits_{z\rightarrow 0^{+}}z\frac{r^{\prime\prime}(z)}{r^{\prime}(z)^{2}}\to 0.$
+To show the smoothness of $S$, we first calculate the Jacobian of $S$. Define the auxiliary function $g:\mathbb{R}_{+}\to\mathbb{R}$:
+$g(z):=\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(z)}{z}}\ \text{ with }g(0):=0.$
 We can write $S_{i}^{x}(x,\epsilon)=g(z)(x_{i}-\tfrac{1}{\beta}\nabla_{i}f(x))$ with $z=\sqrt{x^{2}+\epsilon^{2}}$. It follows that
-$g^{\prime}(z)=\begin{cases}-\tfrac{\lambda}{\beta}\frac{r^{\prime\prime}(|z|)z-\text{sign}(z)r^{\prime}(|z|)}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}},&\text{\ for }z\neq 0,\\ \lim\limits_{\delta\to 0}\left(\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|\delta|)}{|\delta|}}-0\right)\frac{1}{\delta}=0,&\text{\ for }z=0.\end{cases}$ (39)
+(39) $g^{\prime}(z)=\begin{cases}-\tfrac{\lambda}{\beta}\frac{r^{\prime\prime}(z)z-r^{\prime}(z)}{(z+\tfrac{\lambda}{\beta}{r^{\prime}(z)})^{2}},&\text{\ for }z\neq 0,\\ \lim\limits_{\delta\to 0}\left(\frac{1}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|\delta|)}{|\delta|}}-0\right)\frac{1}{\delta}=0,&\text{\ for }z=0.\end{cases}$
 It holds that
-$\lim\limits_{z\to 0}g^{\prime}(z)=$ $-\tfrac{\lambda}{\beta}\lim\limits_{z\to 0}\left(\frac{r^{\prime\prime}(|z|)z}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}}-\frac{\text{sign}(z)r^{\prime}(|z|)}{(|z|+\tfrac{\lambda}{\beta}{r^{\prime}(|z|)})^{2}}\right)$
-$=$ $-\lim\limits_{z\to 0}\left(\tfrac{\beta}{\lambda}\frac{r^{\prime\prime}(|z|)z}{{r^{\prime}(|z|)}^{2}}-\text{sign}(z)\frac{|z|}{r^{\prime}(|z|)}\right)=0=g^{\prime}(0)$
+$\lim\limits_{z\to 0}g^{\prime}(z)=$ $-\tfrac{\lambda}{\beta}\lim\limits_{z\to 0}\left(\frac{r^{\prime\prime}(z)z}{(z+\tfrac{\lambda}{\beta}{r^{\prime}(z)})^{2}}-\frac{r^{\prime}(z)}{(z+\tfrac{\lambda}{\beta}{r^{\prime}(z)})^{2}}\right)$
+$=$ $-\lim\limits_{z\to 0}\left(\tfrac{\beta}{\lambda}\frac{r^{\prime\prime}(z)z}{{r^{\prime}(z)}^{2}}-\frac{z}{r^{\prime}(z)}\right)=0=g^{\prime}(0)$
 by Assumption 4. Therefore, $g$ is continuously differentiable in $\mathbb{R}$.
-We have the following result for $S$ and $T$.
-Theorem 19.
-(Local smoothness and Lipschitz continuity) Suppose Assumption 1-4 hold. The following statements hold:
+We have the following result for maps $S$ and $T$.
+Theorem 4.8.
+(Local smoothness and Lipschitz continuity) Suppose Assumptions 1-4 hold and $x^{*}$ is a stationary point of (P). Then the following statements hold:
   1. (i)
     $S$ is $C^{1}$ smooth and Lipschitz continuous on $\mathcal{L}(F^{0})$ with constant $L_{S}$; moreover
-    $DS(x^{*},\bm{0})=\ \begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&,\bm{0}\\ 0&0&0\\ 0&0&\mu I\end{bmatrix},$ (40)
-    where $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}W_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$.
+    (40) $DS(x^{*},\boldsymbol{0})=\ \begin{bmatrix}H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&\mu I\end{bmatrix},$
+    where $H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{1}{\beta}(I+\tfrac{\lambda}{\beta}U_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$.
   2. (ii)
     $T$ is a lipeomorphism for $\alpha\in(0,\frac{1}{1+L_{S}})$ with
-    $DT(x^{*},\bm{0})=\ \begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&H_{\mathcal{I}^{*}\mathcal{J}^{*}}&\bm{0}\\ \bm{0}&(1-\alpha)I&\bm{0}\\ \bm{0}&\bm{0}&(1-\alpha(1-\mu))I\end{bmatrix}.$
-Proof.
-(i) We simply calculate its partial derivative and verify that it is continuous at any point $(x^{*},\epsilon^{*})\in\mathbb{R}^{2n}$.
+    $DT(x^{*},\boldsymbol{0})=\ \begin{bmatrix}(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}&\alpha H_{\mathcal{I}^{*}\mathcal{J}^{*}}&\boldsymbol{0}\\ \boldsymbol{0}&(1-\alpha)I&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{0}&(1-\alpha(1-\mu))I\end{bmatrix}.$
+Proof 4.9.
+(i) We simply calculate its partial derivative and verify that it is continuous at any point $(x^{*};\epsilon^{*})\in\mathbb{R}^{2n}$.
+The smoothness of $S^{\epsilon}$ is obvious.
+We only verify the smoothness of $S^{x}$.
 Consider $i$ satisfying $z_{i}^{*}:=\sqrt{(x_{i}^{*})^{2}+(\epsilon_{i}^{*})^{2}}>0$. In this case, $S$ is smooth with respect to $x_{i}$ and $\epsilon_{i}$, and
-$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ g(z^{*}_{i})(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))+\tfrac{x_{i}^{*}}{z_{i}^{*}}g^{\prime}(z_{i}^{*})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (41)
-$\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ -g(z^{*}_{i})\tfrac{1}{\beta}\nabla^{2}_{ij}f(x^{*}),\quad j\neq i;$ (42)
-$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ \tfrac{\epsilon_{i}^{*}}{z_{i}^{*}}g^{\prime}(z^{*}_{i})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$ (43)
-$\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*})=$ $\ 0,\quad j\neq i.$ (44)
-Consider $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$. It follows from $S_{i}^{x}(x^{*},\bm{0})=0$ and $r^{\prime}(|0^{+}|)=\infty$ that
-$\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$ (45)
+(41) $\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ g(z^{*}_{i})(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))+\tfrac{x_{i}^{*}}{z_{i}^{*}}g^{\prime}(z_{i}^{*})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$
+(42) $\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ -g(z^{*}_{i})\tfrac{1}{\beta}\nabla^{2}_{ij}f(x^{*}),\quad j\neq i;$
+(43) $\frac{\partial}{\partial\epsilon_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \tfrac{\epsilon_{i}^{*}}{z_{i}^{*}}g^{\prime}(z^{*}_{i})(x_{i}^{*}-\tfrac{1}{\beta}\nabla_{i}f(x^{*}));$
+(44) $\frac{\partial}{\partial\epsilon_{j}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ 0,\quad j\neq i.$
+Consider $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$.
+It follows from $S_{i}^{x}(x^{*},\boldsymbol{0})=0$ and $r^{\prime}(|0^{+}|)=\lim\limits_{\delta\to 0}u(\delta,0)=\lim\limits_{\delta\to 0}u(0,\delta)=\infty$ that
+(45) $\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}u(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{i})]/\delta$
 $=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta$
 $=$ $-\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}g^{\prime}(0)=0,$
-$\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{j})]/\delta$ (46)
+(46) $\frac{\partial}{\partial x_{j}}S^{x}_{i}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}u(\delta,0)}[\delta-\tfrac{1}{\beta}\nabla_{i}f(x^{*}+\delta e_{j})]/\delta$
 $=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta=0,\quad j\neq i;$
-$\frac{\partial}{\partial\epsilon_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}w(0,\delta)}[-\tfrac{1}{\beta}\nabla_{i}f(x^{*})]/\delta$ (47)
+(47) $\frac{\partial}{\partial\epsilon_{i}}S_{i}^{x}(x^{*},\epsilon^{*})=$ $\ \lim\limits_{\delta\rightarrow 0}\frac{1}{1+\frac{2\lambda}{\beta}u(0,\delta)}[-\tfrac{1}{\beta}\nabla_{i}f(x^{*})]/\delta$
 $=$ $\ -\tfrac{1}{\beta}\nabla_{i}f(x^{*})\lim\limits_{\delta\rightarrow 0}[g(\delta)-g(0)]/\delta=0;$
@@
 Now we verify that the Jacobian is continuous at $(x^{*},\epsilon^{*})$.
-It suffices to show this for any $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$; indeed, it holds that It follows from (41), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g(0)=0,g^{\prime}(0)=0$ that
-$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(x^{*},\epsilon^{*}).$
+It suffices to show this for any $i$ satisfying $x_{i}^{*}=0$ and $\epsilon_{i}^{*}=0$.
+It follows from (41), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g(0)=0,g^{\prime}(0)=0$ that
+$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{i}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{i}}S_{i}^{x}(0,0).$
 It follows from (42) and $g(0)=0$ that
-$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{j}}S_{i}^{x}(x^{*},\epsilon^{*}).$
+$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial x_{j}}S_{i}^{x}(x,\epsilon)=0=\frac{\partial}{\partial x_{j}}S_{i}^{x}(0,0).$
 It follows from (43), the fact that $\frac{x_{i}}{z_{i}}\leq 1$ and $g^{\prime}(0)=0$ that
-$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
+$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{i}}S_{i}^{\epsilon}(0,0).$
 It follows from (44) that
-$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x^{*},\epsilon^{*}).$
+$\lim_{x\to x^{*},\epsilon\to\epsilon^{*}}\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(x,\epsilon)=0=\frac{\partial}{\partial\epsilon_{j}}S_{i}^{\epsilon}(0,0).$
 Overall, we have shown that $S$ is continuously differentiable.
-To derive the matrix form of the Jacobian, it follows from (41) that for $i\in\mathcal{I}^{*}$,
-$h_{ii}:=$ $\ \frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})$
+To derive (40), now suppose $x^{*}$ is a stationary point for (P) and $\epsilon^{*}=\boldsymbol{0}$.
+For $i\in\mathcal{I}^{*}$, (5) is equivalent to $x_{i}^{*}-\frac{1}{\beta}\nabla_{i}f(x^{*})=x_{i}^{*}+\frac{\lambda}{\beta}\frac{x_{i}^{*}}{z_{i}^{*}}r^{\prime}(|x^{*}_{i}|)$ and $x_{i}^{*}/z_{i}^{*}=\text{sign}(x_{i}^{*})$.
+It follows from (41) that for $i\in\mathcal{I}^{*}$,
+(49) $h_{ii}=$ $\ \frac{\partial}{\partial x_{j}}S^{x}_{i}(x^{*},\boldsymbol{0})$
 $=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(x^{*}_{i}))\tfrac{\lambda}{\beta}\left(|x^{*}_{i}|r^{\prime\prime}(|x^{*}_{i}|)-r^{\prime}(|x^{*}_{i}|)\right)}{\left(|x^{*}_{i}|+\tfrac{\lambda}{\beta}r^{\prime}(|x^{*}_{i}|)\right)^{2}},$
-$=$ $\ \frac{(1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*}))}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}-\frac{\tfrac{\lambda}{\beta}\left(r^{\prime\prime}(|x^{*}_{i}|)-\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}\right)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x_{i}^{*}|}}$
-$=$ $\ \frac{1-\tfrac{1}{\beta}\nabla^{2}_{ii}f(x^{*})-\tfrac{\lambda}{\beta}r^{\prime\prime}(|x_{i}^{*}|)+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
-$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x_{i}^{*}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}$
+$=$ $\ 1-\tfrac{1}{\beta}\frac{\nabla^{2}_{ii}f(x^{*})+\lambda r^{\prime\prime}(|x_{i}^{*}|)}{1+\tfrac{\lambda}{\beta}\frac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}.$
 It follows from (42) that for $j\neq i$,
-$h_{ij}:=\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\bm{0})=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}.$
-We then combine this with (43)–(48) to derive (40) of the Jacobian $DS(x^{*},\bm{0})$. The Lipschitz continuity of $S$ trivially leads to the boundedness of $\mathcal{L}(F^{0})$.
-(ii) This is straightforward by applying [23, Lemma 2.13 and 2.14]. The Jacobian $DT(x^{*},\bm{0})$ is obtained directly by (i). ∎
+(50) $h_{ij}:=\frac{\partial}{\partial x_{j}}S_{i}(x^{*},\boldsymbol{0})=-\tfrac{1}{\beta}\frac{\nabla^{2}_{ij}f(x^{*})}{1+\tfrac{\lambda}{\beta}\tfrac{r^{\prime}(|x^{*}_{i}|)}{|x^{*}_{i}|}}.$
+We then combine this with (43)–(50) to derive (40). The boundedness of $\mathcal{L}(F^{0})$ trivially leads to the Lipschitz continuity of $S$.
+(ii) This is straightforward by applying [15, Lemma 2.13 and 2.14]. The Jacobian $DT(x^{*},\boldsymbol{0})$ is obtained directly by (i).
 Now we show that every strict saddle point of (P) is an unstable fixed point of $T$.
-Theorem 20.
-(Unstable Fixed Point) Suppose Assumption 1-4 hold.
-Let $x^{*}$ be a stationary point of (P).
-If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=\bm{0}$ is an unstable fixed point of $T$.
-Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible.
-Proof.
-It follows from Theorem 19(ii) that we only have to check the eigenvalues of
-$(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}),$ (49)
-where we have $W^{*}_{ii}=w(x_{i}^{*},0)<\infty,i\in\mathcal{I}^{*}$.
-If $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$, it follows from Lemma [49, Proposition 6.1] that $\lambda_{\min}((I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$.
-Hence, $DT(x^{*},\bm{0})$ has an eigenvalue greater than 1, meaning $(x^{*},\bm{0})$ is the unstable fixed point of $T$.
-Moreover, since $\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}W^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$, it is clear that for $\alpha<\beta/\rho$, $DT(x^{*},\bm{0})$ is invertible.
-∎
-Now we can derive the global escape property of ${\rm DIRL}_{2}$ algorithm.
-Theorem 21.
-(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{2}$ algorithm with Assumption 1-4.
-Let the condition of Proposition 19 holds and $\alpha<\beta/\rho$.
+Theorem 4.10.
+(Unstable Fixed Point) Suppose Assumptions 1-4 hold.
+If $x^{*}$ is a strict saddle point of (P), then the corresponding $(x^{*},\epsilon^{*})$ with $\epsilon^{*}=\boldsymbol{0}$ is an unstable fixed point of $T$.
+Moreover, for $\alpha<\beta/\rho$, $DT(x^{*},\boldsymbol{0})$ is invertible.
+Proof 4.11.
+By Theorem 4.8(ii), it suffices to check the eigenvalues of
+(51) $(1-\alpha)I+\alpha H_{\mathcal{I}^{*}\mathcal{I}^{*}}=I-\tfrac{\alpha}{\beta}(I+\tfrac{\lambda}{\beta}U^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}),$
+where we have $U^{*}_{ii}=u(x_{i}^{*},0)<\infty,i\in\mathcal{I}^{*}$.
+If $\lambda_{\min}(\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$, it follows from Lemma [36, Proposition 6.1] that $\lambda_{\min}((I+\tfrac{\lambda}{\beta}U^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*}))<0$.
+Hence, $DT(x^{*},\boldsymbol{0})$ has an eigenvalue greater than 1, meaning $(x^{*},\boldsymbol{0})$ is the unstable fixed point of $T$.
+Moreover, since $\|(I+\tfrac{\lambda}{\beta}U^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|(I+\tfrac{\lambda}{\beta}U^{*}_{\mathcal{I}^{*}\mathcal{I}^{*}})^{-1}\|\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\|\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})\|\leq\rho$ by (22) and $\nabla^{2}_{\mathcal{I}^{*}\mathcal{I}^{*}}F(x^{*})$ does not have 0 eigenvalue by the strict saddle property, we know by (22) that $DT(x^{*},\epsilon^{*})$ is invertible for $\alpha<\beta/\rho$.
+Now we can derive the global escape property for the ${\rm DIRL}_{2}$ algorithm.
+Theorem 4.12.
+(Global Escape) Consider problem (P) with strict saddle property and the ${\rm DIRL}_{2}$ algorithm with Assumptions 1-4.
+Let $\alpha<\beta/\rho$ and $\alpha\in(0,\frac{1}{1+L_{S}})$ where $L_{S}$ is as defined in Theorem 4.8.
 Let $\{(x^{k},\epsilon^{k})\}$ be the sequence generated by the Algorithm 2.
 Then, the set of initial points attracted by strict saddle points has zero Lebesgue measure.
-Proof.
-It follows directly from Theorem 19 and 20 that: (i) map $T$ is local $C^{1}$-smooth around any stationary points; (ii) all stationary points $x^{*}$ are unstable fixed points; (iii) $T$ is a lipeomorphism.
-Then by Corollary 1, the set of initial conditions attracted by such fixed points has zero Lebesgue measure.
-∎
-We illustrate this property by applying DIRL2 to Example 1.
-We test $\alpha=0.1$ and $0.9$ with different initial points; first is $(3,3)$ representing an initial point far from the global minimizer and the second is $(0,0.043)$ very close to the saddle point.
-The trajectories are shown in Figure 4.
-In four cases, the algorithm always converges to the global minimizer $(0,1)$.
-(a) $\alpha=0.9$ and initial point $(3,3)$.
-(b) $\alpha=0.9$ and initial point $(0,0.043)$.
-(c) $\alpha=0.1$ and initial point $(3,3)$.
-(d) $\alpha=0.1$ and initial point $(0,0.043)$.
-Figure 4: Trajectories of DIRL2.
-Appendix
-Lemma A.1.
-Given $a,b,c\in\mathbb{R}^{n}_{+}$ with $a_{i}\leq b_{i}+c_{i}\quad\text{for all }i$. It holds that $\|a\|\leq\|b\|+\|c\|$.
-Proof.
-It follows that $a_{i}^{2}\leq b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2}$, implying $\sum_{i=1}^{n}a_{i}^{2}\leq\sum_{i=1}^{n}(b_{i}^{2}+2b_{i}c_{i}+c_{i}^{2})$, or equivalently, $\|a\|^{2}\leq\|b\|^{2}+2b^{T}c+\|c\|^{2}$.
-By the Cauchy-Schwarz inequality, $2\sum_{i=1}^{n}b_{i}c_{i}\leq 2\|b\|\|c\|$.
-Therefore, $\|a\|^{2}\leq(\|b\|+\|c\|)^{2}$, completing the proof.
-∎
-Lemma A.2.
-For $x,y\in\mathbb{R}$ and $\alpha\in(0,1]$, it holds that $x^{2}-((1-\alpha)x+\alpha y)^{2}\geq\alpha(x^{2}-y^{2})$.
-Proof.
-It follows that
-$x^{2}-((1-\alpha)x+\alpha y)^{2}$ $=x^{2}-(1-\alpha)^{2}x^{2}-\alpha^{2}y^{2}-2\alpha(1-\alpha)xy$
-$=(2\alpha-\alpha^{2})x^{2}-\alpha^{2}y^{2}-2\alpha(1-\alpha)xy$
-$=\alpha(x^{2}-y^{2})+(\alpha-\alpha^{2})x^{2}+(\alpha-\alpha^{2})y^{2}-2(\alpha-\alpha^{2})xy$
-$=\alpha(x^{2}-y^{2})+(\alpha-\alpha^{2})(x-y)^{2}$
-$\geq\alpha(x^{2}-y^{2}),$
-where the inequality is by $\alpha\in(0,1]$. ∎
+Proof 4.13.
+It follows directly from Theorems 4.8 and 4.10 that: (i) map $T$ is local $C^{1}$-smooth around any stationary point; (ii) all stationary points $x^{*}$ are unstable fixed points; (iii) $T$ is a lipeomorphism.
+Then by Corollary 2.15, the set of initial conditions attracted by such fixed points has zero Lebesgue measure.
+Appendix A Proof supplementary
 Lemma A.1.
@@
 $\|\mathbb{S}_{\hat{w}}(\hat{z})-\mathbb{S}_{\tilde{w}}(\tilde{z})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$
-Proof.
+Proof A.2.
+We consider the following three cases.
 (a) If $|\hat{z}_{i}|\leq\hat{w}_{i}$ and $|\tilde{y}_{i}|\leq\tilde{w}_{i}$, then $\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})=0$ and $\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})=0$.
@@
 $\geq$ $\ |\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|-|\lambda(\hat{w}_{i}-\tilde{w}_{i})/\beta|,$
-implying
-$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
-(c) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|>\tilde{w}_{i}$, then
-$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})-\text{sign}(\tilde{z}_{i})(|\tilde{z}_{i}|-\tilde{w}_{i})|.$
-If $\text{sign}(\hat{z}_{i})=\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|(|\hat{z}_{i}|-\hat{w}_{i})-(|\tilde{z}_{i}|-\tilde{w}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|$.
+implying $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
+(c) If $|\hat{z}_{i}|>\hat{w}_{i}$ and $|\tilde{z}_{i}|>\tilde{w}_{i}$, then $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|\text{sign}(\hat{z}_{i})(|\hat{z}_{i}|-\hat{w}_{i})-\text{sign}(\tilde{z}_{i})(|\tilde{z}_{i}|-\tilde{w}_{i})|.$ If $\text{sign}(\hat{z}_{i})=\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=|(|\hat{z}_{i}|-\hat{w}_{i})-(|\tilde{z}_{i}|-\tilde{w}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|$.
 If $\text{sign}(\hat{z}_{i})=-\text{sign}(\tilde{z}_{i})$, $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|=(|\hat{z}_{i}|-\hat{w}_{i})+(|\tilde{z}_{i}|-\tilde{w}_{i})\leq|\hat{z}_{i}|+|\tilde{z}_{i}|=|\hat{z}_{i}-\tilde{z}_{i}|.$
-In all three cases, we show
-$|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$
-Therefore, by (A.1),
-$\|\mathbb{S}_{\hat{w}}(\hat{w})-\mathbb{S}_{\tilde{w}}(\tilde{w})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$
+In all the three cases, we show $|\mathbb{S}_{\hat{w}_{i}}(\hat{z}_{i})-\mathbb{S}_{\tilde{w}_{i}}(\tilde{z}_{i})|\leq|\hat{z}_{i}-\tilde{z}_{i}|+|\hat{w}_{i}-\tilde{w}_{i}|.$ Therefore, $\|\mathbb{S}_{\hat{w}}(\hat{w})-\mathbb{S}_{\tilde{w}}(\tilde{w})\|\leq\|\hat{z}-\tilde{z}\|+\|\hat{w}-\tilde{w}\|.$